### v1

# Causality Pursuit from Heterogeneous Environments via Neural Adversarial Invariance Learning

Abstract
Statistics suffers from a fundamental problem, “the curse of endogeneity” – the regression function, or more broadly the prediction risk minimizer with infinite data, may not be the target we wish to pursue. This is because when complex data are collected from multiple sources, the biases deviated from the interested (typically causal) association inherited in individuals or sub-populations are not expected to be canceled. Traditional remedies are of hindsight style and restrictive in being tailored to prior knowledge like untestable cause-effect structures, resulting in methods that risk severe model misspecification and lack scalable applicability. This paper seeks to offer a purely data-driven and universally applicable method that only uses the heterogeneity of the biases in the data themselves rather than following pre-offered and task-specific commandments. Such an idea is formulated as a nonparametric invariance pursuit problem, whose goal is to unveil the invariant conditional expectation $m^{\star}(x)\equiv\mathbb{E}[Y^{(e)}|X_{S^{\star}}^{(e)}=x_{S^{\star}}]$ with unknown important variable set $S^{\star}$ across heterogeneous environments $e\in\mathcal{E}$. Under the structural causal model framework, $m^{\star}$ can be interpreted as pragmatic causality in general and it will coincide with the exact causal mechanism when there is enough heterogeneity. The paper contributes to proposing a novel framework, Focused Adversarial Invariance Regularized (FAIR) Estimation, formulated as a single minimax optimization program that can solve the general invariance pursuit problem. The adversary model in the framework serves the role of driving the predictor model towards a prediction-invariant solution adaptively. As illustrated by the unified non-asymptotic analysis, our adversarial estimation framework can attain provable sample-efficient estimation akin to standard regression under a minimal identification condition for various tasks and models. As an application, the Focused Adversarial Invariance Regularized Neural Network (FAIR-NN) estimator realized by two neural network classes is highlighted as the first approach to attain statistically efficient estimation in general nonparametric invariance learning.
Keywords: Adversarial Estimation, Causal Discovery, Conditional Moment Restriction, Gumbel Approximation, Invariance, Neural Networks.
1 Introduction
A fundamental problem in statistics and machine learning is to predict the response variable $Y$ based on explanatory covariates denoted as $X\in\mathbb{R}^{d}$. The objective often centers on estimating the regression function $m_{0}(x)=\mathbb{E}[Y|X=x]$, which minimizes the population $L_{2}$ risk
$\mathsf{R}(m)=\int|y-m(x)|^{2}\mu_{0}(dx,dy).$
based on an i.i.d. sample $\{(X_{i},Y_{i})\}_{i=1}^{n}$ drawn from $\mu_{0}$. Tracing back to pioneering works in the early nineteenth century (Legendre,, 1805; Gauss,, 1809), numerous statistical regression techniques have been developed to estimate the regression function $m_{0}$ from real-world data. By carefully balancing the bias and variance, the fitted model not only offers decent predictions for unseen data that is drawn from $\mu\approx\mu_{0}$ but also furnishes insights into the underlying mechanism between objects grounded in the real world. The above model is referred to as nonparametric regression (Fan & Gijbels,, 1996; Györfi et al.,, 2002; Tsybakov,, 2009) when it is agnostic about the functional form of $m$.
Climbing toward the hill of the efficient estimation of $m_{0}$ is challenging. When there are no additional low-dimension structures, estimating $m_{0}$ would suffer from the “curse of dimensionality” (Stone,, 1982). Unfortunately, the function structure may not often be known to us, whereas, the emergence of algorithmic (Breiman,, 2001) nonparametric techniques, e.g., neural networks, in the age of data make it possible to adaptively circumvent the curse of dimensionality. Compared with previous structural methods like additive models (Stone,, 1985) that pre-assume some low-dimensional structures, the neural network can obtain a comparable estimation error and circumvent the “curse of dimensionality” without the knowledge of the specific low-dimensional function structure ahead (Bauer & Kohler,, 2019; Schmidt-Hieber,, 2020; Fan & Gu,, 2022), and will not suffer from model misspecification. Such an algorithmic nature provides neural network estimators with scalable applicability in various prediction tasks and data structures.
Though the progress we made in climbing towards $m_{0}$ is breathtaking, a more fundamental question is – are we climbing the right hill? To be specific, is $m_{0}$ always the target we wish to pursue, especially when gathering extensive variables for fitting or potentially collecting heterogeneous data from multiple sources? In numerous scenarios, $m_{0}$ itself may offer unstable predictions on unseen data from distributions where $\mu\neq\mu_{0}$, and risks misattributions and possible false scientific conclusions. Consider the simple thought experiment where we aim to classify an object in a picture as either a cow $(Y=1)$ or a camel $(Y=0)$ using two provided features $X_{1}$ (body shape) and $X_{2}$ (background color). Given our dataset drawn from the distribution $\mu_{0}$, the cows usually appear on green grass, while camels often stay on yellow sand. Consequently, the conditional expectation $m_{0}(x_{1},x_{2})=\mathbb{E}_{\mu_{0}}[Y|X_{1}=x_{1},X_{2}=x_{2}]$ might be heavily dependent on $x_{2}$. Such a model is problematic both for prediction and attribution. Its application in a setting with a different background such as zoos would lead to unreliable predictions. Furthermore, attributing the determination of an object to the background that surrounds it also contradicts our understanding of causality. In the above case, we may prefer $m_{\star}(x)=\mathbb{E}[Y|X_{1}=x_{1}]$ for prediction and attribution as we know the underlying causal mechanisms.
We refer to the above problem as the “curse of endogeneity” in that the conditional expectation of the residual for the “potential” interested $m_{\star}$ is not zero given all the explanatory variables, i.e., $\mathbb{E}[Y-m_{\star}|X]\neq 0$, leading to a misalignment between $m_{0}$ and $m_{\star}$, i.e., $m_{0}(X)-m_{\star}(X)\neq 0$. Hence traditional nonparametric regression and machine learning for estimating $m_{0}$ will result in an unsatisfactory solution.
Existing methods to overcome the problem are mainly restricted to causal inference methods, in which some “prior” structural assumptions or a causal graph are often pre-offered before the inference. Such prior knowledge can be formally encoded in estimation within the potential outcome (Rubin,, 1974), or structural causal model (SCM) (Glymour et al.,, 2016) framework. The key idea is that for a specific causal estimation task, one can represent some causal estimand through a statistical estimand given the prior knowledge, i.e., identification conditions. The statistical estimand can be subsequently estimated using regression techniques. Examples include the average treatment effect estimation (Robins et al.,, 1994) under the unconfoundedness condition, and instrument variable regression (Reiersøl,, 1945; Heckman,, 2008) under the valid instrument condition. It is worth pointing out that all of these methods are tailored to prior task-specific and untestable knowledge like cause-effect relationships between the variables, and thus can be categorized as the “structural methods” as a remedy of the “curse of endogeneity”. Specifically, the “causality skeleton” is fully shaped by the prior commandments in the first stage, while the collected data and the statistical estimation methods are only responsible for filling the “association flesh” in the second stage. This limitation prevents the conclusion they drawn from going beyond hindsight in general and hinders them from scalable applicability because it requires domain expert knowledge for very complicated data, and risks more serious model misspecification since such knowledge is not testable and hence impossible to be falsified in general.
In this paper, we consider shaping both the “causality skeleton” and “association flesh” from the data themselves. The goal is to answer the following fundamental question:
Can we design methods that can circumvent the “curse of endogeneity” without thesupervision of cause-effect $\begin{split}&\text{\it Can we design methods that can circumventthe ``curse of endogeneity'' without the}\\ &\text{\it supervision of cause-effect knowledge?}\end{split}$ (Q)
Without the supervision of prior causal structural knowledge, we leverage the principle of how humans understand causality: the causal association is the one that consistently occurs in the past, now, and (potentially) in the future, or more broadly, in diverse environments. The main contribution of this paper is to propose a unified, algorithmic methodological framework to circumvent the “curse of endogeneity” in a purely data-driven way. This algorithmic nature provides it with scalable applicability and is free of model misspecification in a similar manner to neural networks as the remedy of the “curse of dimensionality”.
1.1 The Canonical Model under Study
Let us revisit the thought experiment from the perspective of a hyper-intelligent alien, Alice. Alice knows nothing about cows and camels except for 1000 images with annotated labels highly associated with the background, for example, $r=90\%$ cows/camels on grass/sand. It’s impossible for her to claim that the background cannot determine the object given this limited information. However, if she receives another set of 1000 images, where $r=70\%$ cows/camels on grass/sand, she might begin to question the determinative role of the background. Her skepticism will intensify when more and more datasets with varying $r$ are offered.
When there is no supervision of the cause-effect relationship, the observation from heterogeneous sources is essential in the nonparametric setting. In this context, this paper considers the following multi-environment regression model. Let $\mathcal{E}$ be the set of sources/environments. For each environment $e\in\mathcal{E}$, we observe $n$ [^1] We point out that the same $n$ is just for expository simplicity, the extension to varying $n^{(e)}$ is straightforward. i.i.d. pairs $(X_{1}^{(e)},Y_{1}^{(e)}),\ldots,(X_{n}^{(e)},Y_{n}^{(e)})\sim\mu^{(e)}$, where $\mu^{(e)}$, the joint distribution of $(X^{(e)},Y^{(e)})$, satisfies
$Y^{(e)}=m^{\star}(X_{S^{\star}}^{(e)})+\varepsilon^{(e)}\qquad\text{with}\qquad\mathbb{E}[\varepsilon^{(e)}|X_{S^{\star}}^{(e)}]\equiv 0.$ (1.1)
Here $S^{\star}$, the unknown true important variable set, and $m^{\star}:\mathbb{R}^{|S^{\star}|}\to\mathbb{R}$, the target regression function, are both invariant across different environments; the joint distributions $\mu^{(e)}$ can vary. We aim to estimate the invariant regression function $m^{\star}$ using data $\{\{(X_{i}^{(e)},Y_{i}^{(e)})\}_{i=1}^{n}\}_{e\in\mathcal{E}}$ from $|\mathcal{E}|$ heterogeneous environments. We refer to the above problem as nonparametric invariance pursuit.
Here we temporarily refrain from causal discussions and frame it as a pure statistical estimation problem. Under particular scenarios, such a problem can be instantiated to causal discovery in the SCM framework (Peters et al.,, 2016), and transfer learning with a more realistic assumption (Rojas-Carulla et al.,, 2018); see the details in Section A.1. We offer a rigorous and comprehensive interpretation of what $S^{\star}$ is in the SCM with interventions on $X$ in Section 4.2. It is also notable to mention that model (1.1) only requires invariance in the first moment instead of full distributional invariance, i.e., $\varepsilon^{(e)}\sim F_{\varepsilon}\perp\!\!\!\perp X_{S^{\star}}^{(e)}$, as typically required for causal discovery (Peters et al.,, 2016). The invariance structure in (1.1) is more realistic in that it allows for between-environment heteroscedastic errors. The relaxation we adopted broadens the scope for various applications, including binary classification.
It is important to note that the objective of standard nonparametric regression generally diverges from our target $m^{\star}$, i.e., $\mathbb{E}[Y^{(e)}|X^{(e)}=x]\neq m^{\star}(x_{S^{\star}})$. This discrepancy arises because the exogeneity condition $\mathbb{E}[\varepsilon^{(e)}|X_{S^{\star}}^{(e)}]=0$ only applies to the true important variables, while $\mathbb{E}[\varepsilon^{(e)}|X^{(e)}]\neq 0$. Such a “curse of endogeneity” problem is the main challenge we need to address. To identify $m^{\star}$, we exploit the invariance structure
$\forall e,e^{\prime}\in\mathcal{E}\qquad\mathbb{E}[Y^{(e)}|X_{S^{\star}}^{(e)}]\equiv\mathbb{E}[Y^{(e^{\prime})}|X_{S^{\star}}^{(e^{\prime})}].$ (1.2)
1.2 Contributions of the Paper
In this paper, we propose a unified estimation framework – the Focused Adversarial Invariance Regularized (FAIR) estimator. This estimator regularizes the user-specified risk loss $\ell(y,v)$ using a novel regularizer. Specifically, the FAIR estimator is the solution of the following minimax optimization program
$\min_{g\in\mathcal{G}}\max_{f^{(e)}\in\mathcal{F}_{S_{g}},\forall e\in\mathcal{E}}\underbrace{\sum_{e\in\mathcal{E}}\mathbb{E}_{\mu^{(e)}}\left[\ell(Y,g(X))\right]}_{\mathsf{R}(g)}+\gamma\underbrace{\sum_{e\in\mathcal{E}}\mathbb{E}_{\mu^{(e)}}\left[\{Y-g(X)\}f^{(e)}(X)-\{f^{(e)}(X)\}^{2}/2\right]}_{\mathsf{J}(g,\{f^{(e)}\}_{e\in\mathcal{E}})}.$
Here $\ell(\cdot,\cdot)$ is a loss whose population solution leads to the conditional expectation, $\gamma>0$ is the regularization hyper-parameter to be determined, $(\mathcal{G},\mathcal{F})$ are the function classes to be specified by the user satisfying $\mathcal{G}\subseteq\mathcal{F}$, $\mathcal{F}_{S_{g}}=\{f\in\mathcal{F}:f(x)=h(x_{S_{g}})\text{ for some }h:\mathbb{R}^{|S_{g}|}\to\mathbb{R}\}$, and $S_{g}=\mathrm{supp}(g)$ represents the index set of the variables $g$ depends on. The function class $\mathcal{F}$ serves as the testing function class for the prediction functions in $\mathcal{G}$. From a high-level viewpoint, solving the minimax optimization program of $\mathsf{J}$ promotes the invariance in a similar spirit to (1.1), while minimizing $\mathsf{R}$ with respect to $g$ prevents it from collapsing to conservative solutions. For the empirical counterpart, we solve a similar minimax optimization program that substitutes $\mathbb{E}_{\mu^{(e)}}[\cdot]$ with the corresponding sample means. We highlight our core contributions as follows.
At the technical level, we propose a new method that can discover the invariant regression function, i.e. to solve a generalized version of the problem in Section 1.1, in an algorithmic manner. Specifically, the newly proposed FAIR estimator can be obtained by simply solving a minimax optimization program blind to the cause-effect knowledge. The method is simple – it is just one optimization objective complemented by one extra hyper-parameter $\gamma$. Moreover, it is universal – it accommodates any loss $\ell(\cdot,\cdot)$ whose unique optimal solution aligns with the conditional expectation, and can be seamlessly integrated by a variety of machine learning algorithms. Therefore, our method, which can also be regarded as a general estimation framework, homogenizes various tasks such as regression and classification, and diverse models including neural networks (Barron,, 1994) and the spline-based MARS algorithm (Friedman,, 1991). The framework encapsulates previously studied methods as particular instances. For example, with both $\mathcal{F}$ and $\mathcal{G}$ as linear function classes, it turns into a stabilized variant of the EILLS estimator (Fan et al.,, 2023). When $|\mathcal{E}|=1$, it is similar to the FGMM estimator (Fan & Liao,, 2014) under a certain choice of $(\mathcal{G},\mathcal{F})$. We also conduct a unified non-asymptotic theoretical analysis for our method of estimation. Our main result extends the theoretical findings in Fan et al., (2023) to a broad spectrum of $(\mathcal{G},\mathcal{F},\ell)$ specifications.
As a special instance in our framework, the FAIR neural network (FAIR-NN) estimator is proposed for which $\mathcal{G}$ and $\mathcal{F}$ are neural network classes to unveil $m^{\star}$ in the above nonparametric invariance pursuit problem. To the best of our knowledge, it is the first theoretically guaranteed estimator that can efficiently recover $m^{\star}$ under a single general and minimal identification condition associated with the heterogeneity of the environments. Specifically, the sample efficiency of FAIR-NN can be understood in several notable aspects: it requires the minimal identification condition, leading to fewer required environments; it exhibits the same $L_{2}$ error rate as if directly regressing $Y$ on known $X_{S^{\star}}$, regardless of the complexity of spurious associations; and it adapts to the unknown low-dimension structure of the invariant association $m^{\star}$ in a same manner as Kohler & Langer, (2021). In a nutshell, the FAIR-NN estimator circumvents the “curse of dimensionality” and “curse of endogeneity” simultaneously in a completely algorithmic manner, which does not rely on the prior knowledge of $m^{\star}$ structure or cause-effect relationships among variables. After attaining statistical estimation for the problem (1.1), we also exactly characterize the causal interpretation of $S^{\star}$ in the SCM with interventions setting under nondegenerate arbitrary interventions. This has further implications in robust transfer learning. A sufficient and necessary condition under which the direct causes of $Y$ can be recovered in general is also derived. These results are, to the best of our knowledge, the first in the literature and are of independent interest. This closes the crucial gap underscored in Section 1.1 and further affirms a positive answer to the previously posed Question Q to some extent.
At the methodology level, our method bridges the invariance framework (Peters et al.,, 2016) and asymmetry principle (Janzing et al.,, 2016) for observational data into a unified framework. Besides using the invariance structure to identify $m^{\star}$, our framework also allows the users to incorporate their strong prior structural knowledge such as linearity or additivity of $m^{\star}$ into the FAIR estimation. This can be realized by restricting the function class $\mathcal{G}$ within this known structure and designating $\mathcal{F}$ as a more expansive class. We also propose the FAIR-AugLinear ($\mathcal{G}$ is linear and $\mathcal{F}$ is linear augmented with one transformed feature) and FAIR-ANN ($\mathcal{G}$ is additive neural network, $\mathcal{F}$ is neural network) estimators in Appendix B which leverage the prior knowledge that $m^{\star}$ is linear or additive, respectively. We demonstrate that harnessing such strong structural knowledge can relax the condition for identification. It is worth pointing out that identification is viable even when $|\mathcal{E}|=1$ for the FAIR-AugLinear estimator in Section B.6.
1.3 Related Works
Causal Invariance and Invariant Risk Minimization. Starting from the pioneering work of Peters et al., (2016), there is considerable literature proposing methods to estimate $m^{\star}$, predominantly when $m^{\star}$ is linear. These methods broadly fall into two categories: hypothesis test-based methods and optimization-based methods. For the hypothesis test-based methods (Peters et al.,, 2016; Heinze-Deml et al.,, 2018; Pfister et al.,, 2019), the Type-I error is controlled for an estimator $\widehat{S}$ with $\mathbb{P}(\widehat{S}\subseteq S^{\star})\geq 1-\alpha$. Nonetheless, these procedures may result in missing important variables or conservative solutions like $\widehat{S}=\emptyset$ due to the inherent worst-case construction in the algorithm. Additionally, the introduction of hypothesis tests also hinders its seamless integration by machine learning algorithms, limiting their scalability. On the other hand, some optimization-based methods (Ghassami et al.,, 2017; Rothenhäusler et al.,, 2019, 2021) focus on linear $m^{\star}$ and tackle the problem under additional structures such as linear SCMs with additive interventions (Rothenhäusler et al.,, 2019). The additional structure imposed curtails its applicability to a broader nonparametric setting. There are also optimization-based methods (Pfister et al.,, 2021; Yin et al.,, 2021) designed for linear models, but they are heuristic and lack finite sample guarantees. In summary, there is still a crucial gap towards efficiently estimating $m^{\star}$ without additional assumptions on the underlying model. Although Fan et al., (2023) recently bridged this gap for linear $m^{\star}$ through an optimization-based method, it is still unclear under the general nonparametric setting. This paper is the first to attain sample-efficient estimation for the general model with non-asymptotic guarantees in terms of both $|\mathcal{E}|$ and $n$.
Arjovsky et al., (2019) considers a general task, which aims to search for a data representation such that the optimal solution given that representation is optimal across diverse environments. They propose an optimization-based approach called invariant risk minimization (IRM), with many subsequent variants proposed later. However, their method requires at least $d$ environments even for the linear model, and the improvement over standard empirical risk minimization is not clear (Rosenfeld et al.,, 2021; Kamath et al.,, 2021). Our paper is the first to offer a comprehensive theoretical analysis of general invariance learning when the representation class is $\{(x_{1},\ldots,x_{d})\to(a_{1}x_{1},\ldots,a_{d}x_{d}):a_{1},\ldots,a_{d}\in\{0,1\}\}$ and to show that sample efficient estimation is in general viable even when $|\mathcal{E}|=2$.
Causal Discovery without Active Intervention. There is considerable literature on causal discovery using observational data (Spirtes et al.,, 2000; Richardson,, 1996; Chickering,, 2002; Hyttinen et al.,, 2013, 2014), but they only work on the Markov equivalence class (Meek,, 1995; Geiger & Pearl,, 1990), and cannot identify the specific cause-effect relationship in general. Under the SCM framework, there is also considerable literature capable of determining the cause-effect direction under additional assumptions imposed under the circumstances that the algorithm can only passively observe data rather than performing intervention actively. These methods can be divided into two categories – one based on the invariance principle and one based on the asymmetry principle. The invariance-based approaches (Peters et al.,, 2016) consider collecting samples from multiple experiments where some unknown intervention may apply to the variables other than $Y$. It leverages the idea that the cause-effect mechanism will remain consistent while the reverse effect-cause association may vary to identify the cause-effect relationship. On the other hand, the asymmetry-based approaches (Shimizu et al.,, 2006; Hoyer et al.,, 2008; Zhang & Hyvärinen,, 2009; Janzing et al.,, 2012; Peters et al.,, 2014) only observe one sample of observational data and identify the cause-effect direction using the idea that the cause-effect mechanism admits a simple prior known structure, whereas its inverse does not, example includes the additive noise structure in Hoyer et al., (2008). These two principles for causal discovery seem to have been orthogonal before. To the best of our knowledge, our estimation framework is the first to offer a unified methodological perspective on these two principles. It demonstrates the ability to simultaneously leverage both principles for identification and estimation; see Section B.6.
1.4 Organization
The paper is structured as follows. We provide a generalized invariance pursuit problem in Section 2 and then present our unified estimation framework in Section 3. The main text afterward mainly focuses on the nonparametric invariance pursuit in (1.1), a special instance of our framework, to illustrate the main idea and philosophy: we show in Section 4.1 that applying our framework using neural networks can attain adaptive sample-efficient estimation therein, and show that $m^{\star}$ in (1.1) can be interpreted as a type of data-driven causality via characterizing the semantic meaning of $S^{\star}$ under the SCM with intervention setting. We provide a sketch of the abstract unified theoretical result, from which all non-asymptotic results are derived as corollaries, along with its other applications in Section 5 and defer the detailed statements to the appendix. Several extensions of our framework are discussed in Section A.3. We provide a computationally efficient implementation using variants of gradient descent and Gumbel approximation, followed by its application to the simulation and real data analysis in Section 6. All the proofs are collected in the supplemental material.
1.5 Notations
We use upper case $(X,Y,Z)$ to represent random variables/vectors and denote their instances as $(x,y,z)$. Define $[n]=\{1,\ldots,n\}$. For a vector $x=(x_{1},\ldots,x_{d})^{\top}\in\mathbb{R}^{d}$, we let $\|x\|_{2}=(\sum_{j=1}^{d}x_{j}^{2})^{1/2}$. For given index set $S=\{j_{1},\ldots,j_{|S|}\}\subseteq[d]$ with $j_{1}<\cdots<j_{|S|}$, we denote $[x]_{S}=(x_{j_{1}},\ldots,x_{j_{|S|}})^{\top}\in\mathbb{R}^{|S|}$ and abbreviate it as $x_{S}$ if there is no ambiguity. We let $a\lor b=\max\{a,b\}$ and $a\land b=\min\{a,b\}$. We use $a(n)\lesssim b(n)$, $b(n)\gtrsim a(n)$, or $a(n)=O(b(n))$ if there exists some constant $C>0$ such that $a(n)\leq Cb(n)$ for any $n\geq 3$. Denote $a(n)\asymp b(n)$ if $a(n)\lesssim b(n)$ and $a(n)\gtrsim b(n)$. In the theorem statement and proof, we will use $C$ to represent the universal constants that may vary from line to line and will use $\widetilde{C},\widetilde{C}_{1},\ldots$ to represent the constant that may depend on the other constants defined in the paper.
In the context of the multi-environment setup, consider the following notations. For each $e\in\mathcal{E}$, let $\Theta^{(e)}=L_{2}(\mu^{(e)}_{x})$ be the set of $L_{2}$ functions with respect to $\mu_{x}^{(e)}$, i.e., $\Theta^{(e)}=\{f:\int f^{2}(x)\mu_{x}^{(e)}(dx)<\infty\}$, and denote $\|f\|_{2,e}=\{\int f^{2}(x)\mu^{(e)}_{x}(dx)\}^{1/2}$. Moreover, suppose we have $n$ observations $(X_{1}^{(e)},Y_{1}^{(e)}),\ldots,(X_{n}^{(e)},Y_{n}^{(e)})\in\mathbb{R}^{d}\times\mathbb{R}$ drawn i.i.d. from $\mu^{(e)}$, we define $\mathbb{E}[f(X^{(e)},Y^{(e)})]=\int f(x,y)\mu^{(e)}(dx,dy)$ and $\widehat{\mathbb{E}}[f(X^{(e)},Y^{(e)})]=\frac{1}{n}\sum_{i=1}^{n}f(X^{(e)}_{i},Y^{(e)}_{i})$ for any $f\in\Theta^{(e)}$. We define $\bar{\mu}=\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\mu^{(e)}$, and $\Theta=L_{2}(\bar{\mu}_{x})$ equipped with the norm $\|\cdot\|_{2}=\{\int f^{2}(x)\bar{\mu}_{x}(dx)\}^{1/2}$. It is easy to verify that $\Theta=\bigcap_{e\in\mathcal{E}}\Theta^{(e)}$. We assume $\mathbb{E}[|Y^{(e)}|^{2}]<\infty$.
Let $S\subseteq[d]$ be any index set. Given a function class $\mathcal{H}\subseteq\{h:\mathbb{R}^{d}\to\mathbb{R}\}$, we define $\mathcal{H}_{S}$ be the class of functions in $\mathcal{H}$ that only depend on variables $x_{S}$, i.e., $\mathcal{H}_{S}=\{h\in\mathcal{H},h(x)\equiv u(x_{S})\text{ for some }u:\mathbb{R}^{|S|}\to\mathbb{R}~{}~{}\mu^{(e)}\text{-}a.s.\forall e\in\mathcal{E}\}$. We sometimes also write $h(x_{S})$ instead of $h(x)$ for $h\in\mathcal{H}_{S}$ since $h$ only depends on $x_{S}$. For any $h\in\mathcal{H}$, we use $S_{h}\subseteq[d]$ to represent the index set of the variables $h$ depends on. We let {}$\{\mathcal{H}\}^{k}=\{(h_{1},\ldots,h_{k}):h_{i}\in\mathcal{H}~{}\forall i\in[k]\}$. For any $(X,Y)$’s joint distribution $\nu$, we use $\nu_{x}$ to denote the marginal distribution of $X$, and $\nu_{x,S}$ to denote the marginal distribution of $X_{S}$.
Neural Networks. We will use neural networks as a scalable nonparametric technique: we adopt the fully connected deep neural network with ReLU activation $\sigma(\cdot)=\max\{0,\cdot\}$, and call it deep ReLU network for short. Let $L,N$ be any positive integer, a deep ReLU network with depth $L$ width $N$ admits the form of
$g(x)=T_{L+1}\circ\bar{\sigma}_{L}\circ T_{L}\circ\bar{\sigma}\circ\cdots\circ T_{2}\circ\bar{\sigma}_{1}\circ T_{1}(x).$ (1.3)
Here $T_{l}(z)=W_{l}z+b_{l}:\mathbb{R}^{d_{l}}\to\mathbb{R}^{d_{l+1}}$ is a linear map with weight matrix $W_{l}\in\mathbb{R}^{d_{l}\times d_{l-1}}$ and bias vector $b_{l}\in\mathbb{R}^{d_{l}}$, where $(d_{0},d_{1}\ldots,d_{L},d_{L+1})=(d,N,\ldots,N,1)$, and $\bar{\sigma}_{l}:\mathbb{R}^{d_{l}}\to\mathbb{R}^{d_{l}}$ applies the ReLU activation $\sigma(\cdot)$ to each entry of a $d_{l}$-dimensional vector. Here the equal width is for presentation simplicity.
Definition 1 (Deep ReLU network class).
[…]
2 General Invariance Pursuit from Heterogeneous Environments
In this section, we formalize the problem of invariance pursuit using data from multiple environments. This formalization admits the canonical nonparametric invariance pursuit in Section 1.1 as a special case.
Let $Y\in\mathbb{R}$ be the response variable and $X\in\mathbb{R}^{d}$ be the explanatory variable. We consider the general setting in which we have collected data from multiple environments =$\mathcal{E}=\{e_{1},\ldots,e_{|\mathcal{E}|}\}$, where $\mathcal{E}$ is the set of a finite number of environments. In each environment $e\in\mathcal{E}$, we observe $n$ i.i.d. observations $\{(X^{(e)}_{i},Y^{(e)}_{i})\}_{i=1}^{n}$ that follow from some distribution $\mu^{(e)}$. Let $\Theta_{g},\Theta_{f}\subseteq\Theta$ be the class of prediction functions and testing functions, respectively. Our goal is to estimate the underlying invariant regression function $g^{\star}\in\Theta_{g}$ satisfying the invariance structure
$\forall e\in\mathcal{E}\qquad\mathbb{E}\left[\left(Y^{(e)}-g^{\star}(X_{S^{\star}}^{(e)})\right)f(X^{(e)}_{S^{\star}})\right]=0\qquad\forall f\in[\Theta_{f}]_{S^{\star}},$
where $S^{\star}$ is the unknown set of true important variables. We refer to the above problem as invariance pursuit.
The problem of estimating $g^{\star}$ in (2) is a generalized version of the canonical nonparametric invariance pursuit with $g^{\star}=m^{\star}$ in (1.1) and $\Theta_{f}=\Theta_{g}=\Theta$. It depicts a general form and unifies several problems of interest in predecessors. For example, when $\Theta_{g}$ and $\Theta_{f}$ are all linear function classes, it reduces to the linear invariance pursuit problem, i.e., estimating $g^{\star}(x)=(\beta^{\star})^{\top}x=(\beta^{\star}_{S^{\star}})^{\top}x_{S^{\star}}$ with $\beta^{\star}\in\mathbb{R}^{d}$ satisfying $\mathrm{supp}(\beta^{\star})=S^{\star}$ in the multi-environment linear regression (Fan et al.,, 2023) with linear invariance structure
$\mathbb{E}\left[\left(Y^{(e)}-(\beta^{\star}_{S^{\star}})^{\top}X_{S^{\star}}^{(e)}\right)X_{j}^{(e)}\right]=0\qquad\forall e\in\mathcal{E},j\in S^{\star}.$ (2.1)
Another example is the augmented linear invariance pursuit where $\Theta_{g}$ is linear and $\Theta_{f}=\{f(x)=\sum_{j=1}^{d}\beta_{0,j}x_{j}+\beta_{1,j}\phi(x_{j})\}$ with some transform function $\phi:\mathbb{R}\to\mathbb{R}$. The augmented linear invariance structure that realizes (2) in this case is
$\mathbb{E}\left[\left(Y^{(e)}-(\beta^{\star}_{S^{\star}})^{\top}X_{S^{\star}}^{(e)}\right)X_{j}^{(e)}\right]=\mathbb{E}\left[\left(Y^{(e)}-(\beta^{\star}_{S^{\star}})^{\top}X_{S^{\star}}^{(e)}\right)\phi(X_{j}^{(e)})\right]=0\qquad\forall e\in\mathcal{E},j\in S^{\star}$
It coincides with the problem considered by Fan & Liao, (2014) when $|\mathcal{E}|=1$. Compared with the linear invariance pursuit that searches for the invariant best linear predictor across diverse environments while $\mathbb{E}[Y^{(e)}|X_{S^{\star}}^{(e)}]$ not necessary be linear, augmented linear invariance pursuit further leverages the structural knowledge that $\mathbb{E}[Y^{(e)}|X_{S^{\star}}^{(e)}]$ admits a near linear form, or the noise is at least uncorrelated with the transformed covariates in $S^{\star}$. Identification is possible in this case even when $|\mathcal{E}|=1$.
Similar to the discussion in Section 1.1, the main challenge here is the curse of endogeneity, that the population-level least squares minimizer in one environment $e\in\mathcal{E}$ diverges from $g^{\star}$, $g^{(e)}:=\mathop{\mathrm{argmin}}_{g\in\Theta_{g}}\mathbb{E}[|Y^{(e)}-g(X^{(e)})|^{2}]\neq g^{\star}$, the discrepancy can be arbitrary large; for a formal declaration, see Proposition 2 in Fan et al., (2023). To address this issue, we will harness the insight that the distributions of $(X,Y)$ across diverse environments capture the invariance structure (2). The central idea of this paper is to exploit both the heterogeneity among different environments, i.e., the shifts in population distributions $\mu^{(e)}$, in conjunction with the above invariance structure (2) to pinpoint the invariant regression function $g^{\star}$.
It should be noted that both $g^{\star}$ and $S^{\star}$ are determined by $(\Theta_{g},\Theta_{f})$ and $\mathcal{E}$ through the structure (2). It is required that $\partial\Theta_{g}=\{g-g^{\prime}:g,g^{\prime}\in\Theta_{g}\}\subseteq\Theta_{f}$. In the case of $\Theta_{f}=\partial\Theta_{g}$, one uses only heterogeneity among different environments, or the “invariance principle”, to identify the invariant regression function $g^{\star}$. Heterogeneous environments are essential in this case. By choosing substantially large $\Theta_{f}\supsetneq\partial\Theta_{g}$, one further injects the strong prior structural assumption that the invariant regression function lies in the class $\Theta_{g}$ rather than $\Theta_{f}\setminus\Theta_{g}$. In this case, one leverages both heterogeneity among environments, i.e., the “invariance principle”, and the mentioned prior structure knowledge, i.e., the “asymmetry principle”, to jointly identify $g^{\star}$. Only one environment may be enough for identifying $g^{\star}$ in some cases.
3 Focused Adversarial Invariance Regularized Estimator
In this section, we describe our proposed method of estimation: the Focused Adversarial Invariance Regularized Estimator. The estimation involves an optimization objective function, a linear combination of a user-designated risk function and our proposed focused adversarial invariance regularizer that exploits (2). The regularizer’s optimization involves a minimax/adversarial estimation among user-specified predictor function class $\mathcal{G}$ and discriminator function class $\mathcal{F}$ that approximate $\Theta_{g}$ and $\Theta_{f}$ respectively in Section 2. We first introduce the population-level regularizer to illustrate the underlying intuitions, which enforce the predictor to achieve fair optimal predictions, given the variables used, across diverse populations. The introduction of the complete framework follows this.
3.1 Imposing Invariance via Focused Adversarial Invariance Regularizer
We illustrate the idea of our regularizer for nonparametric invariance pursuit where $\Theta_{g}=\Theta_{f}=\Theta$. According to the discussion in Section 2, we hope to impose invariance structure akin to (1.1) on the predictor function $g$ we search in the class $\mathcal{G}$. That is, given the underlying ground truth $g$, we expect that
$\forall e\in\mathcal{E}\qquad\mathbb{E}[Y^{(e)}-g(X^{(e)})|X^{(e)}_{S_{g}}]=0.$ (3.1)
The above constraint is called conditional moment restriction, and is equivalent to
$\forall e\in\mathcal{E}\qquad\mathbb{E}\big{[}\{Y^{(e)}-g(X^{(e)})\}f(X^{(e)})\big{]}=0~{}~{}~{}~{}\forall\mbox{ testing function }f\in\Theta^{(e)}_{S_{g}}.$
Based on the above observation, we propose the following focused adversarial invariance regularizer. Let $\mathcal{F}$ be a function class that $\mathcal{F}\supseteq\mathcal{G}$. At the population level, it can be expressed as a minimax optimization objective function as
$\inf_{g\in\mathcal{G}}\sup_{f^{\mathcal{E}}\in\{\mathcal{F}_{S_{g}}\}^{|\mathcal{E}|}}\underbrace{\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\mathbb{E}\left[\big{\{}Y^{(e)}-g(X^{(e)})\big{\}}f^{(e)}(X^{(e)})-\frac{1}{2}\big{\{}f^{(e)}(X^{(e)})\big{\}}^{2}\right]}_{\mathsf{J}(g,f^{\mathcal{E}})}.$ (3.2)
The term “adversarial invariance” suggests that the candidate prediction function $g$ is penalized in a way such that the residual $Y^{(e)}-g(X^{(e)})$ defined by $g$ is uncorrelated with any function in $\mathcal{F}_{S_{g}}$. The term “focused” underscores that this penalty is applied in the direction of the selected variables of $g$. In other words, the variables that the test function $f^{(e)}$ works on are restricted within the variables used by $g$. The last $-\frac{1}{2}f^{2}$ term is incorporated to manage the scale automatically.
The above minimax optimization can be interpreted as a min-max game between the predictor $g\in\mathcal{G}$ and $|\mathcal{E}|$ discriminators $f^{\mathcal{E}}=(f^{(e)})_{e\in\mathcal{E}}\in\{\mathcal{F}\}^{|\mathcal{E}|}$. Intuitively, the discriminators from different environments act like police and attempt to discern whether the predictor satisfies the conditional moment restrictions within each environment $e\in\mathcal{E}$. On the other hand, the predictor, akin to a team of counterfeiters, strives to fool the discriminators across all the environments simultaneously.
Let us formally realize the above intuitions by showing that $\breve{g}$, the predictor in the solution of the above minimax objective with function classes $\mathcal{G}=\mathcal{F}=\Theta$, satisfies the conditional moment restriction (3.1). To this end, for a fixed $g\in\Theta$, by taking supremum over all the $f^{\mathcal{E}}\in\{\Theta_{S_{g}}\}^{\mathcal{E}}$, we obtain
$\mathsf{J}(g,\star):=\sup_{f^{\mathcal{E}}\in\{\Theta_{S_{g}}\}^{|\mathcal{E}|}}\mathsf{J}(g,f^{\mathcal{E}})$ $=\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\mathbb{E}\left[\mathbb{E}\left[\sup_{u\in\mathbb{R}}\left\{\left(\mathbb{E}[Y^{(e)}|X_{S_{g}}^{(e)}]-g(X_{S_{g}}^{(e)})\right)u-\frac{1}{2}u^{2}\right\}\Big{|}X_{S_{g}}^{(e)}\right]\right]$
$=\frac{1}{2|\mathcal{E}|}\sum_{e\in\mathcal{E}}\mathbb{E}\left[|\mathbb{E}[Y^{(e)}|X_{S_{g}}^{(e)}]-g(X_{S_{g}}^{(e)})|^{2}\right].$
Given that $\mathsf{J}(g,\star)$ is non-negative, we can assert that the global minimal value of the objective $\mathsf{J}(g,\star)$ is 00, attained by the invariant regression function $m^{\star}$ in (1.1). Consequently, we can conclude that
$\breve{g}\in\mathop{\mathrm{argmin}}_{g\in\Theta}\mathsf{J}(g,\star)\qquad\Longleftrightarrow\qquad\breve{g}(X_{S_{g}}^{(e)})=\mathbb{E}[Y^{(e)}|X_{S_{\breve{g}}}^{(e)}=x_{S_{\breve{g}}}]~{}~{}\forall e\in\mathcal{E}.$ (3.3)
At the same time, the null $m^{\emptyset}(x)=0$ can also minimize $\mathsf{J}(g,\star)$ provided $Y$ is zero-mean across $\mathcal{E}$. It is easy to see if $S^{\star}$ can be divided into $K$ groups of mutually independent variables, then there are $2^{K}$ distinct global minima of $\mathsf{J}(g,\star)$. Here is where the risk loss $\ell(\cdot,\cdot)$ comes into play: it will select the most predictive variable set that preserves the invariance structure.
In practice, it is impossible to run the optimization program over large $\Theta_{g},\Theta_{f}$ given finitely many data, we will choose $\mathcal{G}\subseteq\Theta_{g}$ and $\mathcal{F}\subseteq\Theta_{f}$ that approximate $\Theta_{g},\Theta_{f}$, respective well. For example, when $\Theta_{g}=\Theta_{f}=\Theta$, we may let $\mathcal{G}$ and $\mathcal{F}$ be neural network classes with diverging $NL$.
3.2 An Unified Estimation Framework: FAIR Estimator
Let $\ell:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ be a user-determined risk loss such that
$\frac{\partial\ell(y,v)}{\partial v}=(v-y)\psi(v)\qquad\text{and}\qquad\frac{\partial^{2}\ell(y,v)}{\partial v^{2}}>0.$ (3.4)
The constraints in (3.4) ensure that the conditional expectation aligns with the unique global minima and can be satisfied by various risk losses. Two leading examples are the least square loss $\ell(y,v)=\frac{1}{2}(y-v)^{2}$ with $\psi(v)=1$ for regression, and the cross-entropy loss $\ell(y,v)=-\log(1-v)-y\log\{v/(1-v)\}$ with $\psi(v)=1/\{v(1-v)\}$ for classification.
Given all the data $\{\{(X_{i}^{(e)},Y_{i}^{(e)})\}_{i=1}^{n}\}_{e\in\mathcal{E}}$ from heterogeneous environments together with $(\Theta_{g},\Theta_{f})$ that may encode part of the prior information when $\Theta_{g}\neq\Theta$, our proposed focused adversarial invariance regularized estimator (FAIR estimator) is the solution to the subsequent minimax optimization objective
$\widehat{g}\in\mathop{\mathrm{argmin}}_{g\in\mathcal{G}}\sup_{f^{\mathcal{E}}\in\{\mathcal{F}_{S_{g}}\}^{|\mathcal{E}|}}\underbrace{\widehat{\mathsf{R}}(g)+\gamma\widehat{\mathsf{J}}(g,f^{\mathcal{E}})}_{=:\widehat{\mathsf{Q}}_{\gamma}(g,{f}^{\mathcal{E}})}.$ (3.5)
where $\mathcal{G}\subseteq\Theta_{g}$ and $\mathcal{F}\subseteq\Theta_{f}$ are function classes that approximates $\Theta_{g}$ and $\Theta_{f}$, respectively. Here $\widehat{\mathsf{R}}(g)$, the first part of the objective $\widehat{\mathsf{Q}}_{\gamma}(g,f^{\mathcal{E}})$, is the pooled empirical risk loss. This is defined as the pooled sample mean of the user-specified loss across all the environments $\mathcal{E}$:
$\widehat{\mathsf{R}}(g)=\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\widehat{\mathbb{E}}\left[\ell(Y^{(e)},g(X^{(e)}))\right]=\frac{1}{(|\mathcal{E}|\cdot n)}\sum_{e\in\mathcal{E}}\sum_{i=1}^{n}\ell(Y^{(e)}_{i},g(X^{(e)}_{i})).$ (3.6)
Integrating this loss prevents the estimator from collapsing to conservative solutions since both 00 and $m^{\star}$ attain the global minimum of the population-level regularizer. $\gamma$ is the hyper-parameter to be determined, and $\widehat{\mathsf{J}}(g,f^{\mathcal{E}})$ is the empirical counterpart of the regularizer in (3.2), that
$\widehat{\mathsf{J}}(g,f^{\mathcal{E}})=\frac{1}{(|\mathcal{E}|\cdot n)}\sum_{e\in\mathcal{E}}\sum_{i=1}^{n}\left[\big{\{}Y^{(e)}_{i}-g(X_{i}^{(e)})\big{\}}f^{(e)}(X^{(e)}_{i})-\frac{1}{2}\big{\{}f^{(e)}(X^{(e)}_{i})\big{\}}^{2}\right].$ (3.7)
In practice, we may only get access to the approximate solution. In our theoretical analysis, we focus on the performance of the approximate solution $(\widehat{g},\widehat{f}^{\mathcal{E}})$ satisfying
$\sup_{f^{\mathcal{E}}\in\{\mathcal{F}_{S_{\widehat{g}}}\}^{|\mathcal{E}|}}\widehat{\mathsf{Q}}_{\gamma}(\widehat{g},f^{\mathcal{E}})-(\gamma+1)\delta_{\mathtt{opt}}^{2}\leq\widehat{\mathsf{Q}}_{\gamma}(\widehat{g},\widehat{f}^{\mathcal{E}})\leq\inf_{g\in\mathcal{G}}\sup_{f^{\mathcal{E}}\in\{\mathcal{F}_{S_{g}}\}^{|\mathcal{E}|}}\widehat{\mathsf{Q}}_{\gamma}(g,f^{\mathcal{E}})+(1+\gamma)\delta_{\mathtt{opt}}^{2}$ (3.8)
with some optimization error $\delta_{\mathtt{opt}}^{2}>0$, here $\gamma$ in $(1+\gamma)$ the same as that in $\widehat{\mathsf{Q}}_{\gamma}$?
Discussions and Extensions. From a high-level perspective, our proposed FAIR estimator searches for the most predictive variable set $S$ that preserves some invariance structure imposed by the specification of $(\Theta_{g},\Theta_{f})$. The FAIR estimation framework presented has several limitations: (1) the loss $\ell$ has restrictions in that the conditional expectation must uniquely minimize it; (2) the environment label is discrete; and (3) the discussion still lies within the variable selection level invariance rather than general representation level invariance. We will discuss in Section A.3 that our entire framework can be easily extended to the cases where (1) and (2) fail to hold. We add some discussions on the rationale, comparison with IRM, and extension on (3) in Section A.2.
4 Nonparametric Invariance Pursuit
In this section, we apply our proposed framework to the nonparametric invariance pursuit, whose data-generating process can be summarized as the following two conditions 4.1 and 4.2. In this case, we use deep ReLU networks with growing depth and width to realize the predictor and discriminator function class and choose the $L_{2}$ loss $\ell(y,v)=\frac{1}{2}(y-v)^{2}$, resulting in a Focused Adversarial Invariance Regularized Neural Network (FAIR-NN) least squares estimator.
From a methodological perspective, we use such a canonical estimator to illustrate the main idea and key features of our proposed framework. Firstly, one can seamlessly integrate black-box machine learning models (e.g. neural networks) into it and fully exploit these models’ sample efficiency and capability in being adaptive to low-dimension structures. Moreover, this framework allows to differentiate true important and spurious information from the covariates using only the data itself given its fully “nonparametric” nature. The above claims are supported by our theoretical analysis in Section 4.1 and Section 4.2, respectively.
From a theoretical perspective, Section 4.1 focus on the sample-efficiency of the pure statistical estimation problem nonparametric invariance pursuit: it shows that sample-efficiency (in both $n$ and $|\mathcal{E}|$) is viable under a minimal identification condition related to the heterogeneity of the environments. Meanwhile, Section 4.2 focus on the semantic meaning of $S^{\star}$ and $m^{\star}$ under the SCM with intervention setting. It shows that such an invariance structure (1.2) with some $S^{\star}$ always exists in this setting and the identification condition in Section 4.1 always holds for such a $S^{\star}$ under non-degenerate interventions. Such an identified $(m^{\star},S^{\star})$ has a simple and clear semantic meaning, and unveiling them from $\mathcal{E}$ has certain pragmatic utility. This provides a first yet comprehensive theoretical understanding of the nonparametric invariance pursuit.
To start with, we introduce some additional notations for theoretical analysis. Recall that $\mu^{(e)}$ is the joint distribution of $(X,Y)$ in environment $e$ and $\bar{\mu}=\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\mu^{(e)}$ is the pooled joint distribution. We let $m^{(e,S)}\in\Theta_{S}^{(e)}$ be the conditional expectation of $Y$ given $X_{S}$ in environment $e$, i.e., $m^{(e,S)}(x)=\mathbb{E}[Y^{(e)}|X_{S}^{(e)}=x_{S}]$. Recall that $\nu_{x,S}$ is the marginal distribution of $X_{S}$ for $(X,Y)\sim\nu$. It is easy to see that $\mu^{(e)}_{x,S}$ is absolutely continuous with respect to $\bar{\mu}_{x,S}=[\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\mu^{(e)}]_{x,S}$ for any $S\subseteq[d]$, and we use $\rho^{(e)}_{S}$ to denote the Radon–Nikodym derivative of $\mu_{x,S}^{(e)}$ with respect to $\bar{\mu}_{x,S}$. We define $\bar{m}^{(S)}(x)=\sum_{e\in\mathcal{E}}\rho^{(e)}_{S}(x_{S})m^{(e,S)}(x)$, which can be interpreted as the population-level least-squares that regress $Y$ on $X_{S}$ using all the data.
Condition 4.1 (Data Generating Process).
We collect data from $|\mathcal{E}|\in\mathbb{N}^{+}$ environments. For each environment $e\in\mathcal{E}$, we observe $(X_{1}^{(e)},Y_{1}^{(e)}),\ldots,(X_{n}^{(e)},Y_{n}^{(e)})\overset{i.i.d.}{\sim}\mu^{(e)}$.
Condition 4.2 (Invariance and Nondegenerate Covariate).
There exists some set $S^{\star}$ and $m^{\star}:\mathbb{R}^{|S^{\star}|}\to\mathbb{R}$ such that $m^{(e,S^{\star})}(x)\equiv m^{\star}(x_{S^{\star}})$ for any $e\in\mathcal{E}$. Moreover, for any $S\subseteq[d]$ with $S^{\star}\setminus S\neq\emptyset$, $\inf_{m\in\Theta_{S}}\|m-m^{\star}\|_{2}^{2}\geq s_{\min}>0$.
As stated in 4.1, we consider a more general setting where only the data within each environment are independent; data across different environments may be dependent. In 4.2, the “Invariance” condition is just a restatement of (1.1). The “Nondegenerate Covariate” condition rules out some degenerate cases, for example, $m^{\star}(x_{1})=x_{1}^{2}$ with $S^{\star}=\{1\}$ and $X_{2}=X_{1}^{4}$, or $m^{\star}(x_{1},x_{2})=x_{1}$ with $S^{\star}=\{1,2\}$, and is imposed for technical convenience. The target (invariant) regression function in nonparametric invariance pursuit is $m^{\star}$.
We also introduce the following two standard regularity conditions in nonparametric regression.
Condition 4.3 (Sub-Gaussian Response).
For any $e\in\mathcal{E}$ and $t\geq 0$, $\mathbb{P}\left[|Y^{(e)}|\geq t\right]\leq C_{y}e^{-t^{2}/(2\sigma_{y}^{2})}$, where $\sigma_{y}>0$ and $C_{y}>0$ are some constants independent of $e$ and $t$.
Condition 4.4 (Boundedness in Nonparametric Regression).
There exists some constants $b_{x}$ and $b_{m}$ such that (1) $X\in[-b_{x},b_{x}]^{d}$ $\bar{\mu}$-a.s. and (2) $\|m^{(e,S)}\|_{\infty}\leq b_{m}$ for any $S\subseteq[d]$ and $e\in\mathcal{E}$.
We consider realizing the predictor function class and discriminator function class as
$\mathcal{G}=\mathcal{H}_{\mathtt{nn}}(d,L,N,b_{m})\qquad\text{and}\qquad\mathcal{F}=\mathcal{H}_{\mathtt{nn}}(d,L+2,2N,2b_{m})$ (4.1)
for neural network architecture hyper-parameter $N,L$. A larger width or depth can be adopted for $\mathcal{F}$ in general. Our specification of width and depth for $\mathcal{F}$ here is for technical purposes, that is, to let $f\in\mathcal{F}$ approximate any $m^{(e,S)}-g$ well for any $g\in\mathcal{G}$.
4.1 Non-Asymptotic Result for FAIR-NN
Condition 4.5 (Identification for Nonparametric Invariance Pursuit).
For any $S\subseteq[d]$ such that $\bar{\mu}(\{m^{\star}\neq\bar{m}^{(S\cup S^{\star})}\})>0$, there exists some $e,e^{\prime}\in\mathcal{E}$ such that $\min\{\mu^{(e)},\mu^{(e^{\prime})}\}(\{m^{(e,S)}\neq m^{(e^{\prime},S)}\})\}>0$.
Remark 1 (Minimal Heterogeneity Condition for Identification).
The above identification condition necessitates that whenever a bias emerges when regressing $Y$ on $X_{S\cup S^{\star}}$ using least squares, there should be noticeable shifts in conditional expectation $m^{(e,S)}$ across environments. In other words, $S^{\star}$ the maximum set that preserves the invariant associations. This condition is minimal. If it is violated, it would imply
$\exists\widetilde{S}\subseteq[d]~{}\text{with}~{}\widetilde{S}\setminus S^{\star}\neq\emptyset\qquad s.t.\qquad\forall e\in\mathcal{E}~{}~{}\mathbb{E}[Y^{(e)}|X_{\widetilde{S}}^{(e)}]\equiv g(X_{\widetilde{S}}^{(e)})~{}~{}\mu^{(e)}\text{-}a.s.~{}~{}\text{with}~{}g:\mathbb{R}^{|S|}\to\mathbb{R},$
in which both set $S^{\star}$ and $\widetilde{S}$ embody the invariant conditional expectation structure, thus more environments are needed in this case to pinpoint $S^{\star}$. Such a minimal identification condition underscores that our proposed FAIR-NN estimator is “sample efficient” regarding the number of environments $|\mathcal{E}|$ required; see the discussions in Section 4.2. Notably, such an identification condition relaxes those employed in approaches using intersections like ICP (Peters et al.,, 2016; Heinze-Deml et al.,, 2018). These approaches require the shifts of conditional distributions for all the $S$ with $\bar{m}^{(S)}\neq m^{\star}$ for identifying $S^{\star}$.
Remark 2.
We claim that 4.5 can be further slightly relaxed given our algorithm searches for the most predictive variable set that preserves the invariance structure. But it is of a technical style and lacks semantic meaning; see discussions in Section A.4.
The following theorem provides an oracle-type inequality for the FAIR-NN estimator in a structure-agnostic manner.
Theorem 4.1 (Oracle-type Inequality for FAIR-NN Least Squares Estimator).
Assume 4.14.5 hold. Then $\gamma^{\star}_{\mathtt{NN}}=\sup_{S\subseteq[d]:\mathsf{b}_{\mathtt{NN}}(S)>0}(\mathsf{b}_{\mathtt{NN}}(S)/\bar{\mathsf{d}}_{\mathtt{NN}}(S))<\infty$, where
$\mathsf{b}_{\mathtt{NN}}(S)=\|m^{\star}-\bar{m}^{(S\cup S^{\star})}\|_{2}^{2}\qquad\text{and}\qquad\bar{\mathsf{d}}_{\mathtt{NN}}(S)=\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\|m^{(e,S)}-\bar{m}^{(S)}\|_{2,e}^{2}.$ (4.2)
Let $C_{1}$ be some large universal constant. Assume further that $|\mathcal{E}|\leq n^{C_{1}}$. Consider the FAIR-NN estimator that solves (3.8) with $\ell(y,u)=\frac{1}{2}(y-v)^{2}$ using $\gamma\geq 8\gamma^{\star}_{\mathtt{NN}}$ and function classes (4.1) with $L,N$ satisfying $NL\leq n^{C_{1}}$ and $N\geq 4$. Then, there exists some constant $\widetilde{C}$ depending on $(d,\sigma_{y},C_{y},b_{m},C_{1})$ such that for any $n\geq 3$,
$\|\widehat{g}-m^{\star}\|_{2}\leq\widetilde{C}(1+\gamma)\left(\delta_{\mathtt{opt}}+\max_{e\in\mathcal{E},S\subseteq[d]}\inf_{h\in[\mathcal{H}_{\mathtt{nn}}(d,L,N,b_{m})]_{S}}\|m^{(e,S)}-h\|_{2,e}+\frac{NL\log^{3/2}n}{\sqrt{n}}\right)=:{\delta}_{\mathtt{NN},1}$
occurs with probability at least $1-\widetilde{C}n^{-100}$. In addition, if $\delta_{\mathtt{NN},1}\leq\widetilde{C}^{-1}\left(1\land s_{\min}\land\{\gamma\inf_{S:\bar{\mathsf{d}}_{\mathtt{NN}}(S)>0}\bar{\mathsf{d}}_{\mathtt{NN}}(S)\}\right)$, then
$\|\widehat{g}-m^{\star}\|_{2}\leq\widetilde{C}\left(\delta_{\mathtt{opt}}+\max_{e\in\mathcal{E}}\inf_{h\in[\mathcal{H}_{\mathtt{nn}}(d,L,N,b_{m})]_{S^{\star}}}\|m^{\star}-h\|_{2,e}+\frac{NL\log^{3/2}n}{\sqrt{n}}\right)$
occurs with probability at least $1-\widetilde{C}n^{-100}$.
Remark 3 (Interpretation of $\mathsf{b}_{\mathtt{NN}}(S)$ and $\bar{\mathsf{d}}_{\mathtt{NN}}(S)$).
The two quantities $\mathsf{b}_{\mathtt{NN}}(S)$ and $\bar{\mathsf{d}}_{\mathtt{NN}}(S)$ can be interpreted as bias mean and bias variance, respectively. It generalizes the quantities defined in Fan et al., (2023) in a nonparametric setting. It matches those up to constant-level factors when $(X^{(e)},Y^{(e)})$ follows a multivariate Gaussian distribution for each $e\in\mathcal{E}$. We refer to $\mathsf{b}_{\mathtt{NN}}(S)$ as bias mean since it exactly characterizes the bias of the least squares estimator in the presence of spurious/endogenous variables like the background color in the thought experiment. In particular, let $\widehat{g}_{\mathtt{LSE}(S)}$ be the least squares estimator that regresses $Y$ on $X_{S}$ using all the data, applying Proposition B.1 gives
$\left|\frac{\|\widehat{g}_{\mathtt{LSE}(S)}-m^{\star}\|_{2}^{2}}{\mathsf{b}_{\mathtt{NN}}(S)}-1\right|=o_{\mathbb{P}}(1)\qquad\text{if}~{}~{}S^{\star}\subseteq S~{}\text{and}~{}\mathsf{b}_{\mathtt{NN}}(S)>0.$
We refer to $\bar{\mathsf{d}}_{\mathtt{NN}}(S)$ as the bias variance because it measures the variations of bias across environments. Specifically, when $S^{\star}\subseteq S$, the bias in environment $e$ is $(m^{(e,S)}-m^{\star})$, and $\bar{\mathsf{d}}_{\mathtt{NN}}(S)$ can be viewed as the variance of the bias concerning the uniform distribution on $\mathcal{E}$ since $\bar{\mathsf{d}}_{\mathtt{NN}}(S)=\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\|(m^{(e,S)}-m^{\star})-(\bar{m}^{(S)}-m^{\star})\|_{2,e}^{2}$. We have $\bar{\mathsf{d}}_{\mathtt{NN}}(S^{\star})=0$ by the invariance structure 4.2.
Remark 4 (Main Identification Message).
Theorem 4.1 combines the identification result, which characterizes when it is possible to consistently estimate $m^{\star}$, and the finite-sample estimation error result, which characterizes how accurately we can estimate $m^{\star}$. The main identification message disentangled from the above theorem is that if the minimal heterogeneity condition 4.5 holds, then one can consistently estimate $m^{\star}$ provided $\gamma$ is larger than some threshold $8\gamma^{\star}_{\mathtt{NN}}$ that is independent of $n$.
To present the exact $\ell_{2}$ error rate under a specific nonparametric setup, we first introduce the concept of $(\beta,C)$-smooth function.
Definition 2 ($(\beta,C)$-smooth Function).
Let $\beta=r+s$ for some nonnegative integer $r\geq 0$ and $0<s\leq 1$, and $C>0$. A $d$-variate function $f$ is $(\beta,C)$-smooth if for every non-negative sequence $\alpha\in\mathbb{N}^{d}$ such that $\sum_{j=1}^{d}\alpha_{j}=r$, the partial derivative $\partial^{\alpha}f=(\partial f)/(\partial x_{1}^{\alpha_{1}}\cdots x_{d}^{\alpha_{d}})$ exists and satisfies $|\partial^{\alpha}f(x)-\partial^{\alpha}f(z)|\leq C\|x-z\|_{2}^{s}$. We use $\mathcal{H}_{\mathtt{HS}}(d,\beta,C)$ to denote the set of all the $d$-variate $(\beta,C)$-smooth functions.
One significant advantage of neural networks over traditional nonparametric methods is their intrinsic capability for algorithmic nonparametric regression. This enables them to learn low-dimensional structures with little or no explicit guidance regarding the forms of functions (Bauer & Kohler,, 2019; Schmidt-Hieber,, 2020; Kohler & Langer,, 2021). The following corollary claims that the $\ell_{2}$ estimation error is guarantee for any $n\geq 3$, while for sufficiently large $n$, our proposed estimator can be efficiently adaptive to the hierarchical composition structure of $m^{\star}$ in the same manner. We begin by elucidating the concept of the Hierarchical Composition Model (HCM).
Definition 3 (Hierarchical Composition Model $\mathcal{H}_{\mathtt{HCM}}(d,l,\mathcal{O},C)$).
We define function class of hierarchical composition model $\mathcal{H}_{\mathtt{HCM}}(d,l,\mathcal{O},C)$ (Kohler & Langer,, 2021) with $l,d\in\mathbb{N}^{+}$, $C\in\mathbb{R}^{+}$, and $\mathcal{O}$, a subset of $[1,\infty)\times\mathbb{N}^{+}$ satisfying $\sup_{(\beta,t)\in\mathcal{O}}(\beta\lor t)<\infty$, in a recursive way as follows. Let $\mathcal{H}_{\mathtt{HCM}}(d,0,\mathcal{O},C)=\{h(x)=x_{j},j\in[d]\}$, and
[…]
$~{}~{}~{}~{}~{}g\in\mathcal{H}_{\mathtt{HS}}(t,\beta,C)\text{ with }(\beta,t)\in\mathcal{O}\text{ and }f_{i}\in\mathcal{H}_{\mathtt{HCM}}(d,l-1,\mathcal{O},C)\big{\}}.$
Following Kohler & Langer, (2021), we assume all the compositions are at least Lipschitz functions to simplify the presentation. The minimax optimal $L_{2}$ estimation risk over $\mathcal{H}(d,l,\mathcal{O},C_{h})$ is $n^{-\alpha^{\star}/(2\alpha^{\star}+1)}$, where $\alpha^{\star}=\min_{(\beta,t)\in\mathcal{O}}(\beta/t)$ is the smallest dimensionality-adjusted degree of smoothness that represents the hardest component in the hierarchical composition (Fan et al.,, 2022; Schmidt-Hieber,, 2020). Now we are ready to present the optimal rate for FAIR-NN when all the associations $m^{(e,S)}$ lies in potentially larger HCM, and $m^{\star}$ resides within smaller HCM.
Condition 4.6 (Function Complexity and Neural Network Architecture).
The following holds:
(1) $m^{(e,S)}\in\mathcal{H}_{\mathtt{HCM}}(|S|,l,\mathcal{O},C_{h})$ for any $e\in\mathcal{E}$ and $S\subseteq[d]$ with $\alpha_{0}=\inf_{(\beta,t)\in\mathcal{O}}(\beta/t)$.
(2) $m^{\star}\in\mathcal{H}_{\mathtt{HCM}}(|S^{\star}|,l,\mathcal{O}^{\star},C_{h})$ with $\alpha^{\star}=\inf_{(\beta,t)\in\mathcal{O}^{\star}}(\beta/t)$.
(3) $\delta_{\mathtt{opt}}=o(1)$, we choose $N,L$ satisfying $LN\asymp\{n(\log n)^{8\alpha^{\star}-3}\}^{\frac{1}{2(2\alpha^{\star}+1)}}$ and $(\log n)/(N\land L)=o(1)$
(4) $\max\{C_{1},d,l,C_{h},\sigma_{y},C_{y},b_{x},b_{m},\sup_{(\beta,t)\in\mathcal{O}}(\beta\lor t),\sup_{(\beta,t)\in\mathcal{O}^{\star}}(\beta\lor t)\}\leq C_{2}$ for some constant $C_{2}>0$.
Corollary 4.2 (Optimal Rate for FAIR-NN).
Under the setting of Theorem 4.1, assume further that 4.6 holds. Then, for any $n\geq 3$, with probability at least $1-\widetilde{C}n^{-100}$, the following holds
$\|\widehat{g}-m^{\star}\|_{2}\leq\widetilde{C}\left[\delta_{\mathtt{opt}}+\left(\frac{\log^{7}n}{n}\right)^{-\frac{\alpha^{\star}}{2\alpha^{\star}+1}}\right]+1_{\{n<n_{0}\}}\widetilde{C}\gamma\left[\delta_{\mathtt{opt}}+\left(\frac{\log^{7}n}{n}\right)^{-\frac{\alpha_{0}}{2\alpha^{\star}+1}}\right],$ (4.3)
where $n_{0}$ is a constant dependent on $(C_{2},\delta_{\mathtt{opt}},\gamma,s_{\min},\inf_{S:\bar{\mathsf{d}}_{\mathtt{NN}}(S)>0}\bar{\mathsf{d}}_{\mathtt{NN}}(S))$, and $\widetilde{C}$ is a constant dependent on $C_{2}$ but independent of $\gamma,\delta_{\mathtt{opt}}$ and $n$.
The error bound (4.3) is applicable for any $n\geq 3$. Notably, the error bound will not inflate if the signal of important variables $s_{\min}$ and the heterogeneity signal $\inf_{S\subseteq[d]:\bar{\mathsf{d}}_{\mathtt{NN}}(S)>0}\bar{\mathsf{d}}_{\mathtt{NN}}(S)$ is small. Though the general error bound scales linearly with $\gamma$, the estimator we propose is not vulnerable to “weak spurious” variables, for example, $x_{j}$ where $\|m^{(e,S^{\star}\cup\{j\})}-m^{\star}\|_{2,e}\leq\epsilon$ for any $e\in\mathcal{E}$, provided all the ratio of the bias $\mathsf{b}_{\mathtt{NN}}(S)$ to heterogeneity $\bar{\mathsf{d}}_{\mathtt{NN}}(S)$ gets controlled. When $n$ larger than some constant $n_{0}$ that may inflate as the signals $(s_{\min},\inf_{S:\bar{\mathsf{d}}_{\mathtt{NN}}(S)>0}\bar{\mathsf{d}}_{\mathtt{NN}}(S))$ goes to zero, the second term in (4.3) disappears and the rate is (1) independent of $\gamma$ and $\alpha_{0}$; and (2) only dependent on the $\alpha^{\star}$.
Utilizing neural network class for both predictor and discriminator function classes allows the estimator to adapt to the invariant regression function $m^{\star}$ efficiently from two crucial perspectives. Firstly, similar to using neural networks in nonparametric regression (Bauer & Kohler,, 2019; Kohler & Langer,, 2021; Schmidt-Hieber,, 2020), implementing the predictor function using neural network endows the estimator with the capability of being adaptive to the low-dimensional hierarchical structure algorithmically. Secondly, the choice of model parameter $(N,L)$, and the convergence rate only depends on $m^{\star}$. The (spurious) conditional expectations $m^{(e,S)}$ can be much more complex than $m^{\star}$. Notably, this complexity will not affect the convergence rate. This can be credited to the scalability of neural networks used as discriminators, i.e., their adaptivity capability in the role of regularization in FAIR.
At the same time, though we have to choose a hyper-parameter $\gamma$ larger than a certain threshold to attain such a rate, the convergence rate is independent of $\gamma$. This implies that when there is a lot of data available, we do not need to tune the hyper-parameter $\gamma$ for optimal performance carefully. Instead, we can choose some conservative (large) $\gamma$ such that the lower bound requirement $\gamma\geq 8\gamma^{\star}_{\mathtt{NN}}$ is guaranteed.
The statistical rate of convergence (4.3) is minimax optimal (up to logarithmic factors). It is noteworthy that this rate matches the case where we have specific prior cause-effect knowledge such that all the variables leading to a biased estimation ($\bar{m}^{(S\cup S^{\star})}\neq m^{\star}$) are all screened out before fitting the model. We impose only one minimal 4.5 besides invariance assumption and standard assumptions used in nonparametric regression to achieve this. The discussion in the following section asserts that this assumption can be easily satisfied with $S^{\star}$ coinciding with the direct causes of $Y$ in the setting of SCM with interventions on $X$ even when $|\mathcal{E}|=2$, which highlights the capability of our proposed estimator to unveil the causality from multi-source observational data in a complete algorithmic manner, which can learn the causal mechanism between the target variable and its direct causes with no explicit supervision of cause-effect relationship and knowledge of intervention.
4.2 The Semantics of $S^{\star}$ in the SCM with Interventions Setting
The results in Section 4.1 are for the problem nonparametric invariance pursuit itself. In a population-level view, it shows that if $S^{\star}$ satisfies 4.2 and 4.5, or in other (informal) words, $S^{\star}$ is the “maximum invariant set” in that it satisfies
$m^{(e,S^{\star})}\equiv\bar{m}^{(S^{\star})}~{}(\text{invariant})\qquad\text{and}\qquad\forall S\subseteq[d],~{}m^{(e,S)}\equiv\bar{m}^{(S)}\Longrightarrow\bar{m}^{(S\cup S^{\star})}=\bar{m}^{(S^{\star})}~{}(\text{maximum})$
simultaneously, the sample-efficiency estimation can be attained by our proposed FAIR-NN estimator. It is natural to ask
Does such a maximum invariant set $S^{\star}$ exist? What’s the semantic meaning of it?
This section is devoted to offering a clean yet general answer to the problem under the SCM with intervention (on $X$) setting. The short answer is: Yes, and it is a type of data-driven causality.
To start with, we first introduce the concept of the structural causal model (Glymour et al.,, 2016).
Definition 4 (Structural Causal Model).
A structural causal model $M=(\mathcal{S},\nu)$ on $p$ variables $Z_{1},\ldots,Z_{p}$ can be described using $p$ assignment functions $\{f_{1},\ldots,f_{p}\}=\mathcal{S}$:
$Z_{j}\leftarrow f_{j}(Z_{\mathtt{pa}(j)},U_{j})\qquad j=1,\ldots,p,$
where $\mathtt{pa}(j)\subseteq\{1,\ldots,p\}$ is the set of parents, or the direct causes, of the variable $Z_{j}$, and the joint distribution $\nu(du)=\prod_{j=1}^{p}\nu_{j}(du_{j})$ over $p$ independent exogenous variables $(U_{1},\ldots,U_{p})$. For a given model $M$, there is an associated directed graph $G(M)=(V,E)$ that describes the causal relationships among variables, where $V=[p]$ is the set of nodes, $E$ is the edge set such that $(i,j)\in E$ if and only if $i\in\mathtt{pa}(j)$. $G(M)$ is acyclic if there is no sequence $(v_{1},\ldots,v_{k})$ with $k\geq 2$ such that $v_{1}=v_{k}$ and $(v_{i},v_{i+1})\in E$ for any $i\in[k-1]$.
As in Peters et al., (2016), we consider the following data-generating process in $|\mathcal{E}|$ environments. For each $e\in\mathcal{E}$, the process governing $p=d+1$ random variables $Z^{(e)}=(Z_{1}^{(e)},\ldots,Z_{d+1}^{(e)})=(X_{1}^{(e)},\ldots,X_{d}^{(e)},Y^{(e)})$ is derived from an SCM $M^{(e)}(\mathcal{S}^{(e)},\nu)$, whose induced graph $G(M^{(e)})$ is acyclic, and structural assignments are as follows,
$\begin{split}X_{j}^{(e)}&\leftarrow f_{j}^{(e)}(Z_{\mathtt{pa}(j)}^{(e)},U_{j}),\qquad\qquad j=1,\ldots,d\\ Y^{(e)}&\leftarrow f_{d+1}(X_{\mathtt{pa}(d+1)}^{(e)},U_{d+1}).\end{split}$ (4.4)
Here the distribution of exogenous variables $(U_{1},\ldots,U_{d+1})$, the cause-effect relationship $G$, and the structural assignment $f_{d+1}$ are invariant across $e\in\mathcal{E}$, while the structural assignments for $X$ may vary among $e\in\mathcal{E}$. We use superscript $(e)$ to highlight this heterogeneity. This heterogeneity may arise from performing arbitrary intervention on the variables $X$. We use $Z_{\mathtt{pa}(j)}$ to emphasize that $Y$ can be the direct cause of some variables in the covariate vector. See an example in Fig. 1 (a).
To present the result, we consider an augmented SCM that incorporates the environment label $e$ as a variable $E$. We consider the case where =$\mathcal{E}=\{0,\ldots,|\mathcal{E}|-1\}$. We let 00 be the observational environment, and the rest are the interventional environments where some unknown, arbitrary interventions are applied to the variables in some given set (but is agnostic to our estimation method) $I\subseteq[d]$. The shared cause-effect relationships in all the environments are encoded by $G$, or $\{\mathtt{pa}(j)\}_{j=1}^{d}$, based on which we define $\mathtt{ch}(k):=\{j:k\in\mathtt{pa}(j)\}$ as the set of children of variable $k$ and $\mathtt{at}(k)$ as the set of all the ancestors of the variable $Z_{k}$, defined recursively as $\mathtt{at}(k)=\mathtt{pa}(k)\cup\cup_{j\in\mathtt{pa}(k)}\mathtt{at}(j)$ in the topological order of $G$.
We can define a unified SCM $\widetilde{M}=(\widetilde{\mathcal{S}},\widetilde{\nu})$ on $d+2$ variables $Z=(Z_{1},\ldots,Z_{d},Z_{d+1},Z_{d+2})=(X_{1},\ldots,X_{d},Y,E)$ encoding all the information of $|\mathcal{E}|$ models $M^{(0)}(\mathcal{S}^{(0)},\nu),\ldots,M^{(|\mathcal{E}|-1)}(\mathcal{S}^{(|\mathcal{E}|-1)},\nu)$ defined in (4.4). Here $\widetilde{\nu}(du_{1},\ldots,du_{d+2})=\nu(du_{1},\ldots,du_{d+1})\nu_{b}(du_{d+2})$ with $\nu_{b}\sim\mathrm{Uniform}(\mathcal{E})$, and the structural assignments $\widetilde{\mathcal{S}}=\{\widetilde{f}_{1},\ldots,\widetilde{f}_{d+2}\}$ are defined as follows:
$\begin{split}E&\leftarrow\widetilde{f}_{d+2}(U_{d+2}):=U_{d+2}\\ X_{j}&\leftarrow\begin{cases}\widetilde{f}_{j}(X_{\mathtt{pa}(j)},U_{j}):=f^{(0)}_{j}(X_{\mathtt{pa}(j)},U_{j})&\qquad\forall j\in[d]\setminus I\\ \widetilde{f}_{j}(X_{\mathtt{pa}(j)},E,U_{j}):=f^{(E)}_{j}(X_{\mathtt{pa}(j)},U_{j})&\qquad\forall j\in I\end{cases}\\ Y&\leftarrow\widetilde{f}_{d+1}(X_{\mathtt{pa}}(d+1),U_{d+1}):=f_{d+1}(X_{\mathtt{pa(d+1)}},U_{d+1}).\end{split}$ (4.5)
It should be noted that throughout this section, the direct cause map $\mathtt{pa}:[d+1]\to[d+1]$ matches the causal relationship $G$ instead of $\widetilde{G}=G(\widetilde{M})$. See a graphical illustration of the above construction in Fig. 1.
[…]
Figure 1: (a) is an illustration of the two-environment model, the SCMs in the two environments share the same associated graph, that is also plotted in (b): $M^{(0)}$ is an observational environment, and $M^{(1)}$ is an interventional environment where some unknown intervention is applied to $(X_{4},X_{6},X_{7})$, $M^{(0)}$ and $M^{(1)}$ are defined as (4.4). (b) visualizes $\widetilde{G}$, the associated graph of $\widetilde{M}$ constructed based on $(M^{(0)},M^{(1)})$ and (4.5).
We summarize the above construction as a condition.
Condition 4.7 (SCM with Interventions on $X$).
Suppose $M^{(0)},\ldots,M^{(|\mathcal{E}|-1)}$ are defined by (4.4), and $G$ is acyclic. Let $\widetilde{M}$ be the model constructed as (4.5) by $\{M^{(e)}\}_{e\in\mathcal{E}}$ with $I$ be given set of variables intervened.
This section aims to characterize (1) what $S^{\star}$ would satisfy (1.1) and 4.5 given a fixed intervention set $I$, and (2) how large $I$ should be to recover the $Y$’s direct causes under arbitrary types of interventions. To this end, we need the following condition ruling out some degenerated cases.
Condition 4.8 (Nondegenerate Interventions).
[…]
$\forall~{}\text{Disjoint}~{}A,B,C\subseteq[d+2],\qquad Z_{A}\perp\!\!\!\perp Z_{B}|Z_{C}~{}~{}\overset{(a)}{\Longrightarrow}~{}~{}Z_{A}\perp\!\!\!\perp_{\widetilde{G}}Z_{B}|Z_{C},$
where $Z_{A}\perp\!\!\!\perp_{\widetilde{G}}Z_{B}|Z_{C}$ means the node set $A$ and $B$ are d-separated by $C$ in the graph $\widetilde{G}$; see Definition 2.4.1 in Glymour et al., (2016) for a formal definition of $d$-separation.
The condition (2), faithfulness on the graph $\widetilde{G}$ constraining that the graph $\widetilde{G}$ truly depicts all the conditional independence relationships, is widely used in the causal discovery literature. Condition (1) is further imposed since we only leverage the information of conditional expectations instead of conditional distributions. We impose 4.8 such that the dependence on $E$ in conditional expectation of $Y$ given $X_{S}$ with any $S\subseteq[d]$ can be represented by the graph $\widetilde{G}$ itself. The imposed 4.8 rules out the possibility of some degenerated cases; see the justifications for 4.8 and some degenerated examples in Section A.5. It should be noted that our general results in Theorem 4.3 and Proposition 4.5 apply to arbitrary forms of interventions under 4.8, which is a mild condition as violation of faithfulness in 4.8 occurs with probability 00 under some suitable measure on the model (Spirtes et al.,, 2000).
Theorem 4.3 (General Identification under SCM with Interventions on $X$).
Under 4.7, there exists some set $S_{\star}$, defined as
$S_{\star}=\mathtt{pa}(d+1)\cup A(I)\cup\bigcup_{j\in A(I)}\left(\mathtt{pa}(j)\setminus\{d+1\}\right)$ (4.6)
with $A(I)=\{j:j\in\mathtt{ch}(d+1),\mathtt{at}(j)\cap\mathtt{ch}(d+1)\cap I=\emptyset\}$ such that $m^{(e,S_{\star})}\equiv\bar{m}^{(S_{\star})}:=m_{\star}$. Suppose further 4.8 holds, then 4.5 holds with $S^{\star}=S_{\star}$ and $m^{\star}=m_{\star}$.
Theorem 4.3 exactly characterizes what $S^{\star}$ in our nonparametric invariance pursuit in (1.1) is in the setting of SCM with interventions on $X$ for general and nondegenerate intervention. Notably, it doesn’t require intervention to be “sufficient”. Firstly, such a $S^{\star}$ is well-defined in that there exists one maximum set $S_{\star}$ satisfying the invariance condition (1.1) and heterogeneity condition 4.5 simultaneously. Secondly, in the SCM setting, such a $S^{\star}=S_{\star}$ can be represented in a simple way in (4.6), which lies in between the Markov blanket of the variable $Y$ and the set of $Y$’s direct causes. Note $A(I)$ can be interpreted as the “unaffected” children of $Y$ surviving from the interventions $I$. Then $S^{\star}$ is the union of (1) parents of $Y$, (2) unaffected children of $Y$; and (3) parents of these unaffected children. The size of that set $S^{\star}$ will keep decreasing when $I$ enlarges, and it will recover the set of direct causes of $Y$ when $I$ covers “root children set” $I^{\star}$ as stated in the following Proposition 4.5.
Unfortunately, $S_{\star}\supsetneq\mathtt{pa}(d+1)$ when $I^{\star}\not\subseteq I$ in general. However, $S_{\star}$ in this setup can be interpreted as the “contemporary direct causes” or “pragmatic direct causes” of $Y$ based on the observed environments. If the future interventions are made within the set $I$, then $S_{\star}$ can be regarded as the direct causes since the conditional expectation of $Y$ given $X_{S_{\star}}$ will remain invariant in a new environment $t$. Moreover, one can deploy such a predictor in unseen environments because it depicts the most predictive one among all the associations in environment 00 that remains in environment $t$. This can be formally stated in Proposition 4.4.
Proposition 4.4 (Robust Transfer Learning).
Under 4.7, for a new environment $t$ with SCM $M^{(t)}=\{\mathcal{S}^{(t)},\nu\}$ satisfying $f_{j}^{(t)}\equiv f_{j}^{(0)}$ for any $j\in[d+1]\setminus I$, i.e., only $X_{I}$ is intervened, we have $\mathbb{E}[Y^{(t)}|X_{S_{\star}}^{(t)}]\equiv\mathbb{E}[Y^{(0)}|X_{S_{\star}}^{(0)}]$ with $S_{\star}$ in (4.6). If 4.8 holds and $M^{(t)}$ satisfies a condition akin to 4.8 (see the details in Section A.6), then $S_{\star}$ is the maximum set whose conditional expectation is transferable in that for any $S\subseteq[d]$ such that $\mathbb{E}[Y^{(t)}|X_{S_{\star}\cup S}^{(t)}]\neq\mathbb{E}[Y^{(t)}|X_{S_{\star}}^{(t)}]$, one has $\mathbb{E}[Y^{(t)}|X_{S}^{(t)}]\neq\mathbb{E}[Y^{(0)}|X_{S}^{(0)}]$.
Proposition 4.5 (Direct Cause Recovery).
(Sufficiency) Under 4.7, define $I^{\star}=\{j:j\in\mathtt{ch}(d+1),\mathtt{pa}(j)\cap\mathtt{ch}(d+1)=\emptyset\}$. If 4.8 holds and $I\supseteq I^{\star}$, then 4.5 holds with $S^{\star}=\mathtt{pa}(d+1)$.
(Necessity) Moreover, if $\bar{m}^{(S^{\star}\cup S)}\neq m^{\star}$ for any $S$ with $\{j:d+1\in\mathtt{pa}(j)\}\cap S\neq\emptyset$, i.e., $Y$ does not have degenerated children, then 4.5 holds only if $I\supseteq I^{\star}$.
We refer to $I^{\star}$ as the minimal intervention set because it is the exact minimal set of variables that should be intervened on for identification in general, nondegenerated cases. The set $I^{\star}$ is determined by the cause-effect relationship $G$. In particular, $I^{\star}$ is $\{6,7\}$ for the example in Fig. 1. Notably, $X_{8}$ does not require intervention, as $X_{7}$, one of its direct causes, is included in $I^{\star}$.
5 Sketch of the Generic Result and Its Applications
The non-asymptotic results in Section 4.1 can be extended to be the result for the general FAIR estimation framework, formally as Theorem B.2, which unifies the identification condition and $L_{2}$ estimation errors for unspecific $(\Theta_{g},\Theta_{f})$ or $(\mathcal{G},\mathcal{F})$ under the least squares loss $\ell(y,v)=\frac{1}{2}(y-v)^{2}$. We sketch the main idea and informal statement here and defer the complete result and applications besides Theorem 4.1 to Appendix B.
Suppose $[\Theta_{g}]_{S}$ and $[\Theta_{f}]_{S}$ are closed subspaces of $\Theta_{S}$ for any $S\subseteq[d]$ such that one can define
$\bar{g}^{(S)}(x)=\mathop{\mathrm{argmin}}_{g\in[\Theta_{g}]_{S}}\|g-\bar{m}^{(S)}\|_{2}\qquad\text{and}\qquad f^{(e,S)}(x)=\mathop{\mathrm{argmin}}_{f\in[\Theta_{f}]_{S}}\|f-m^{(e,S)}\|_{2,e}.$
In this case, the invariance structure and the invariant regression function in (2) can be simplified as
$f^{(e,S^{\star})}(x)\equiv\bar{g}^{(S^{\star})}(x):=g^{\star}(x).$ (5.1)
Similar to the nonparametric bias mean and bias variance in Remark 3, we can define the generalized bias mean and bias variance w.r.t. $(\Theta_{g},\Theta_{f})$ as $\mathsf{b}(S)=\|\bar{g}^{(S\cup S^{\star})}-g^{\star}\|_{2}^{2}$ and $\bar{\mathsf{d}}(S)=\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\|\bar{g}^{(S)}-f^{(e,S)}\|_{2,e}^{2}$. The general identification condition akin to 4.5 is
$\forall~{}S\subseteq[d],\qquad\mathsf{b}(S)>0~{}~{}\Longrightarrow~{}~{}\bar{\mathsf{d}}(S)>0.$ (5.2)
The above condition requires that whenever incorporating more variables in $S$ will lead to better prediction performance, the set $S$ will not satisfy the invariance structure (2). 4.5 instantiates (5.2) by letting $\bar{\mathsf{d}}(S)=\bar{\mathsf{d}}_{\mathtt{NN}}(S)$ and ${\mathsf{b}}(S)=\mathsf{b}_{\mathtt{NN}}(S)$ with $(\mathsf{b}_{\mathtt{NN}}(S),\bar{\mathsf{d}}_{\mathtt{NN}}(S))$ defined in (4.2).
Theorem 5.1 (Main Result for FAIR Least Squares Estimator, Informal).
Under (5.1), (5.2) and some regularity conditions in regression, one can consistently estimate $g^{\star}$ by choosing $\gamma\geq 8\sup_{S:\mathsf{b}(S)>0}\{\mathsf{b}(S)/\bar{\mathsf{d}}(S)\}$. To be specific, there exists some large constant $C_{1}$ such that for any $n\geq 3$, the FAIR estimator $\widehat{g}$ solving (3.8) with $\ell(y,v)=\frac{1}{2}(y-v)^{2}$ and $\delta_{\mathtt{opt}}=0$ satisfies,
$\frac{\|\widehat{g}-g^{\star}\|_{2}}{C_{1}}\leq\delta_{\mathtt{stoc}}+\delta_{\mathtt{approx}}^{\star}+\gamma(\delta_{\mathtt{stoc}}+\delta_{\mathtt{approx}})1\{(1+\gamma)(\delta_{\mathtt{stoc}}+\delta_{\mathtt{approx}})\geq s\}$ (5.3)
w.h.p.. Here $\delta_{\mathtt{stoc}}$ is the stochastic error characterized by the local Rademacher complexity of $\mathcal{F},\partial\mathcal{G}$ and $n$, $\delta_{\mathtt{approx}}^{\star}$ measures certain approximation error of $(\mathcal{G},\mathcal{F})$ w.r.t. $g^{\star}$, and $\delta_{\mathtt{approx}}$ measures the worst case approximation error of $(\mathcal{G},\mathcal{F})$ w.r.t. all the $\{f^{(e,S)}\}$. Here $s>0$ is the signal strength related to $\min_{S:\bar{\mathsf{d}}(S)>0}\bar{\mathsf{d}}(S)$ and $\min_{S:S^{\star}\setminus S\neq\emptyset}\inf_{g\in[\Theta_{g}]_{S}}\|g-g^{\star}\|_{2}$, while $C_{1}$ is independent of the two quantities.
[…]
Linear Linear Linear Linear None Impossible Thm B.10
Linear Linear w/ $\phi$ Linear Linear w/ $\phi$ Nearly Linear Possible Thm B.11
Linear $\Theta$ Linear NN Linear Possible Thm B.12
Additive $\Theta$ Additive NN NN Additive Impossible Thm B.9
$\Theta$ $\Theta$ NN NN None Impossible Thm 4.1
Table 1: Applications of Theorem B.2. Recall that $\Theta$ is the set of all $L_{2}(\bar{\mu}_{x})$ functions. For the function classes in columns $\Theta_{g},\Theta_{f},\mathcal{G}$ and $\mathcal{F}$, “Linear” is $\{f(x)=\sum_{j=1}^{d}\beta_{j}x_{j}\}$, “Linear w/ $\phi$” is $\{f(x)=\sum_{j=1}^{d}\beta_{j}x_{j}+\alpha_{j}\phi(x_{j})\}$, “NN” is deep ReLU network class, “Additive” is the additive functions $\{f(x)=\sum_{j=1}^{d}f_{j}(x_{j})\}$ and “Additive NN” is a structured neural network approximating additive functions. The column “Priors” indicates what prior structure knowledge is injected by the choice of $(\Theta_{g},\Theta_{f})$. For the second row, it is “nearly linear” given it only requires that the residual is uncorrelated with all the $\phi(x_{j})$ with $j\in S^{\star}$; the prior for the third row is exactly linear provided $\Theta_{f}=\Theta$. The column “$|\mathcal{E}|=1$ Ident” indicates whether identification for $S^{\star}$ in (2) is possible with only one environment.
The complete and rigorous statement is deferred to Theorem B.2 in Section B.1, one can also consider general loss function $\ell$; see Theorem B.3. These generic results can characterize several advantages in sample efficiency of our FAIR framework. Firstly, the error (5.3) is structure-agnostic in that it is represented by the sum of approximation error and stochastic error, indicating that (1) our framework can fully exploit the capability of $(\mathcal{G},\mathcal{F})$ in learning low-dimensional structures, and (2) it has almost no additional cost in sample efficiency compared with standard regression. Moreover, the error rate applies to any $n$, implying the estimation error is guaranteed even when it selects the wrong variable, especially when the signal $s$ is weak. Finally, though a large enough regularization hyper-parameter $\gamma$ is needed to guarantee consistent estimation, the error will be free of $\gamma$ when $n$ is large enough. For completeness, we also apply our unified result to a wide specification of $(\mathcal{G},\mathcal{F})$, including the $\mathcal{E}$-non-asymptotic result in identification and $n$-non-asymptotic result in convergence rate. See a summary in Table 1.
6 Experiments
6.1 An End-to-End Implementation
We realize the minimax optimization using gradient descent ascent, a similar approach adopted in GAN (Goodfellow et al.,, 2014) training. The main challenge here is how to do “focused regularization” which enforces $f^{(e)}\in\mathcal{F}_{S_{g}}$. Here we consider a re-parameterization trick that disentangles the function $g$ and the the variable $S_{g}$ it selects. To start with, we can write $g(x)=g(a\odot x)=g(x_{1}a_{1},\ldots,x_{d}a_{d})$ with $a\in\{0,1\}^{d}$ where $a$ is another parameter to be optimized. Then the objective (3.5) can be written as
$(\widehat{g},\widehat{a})\in\mathop{\mathrm{argmin}}_{g\in\mathcal{G},a\in\{0,1\}^{d}}\sup_{f^{\mathcal{E}}\in\{\mathcal{F}\}^{|\mathcal{E}|}}\widehat{\mathsf{R}}(g(a\odot\cdot))+\gamma\widehat{\mathsf{J}}(g(a\odot\cdot),f^{\mathcal{E}}(a\odot\cdot))$
A naive implementation for the above optimization objective is to first enumerate all the possible $a\in\{0,1\}^{d}$ and then do gradient descent ascent for given $a$, but that is computationally inefficient for large $d$. We consider borrowing the idea of Gumbel softmax as a continuous approximation for the categorical variable. To be specific, let $v\in\mathrm{R}$, and let $U_{1},U_{2}$ be two independent $\mathrm{Gumbel}(0,1)$ random variables, then we have
[…]
where ${\mathrm{sig}(v)}=e^{v}/(1+e^{v})$ is the sigmoid function. Suppose that the predictor $g$ can be parameterized by the weight $\theta\in\mathbb{R}^{N_{g}}$, i.e., $\mathcal{G}=\{g(x;\theta):\mathbb{R}^{d}\to\mathbb{R}~{}\text{with}~{}\theta\in\mathbb{R}^{N_{g}}\}$ and the discriminator can be parameterized by the weight $\phi\in\mathbb{R}^{N_{f}}$, i.e., =$\mathcal{F}=\{f(x;\phi):\mathbb{R}^{d}\to\mathbb{R}~{}\text{with}~{}\phi\in\mathbb{R}^{N_{f}}\}$. We let $w\in\mathbb{R}^{d}$ be another weight used in the Gumbel module such that the following random vector serves as the stochastic approximation of the variable selection parameter $a$, that is,
$A(U,w,\tau)=(V(w_{1},\tau,U_{1,1},U_{1,2}),\ldots,V(w_{d},\tau,U_{d,1},U_{d,2}))^{\top}\in\mathbb{R}^{d}$
with $\{U_{j,1},U_{j,2}\}_{j=1}^{d}$ being i.i.d. Gumbel(0,1) random variables. One can use the following objective as the approximation of the original objective,
$(\widehat{\theta},\widehat{w})$ $\in\mathop{\mathrm{argmin}}_{\theta\in\mathbb{R}^{N_{g}},w\in\mathbb{R}^{d}}\sup_{\forall e\in\mathcal{E},\phi^{(e)}\in\mathbb{R}^{N_{f}}}\underbrace{\mathbb{E}_{A(U,w,\tau)}\left[\widehat{\mathsf{R}}(g(A\odot\cdot;\theta))+\gamma\widehat{\mathsf{J}}(g(A\odot\cdot),f^{\mathcal{E}}(A\odot\cdot;\{\phi^{(e)}\}_{e\in\mathcal{E}}))\right]}_{\mathbb{E}_{A(U,w,\tau)}[\widehat{\mathsf{L}}(A,\theta,\{\phi^{(e)}\}_{e\in\mathcal{E}})]}$ (6.2)
One can use similar implementation tricks widely used in stochastic gradient descent with Gumbel approximation that gradually anneals the Gumbel approximation hyperparameter $\tau$. We defer the formal pseudo-code Algorithm 1 to the Section C.1.
6.2 Simulations
In this section, we present the simulation result for the FAIR-Linear estimator and FAIR-NN estimator implemented by the Gumbel approximation trick and gradient descent ascent algorithm.
6.2.1 Finite Performance of FAIR-Linear Estimator
Data Generating Process. We consider the case where $|\mathcal{E}|=2$ and the data $(X^{(e)},Y^{(e)})$ in each environment $e\in\{0,1\}$ are generated from two SCMs sharing the same causal relationship between variables. For each trial, we first generate the parent-children relationship among the variables. We enumerate all the $i\in[d+1]$. For each $i\in[d+1]$, we randomly pick at most $4$ parents for the variable $Z_{i}$ from $\{Z_{1},\ldots,Z_{i-1}\}$, this step ensures that the induced graph is a DAG. We use fixed $d=70$, and let the variable $Z_{36}$ be $Y$ and the rest variables constitute the covariate $X$, that is, we let $(Z_{1},\ldots,Z_{35},Z_{36},Z_{37},\ldots,Z_{71})=(X_{1},\ldots,X_{35},Y,X_{36},\ldots,X_{70})$. We also enforce that $Y$ has at least $5$ parents and at least $5$ children. The structural assignment for each variable $Z_{j}$ is defined as
$Z_{j}^{(e)}\leftarrow\sum_{k\in\mathtt{pa}(j)}C_{j,k}^{(e)}f_{j,k}^{(e)}(Z_{k}^{(e)})+C_{j,j}^{(e)}\varepsilon_{j}$
where $(\varepsilon_{1},\ldots,\varepsilon_{71})$ are independent standard normal random variables, $f_{j,k}^{(e)}$ sampled randomly from the candidate functions $\{\cos(x),\sin(x),\sin(\pi x),x,1/(1+e^{-x})\}$, $C_{j,k}^{(e)}$ sampled from $\mathrm{Uniform}[-1.5,1.5]$ with $|C_{j,j}^{(e)}|\geq 0.5$ are different for $e\in\{0,1\}$ in general except for $j=36$: we have $f_{36,k}^{(e)}(x)=x$ and $C_{36,k}^{(0)}\equiv C_{36,k}^{(1)}$ instead. The above data generating process can be regarded as one observation environment $e=0$ and an interventional environment $e=1$ where the random and simultaneous interventions are applied to all the variables other than the variable $Y$, while the assignment from $Y$’s parent to $Y$ remains and furnishes the target regression function $m^{\star}(x)=\sum_{k\in\mathtt{pa}(36)}C_{36,k}^{(e)}x_{k}$ in pursuit. In this case, we let $S^{\star}=\mathtt{pa}(36)$ and $\beta^{\star}$ with support set $S^{\star}$ be such that $\beta_{j}^{\star}=C^{(0)}_{36,k}=C^{(1)}_{36,k}$ for any $k\in S^{\star}$. We also let the noise variance be different for the two environments, i.e., $C_{36,36}^{(0)}\neq C_{36,36}^{(1)}$. Now the model only has conditional expectation invariance rather than the full conditional distribution invariance. Fig. 2 (a) visualizes the induced graph in one trial. The complex cause-effect relationships in high-dimensional variables make the problem of estimating $\beta^{\star}$ very challenging.
Figure 2: The visualization of (a) the SCM and (b) the ${\mathrm{sig}(w)}$ during training in one trail for FAIR-Linear estimator. We use different colors to represent the different relationships with $Y$: blue = parent, red = child, orange = offspring, lightblue = other.
Implementation. For the FAIR-Linear estimator, we realize $\mathcal{G}$ and $\mathcal{F}$ by linear function class, i.e., $\mathcal{G}=\{g(x)=\beta_{g}^{\top}x:\beta_{g}\in\mathbb{R}^{d}\}$ and =$\mathcal{F}=\{f(x)=\beta_{f}^{\top}x:\beta_{f}\in\mathbb{R}^{d}\}$, and run gradient descent ascent using Adam optimizer with a learning rate of 1e-3, batch size $64$ for $50k$ iterations. In each iteration, one gradient descent update of the parameters of the predictor $\beta_{g}$ and Gumbel logits parameters $w$ is followed by the three gradient ascent updates of the discriminators’ parameters $(\beta_{f}^{(1)},\beta_{f}^{(2)})$. We adopt fixed hyper-parameter $\gamma=36$ and report the performance of the following estimators using the median of the estimation error $\|\widehat{\beta}-\beta^{\star}\|_{2}^{2}$ over $50$ replications and varying $n\in\{200,500,1000,2000,5000\}$.
  • (1)
    Pool-LS: it simply runs least squares on the full covariate $X$ using all the data.
  • (2)
    FAIR-GB: Our FAIR-Linear estimator with Gumbel approximation that outputs $\beta_{g}\odot{\mathrm{sig}(w)}$.
  • (3)
    FAIR-RF: it selects the variables $x_{j}$ with ${\mathrm{sig}(w_{j})}>0.9$ of the fitted model in (2), i.e., $\widehat{S}=\{j:{\mathrm{sig}(w_{j})}>0.9\}$, and runs least squares again on $X_{\widehat{S}}$ using all the data.
  • (4)
    Oracle: it runs least squares on $X_{S^{\star}}$ using all the data.
  • (5)
    Semi-Oracle: it runs least squares on $X_{G^{c}}$ using all the data, where $G$ is the set of all the descendants of $Y$. Compared with the ERM, it manually removes all the variables that will lead to a biased estimation; but it will also keep uncorrelated variables compared with the full Oracle estimation.
Fig. 2 (b) visualizes how the Gumbel gate values for different covariables ${\mathrm{sig}(w)}$ evolve during training in one trail. We can see that ${\mathrm{sig}(w_{j})}$ for $j\in S^{\star}$ quickly increases and dominates the values for other variables like children/offspring of $Y$ during the whole training process.
[…]
Figure 3: The simulation results for linear models with (a) $p=70$ and (b) $p=15$. Both figures depict how the median estimation errors (based on $50$ replications, shown in log scale) for different estimators (marked with different shapes and colors) change when $n$ varies in (a) $\{200,500,1000,2000,5000\}$ and (b) $\{100,200,500,800,1000\}$, respectively.
6.2.2 Finite Performance of FAIR-NN Estimator
Data Generating Process. We consider the following data generating process with $d=26$ and $|\mathcal{E}|=2$ in each trial as
$X_{i}^{(e)}$$\leftarrow\begin{cases}\varepsilon_{i}^{(e)}&\qquad i\leq 5\\ f_{i,0}^{(e)}(Y^{(e)})+\varepsilon_{i}^{(e)}&\qquad 6\leq i\leq 9\\ \sum_{j\in\mathtt{pa}(i)\subseteq[8]}f_{i,j}^{(e)}(X_{j}^{(e)})+\varepsilon_{i}^{(e)}&\qquad 10\leq i\leq 26\end{cases}$
$Y^{(e)}$ $\leftarrow m_{k}^{\star}(X_{1}^{(e)},\ldots,X_{5}^{(e)})+\varepsilon_{0},$
where the regression function $m^{\star}$ is either $m^{\star}_{1}(x)=\sum_{k=1}^{5}m_{0,j}(x_{j})$ with random chosen $m_{0,j}$ or a hierarchical composition model $m^{\star}_{2}(x)=x_{1}x_{2}^{3}+\log(1+e^{\tanh(x_{3})}+e^{x_{4}})+\sin(x_{5})$; see detailed model and omitted implementation details in Section C.2.2. In the two environments, the cause-effect relationships are shared. The variable $Y$’s parent set is $\{1,2,3,4,5\}$, its children set is $\{6,7,8,9\}$, and may have potential descendants in $\{9,\ldots,26\}$. The above data generating process can be regarded as one observation environment $e=0$ and an interventional environment $e=1$ where the random and simultaneous interventions are applied to all the variables other than the variable $Y$, while the assignment from $Y$’s parent to $Y$ remains and furnishes the target regression function $m_{k}^{\star}(x)$ with $k\in\{1,2\}$ in pursuit. Fig. 4 (a) visualizes the induced graph in one trial.
Figure 4: The visualization of (a) the SCM and (b) the ${\mathrm{sig}(w)}$ during training in one trail for FAIR-NN estimator when $k=1$. We use different colors to represent the different relationships with $Y$: blue = parent, red = child, orange = offspring, lightblue = other.
Implementation. We let $\mathcal{G}$ be the class of ReLU neural network with depth $2$ and width $128$ and $\mathcal{F}$ be the class of ReLU neural network with depth $2$ and width $196$, and run gradient descent ascent using similar experimental configurations. We use the following empirical mean squared square computed using another $2\times n_{\mathrm{test}}=2\times 30000$ i.i.d. sampled data
$\widehat{\mathtt{MSE}}=\frac{1}{2n_{\mathrm{test}}}\sum_{e\in\mathcal{E}}\sum_{i=1}^{n_{\mathrm{test}}}\{m^{\star}(x^{(e)}_{i})-\widehat{m}(x^{(e)}_{i})\}^{2}$
as the evaluation metric. We report the median of $\widehat{\mathtt{MSE}}$ over $100$ replications for the estimators (1) – (4) akin to that for the linear model. For (1), (2), and (4), we also use ReLU neural network width depth $2$ and width $128$ in running least squares. Fig. 4 (b) also visualizes how the Gumbel gate values for different covariables ${\mathrm{sig}(w)}$ evolve during training in one trail. We can see that the training dynamics for ${\mathrm{sig}(w)}$ is much more challenging and interesting than that for the linear model depicted in Fig. 2: the weight for some $Y$’s children quickly increases at a comparable rate than the variables in $S^{\star}$ at the beginning, but such a trend slows down and finally completely reverses in the middle. We leave the rigorous and in-depth analysis behind such dynamics for future studies.
Figure 5: The simulation results for nonlinear models with (a) $m_{1}^{\star}$ and (b) $m_{2}^{\star}$. Both figures depict how the median estimation errors (based on $50$ replications) for different estimators (marked with different shapes and colors) change when $n$ varies in $\{1000,2000,3000,5000\}$ for (a) and $\{1000,2000,3000,5000,10000\}$ for (b).
Results. The results are shown in Fig. 5 and the messages are similar to those for FAIR-Linear estimators. The pooled least squares yield biased estimation, while our proposed FAIR-NN estimator can unveil the invariant association $m^{\star}$ from the two environments. Moreover, the refitted FAIR-NN estimator can obtain a near-oracle performance when $n$ is large.
Appendix
The appendix is organized as follows:
  • Appendix A
    contains the omitted discussions in the main text, including the applicable scenarios for the nonparametric invariance pursuit, some discussions and extensions on the method, and some discussions on the conditions in Section 4.
  • Appendix B
    contains the complete result that is sketched in Section 5.
  • Appendix C
    contains omitted discussions and results in experiments section.
Appendix A Omitted Discussions and Results
A.1 Applicable Scenarios for Nonparametric Invariance Pursuit
This section is devoted to providing a self-contained introduction to the motivation behind the nonparametric invariance pursuit using statements akin to previous literature (Peters et al.,, 2016; Rojas-Carulla et al.,, 2018; Fan et al.,, 2023).
Causal Discovery.
If we can expect $\mathcal{E}$ to be heterogeneous enough, recovering $S^{\star}$ in nonparametric invariance pursuit coincides with discovering the direct cause of $Y$ when the multi-environment data come from SCM with intervention on $X$ setting.
Proposition A.1.
Under the model (4.4), if we further assume that $\mathbb{E}[|Y^{(e)}|^{2}]<\infty$ for any $e\in\mathcal{E}$, then (1.2) holds with $S^{\star}=\mathtt{pa}(d+1)$.
The SCM (4.4) and Proposition A.1 extend the framework described in Peters et al., (2016) (specifically Section 4.1 and Proposition 1). This model accommodates nonlinear structural assignments. Critically, the residuals $\varepsilon^{(e)}=Y^{(e)}-\mathbb{E}[Y^{(e)}|X_{S^{\star}}^{(e)}]$, do not need to be independent of $X_{S^{\star}}^{(e)}$ or remain invariant across various environments as represented by $\varepsilon^{(e)}\sim\mu_{\varepsilon}$. Such flexibility broadens the scope for various applications, including binary classification. According to Proposition A.1, when restricted to model (4.4), a specific instantiation of our generic statistical model (1.1), identifying the true important variable set $S^{\star}$ is tantamount to pinpointing the direct cause of the target variable $Y$. Concurrently, unveiling the invariant association $m^{\star}$ aligns with uncovering the causal mechanism between $Y$ and its direct causes.
Transfer Learning. Consider we collect data $\{(X_{i}^{(e)},Y_{i}^{(e)})\}_{e\in\mathcal{E},i\in[n]}$ from $|\mathcal{E}|$ distinct sources and aim to develop a model that produces decent predictions on the data $\{X^{(t)}_{i}\}_{i\in[n_{t}]}$ in an unseen environment $t$. A significant portion of transfer learning algorithms fundamentally relies on the covariate shift assumption, represented as
$\mathbb{E}[Y^{(t)}|X^{(t)}]\equiv\mathbb{E}[Y^{(e)}|X^{(e)}]\qquad\forall e\in\mathcal{E}.$
However, as illustrated in Fan et al., (2023); Rojas-Carulla et al., (2018), it is hard for this to be true given collecting so many variables. Therefore, a more realistic assumption is that information from true important variables is transferable, articulated as $\mathbb{E}[Y^{(t)}|X^{(t)}_{S^{\star}}]=\mathbb{E}[Y^{(e)}|X^{(e)}_{S^{\star}}]$. The subsequent proposition suggests that though $m^{\star}$ might not be the optimal predictor in the unseen environment $t$, it does minimize the worst-case $L_{2}$ risk, and the associated excess risk can be decomposed as follows.
We suppose both the distribution $\mu^{(e)}$ we observed in $\mathcal{E}$ and the future distributions $\nu$ come from the following distribution family.
$\mathcal{U}_{S^{\star},m^{\star},\sigma^{2}}=\Big{\{}\mu:\mathbb{E}_{\mu}[Y^{2}]<\infty,\mathbb{E}_{\mu}[Y|X_{S^{\star}}]=m^{\star}(X_{S^{\star}}),\mathbb{E}_{\mu}[\mathrm{Var}_{\mu}(Y|X_{S^{\star}})]\lor\max_{1\leq j\leq d}\mathbb{E}_{\mu}[X_{j}^{2}]\leq\sigma^{2}\Big{\}},$
Proposition A.2.
Let $\nu\in\mathcal{U}_{S^{\star},m^{\star},\sigma^{2}}$ be arbitrary. Define $\mathsf{R}_{\mathtt{oos}}(m;\nu_{x})=\sup_{\mu\in\mathcal{U}_{S^{\star},m^{\star},\sigma^{2}},\mu_{x}\sim\nu_{x}}\mathbb{E}_{(X,Y)\sim\mu}[|Y-m(X)|^{2}]$ and $\Theta^{(t)}=L_{2}(\nu_{x})$. We have
$\forall m\in\Theta^{(t)}\qquad\mathsf{R}_{\mathtt{oos}}(m;\nu_{x})-\mathsf{R}_{\mathtt{oos}}(m^{\star};\nu_{x})=\|m-m^{\star}\|_{L_{2}(\nu_{x})}^{2}+2\sigma\|m-\widetilde{m}\|_{L_{2}(\nu_{x})},$
where $\widetilde{m}(x)=\mathbb{E}_{X\sim\nu_{x}}[m(X)|X_{S^{\star}}=x_{S^{\star}}]$. The term $2\sigma\|m-\widetilde{m}\|_{L_{2}(\nu_{x})}$ is zero when $m\in\Theta_{S^{\star}}^{(t)}$.
Given the framework described above, our proposed method solving problem in Section 1.1 can be integrated with the re-weighting technique (Gretton et al.,, 2009), a strategy addressing discrepancies within the marginal distribution of $X$, to yield reliable predictions in the previously unobserved environment $t$.
A.2 Discussion on the Methods
We provide a discussion in a question-and-response manner.
[Q] You are doing “focused regularizer” that are of combinatorial nature, can it be removed?
Answer: The short answer is No. The regularizer will be the same as running least squares if we do not enforce the discriminator using the same variables that the predictor uses. This is also the main computational difficulty in our framework and why we use randomness relaxation and Gumbel approximation in implementation. Indeed, even for linear invariance pursuit, there are certain fundamental computational limits in this such that no polynomial-time algorithm can attain consistent estimation in pursuing invariance without relying on additional structures other than invariance.
[Q] The method has a similar form to IRM, what’s the major difference?
Answer: The main difference is we should at least let $\Theta_{f}\supseteq\Theta_{g}$, such a constraint leverage the idea of over-identification and make identification possible even when $|\mathcal{E}|=2$ provided enough heterogeneity. Suppose our regularizer, which can be seen as a “correct” method to pursue condition expectation invariance, is to make $u^{(1)}=u^{(2)}$ for two $s$-dimensional parameter vectors $u^{(1)},u^{(2)}\in\mathbb{R}^{s}$, what IRM does is to let $\sum_{i=1}^{s}u^{(1)}_{i}=\sum_{i=1}^{s}u^{(2)}_{i}$. It is hard to say the latter constraint will make sense and can obtain a similar effect as the former.
[Q] Could your proposed framework be extended to the representation-level invariance like IRM?
Answer: The short answer is Yes given its algorithmic nature. But identification with two or constant-level environments is impossible now: a linear-in-dimension number of environments is required even for linear representation learning. For example, one can find some linear representation $\Phi:\mathbb{R}^{d}\to\mathbb{R}^{r}$ such that
$\mathbb{E}[Y^{(e)}|\Phi X^{(e)}]\equiv m^{\star}(\Phi X^{(e)})$
However, $|\mathcal{E}|\geq r$ is the necessary condition for identification even when the heterogeneity is enough and $r$ is pre-known to us. We conjecture that any finite number of environments $|\mathcal{E}|<\infty$ may be impossible for identification if $\Phi$ lies in some nonparametric function class.
A.3 Extensions to General Environment Variable and Loss Function
In the main text, we propose an estimation framework leveraging conditional expectation invariance with respect to discrete environment variables. It is worth noticing that our adversarial estimation framework is indeed more versatile than this: one can easily extend it to other conditional point prediction invariance with respect to more general environment covariate. We briefly discuss the direct extension here and leave a rigorous treatment as future work. In the following discussions, suppose we observe data $\{(X_{i},Y_{i},E_{i})\}_{i=1}^{n}$ drawn i.i.d. from some distribution $\mu_{0}$, where $X\in\mathbb{R}^{d}$ is the covariate we used for prediction, $Y\in\mathbb{R}$ is the target response, $E\in\mathbb{R}^{q}$ is the environment covariate we wish our prediction should be invariant with respect to.
Let $\ell(u,y):\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ be the user-defined risk whose population-level minimizer may not necessarily be conditional expectation but satisfying certain regularity conditions. Let $\ell_{u}(u,y)=\partial\ell(u,y)/\partial u$ be the partial sub-gradient with respect to the prediction. Suppose the following general invariance structure with respect to $\ell$ and environment covariate holds, that there exists $S^{\star}\subseteq[d]$ and a function $g^{\star}$ that only depends $x_{S^{\star}}$ such that
$\mathbb{E}\left[\ell_{u}(g^{\star}(X_{S^{\star}}),Y)|X_{S^{\star}},E\right]\equiv 0.$ (A.1)
It coincides with the main problem of study when $E$ is discrete and $\ell$ satisfies (3.4), but also allows for other loss and continuous environment label. Other losses include but not limited to Huber loss for robust regression, or $L_{1}$ loss for median regression.
We consider the following optimization minimax objective containing a min-max game between a predictor $g:\mathbb{R}^{d}\to\mathbb{R}$ and a discriminator $f:\mathbb{R}^{d}\times\mathbb{R}^{q}\to\mathbb{R}$:
$\min_{g\in\mathcal{G}}\max_{f\in\mathcal{F}_{S_{g}}}\underbrace{\frac{1}{n}\sum_{i=1}^{n}\ell(g(X_{i}),Y_{i})}_{\widehat{\mathsf{R}}(g)}+\gamma\underbrace{\frac{1}{n}\sum_{i=1}^{n}\left[\ell_{u}(g(X_{i}),Y_{i})f(X_{i},E_{i})-0.5\{f(X_{i},E_{i})\}^{2}\right]}_{\widehat{\mathsf{J}}(g,f)},$ (A.2)
where $\gamma$ is the hyper-parameter to be determined, and $\mathcal{F}_{S_{g}}=\{f(x,e)\in\mathcal{F},f(x,e)=w(x_{S_{g}},e)\text{ for some }w\}$. Similar to the calculation for (3.3), one can expect that minimizing the population counterpart of the focused adversarial invariance regularizer $\max_{f\in\mathcal{F}_{S_{g}}}\widehat{\mathsf{J}}(g,f)$ shares a similar nature of imposing (A.1). One can derive non-asymptotic identification and estimation error results akin to Theorem B.2 and Theorem B.3 provided strong convexity and certain Lipschitz property of the loss $\ell(u,y)$. We leave this for future studies.
A.4 Discussion on Relaxing Nonparametric Invariance Pursuit Identification Condition
Given our FAIR criterion search for the most predictive variable set whose conditional expectations remain across different environments, that is, when $\gamma\to\infty$ and $n=\infty$, our population-level objective is equivalent to the following program,
$\min_{g\in\Theta}\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\mathbb{E}\left[|Y^{(e)}-g(X^{(e)})|^{2}\right]~{}~{}s.t.~{}~{}g(x)\equiv m^{(e,S_{g})}(x)~{}~{}\forall e\in\mathcal{E}.$
We say a set $S$ is an invariant set if $m^{(e,S)}\equiv\bar{m}^{(S)}$ for any $e\in\mathcal{E}$. Therefore, one can slightly relax the identification condition as: $S^{\star}$ is the most predictive invariant set, that is,
$\forall S\subseteq[d],\qquad\text{if}~{}~{}m^{(e,S)}\equiv\bar{m}^{(S)},\qquad\text{then either }\|\bar{m}^{(S)}\|_{2}<\|\bar{m}^{(S^{\star})}\|_{2}~{}\text{or}~{}\bar{m}^{(S)}=m^{\star}.$ (A.3)
The above condition is definitely weaker than 4.5 because 4.5 essentially requires the set $S^{\star}$ is the maximum invariant set,
$\forall S\subseteq[d],\qquad\text{if}~{}~{}m^{(e,S)}\equiv\bar{m}^{(S)},\qquad\text{then }\bar{m}^{(S\cup S^{\star})}=\bar{m}^{(S^{\star})}.$ (A.4)
Here (A.4) just rewrites 4.5 in a manner similar to (A.3).
It is easy to derive results similar to Theorem 4.1 under (A.3) rather than 4.5. We can construct cases where (A.3) holds but 4.5 does not. Examples include Example 1 below with $s^{(1)}s^{(2)}=1$ under which both $\{1\}$ and $\{2\}$ are invariant set but the set $\{1,2\}$ is not. In this case, 4.5 no longer holds. However, our algorithm can still consistently estimate $m^{\star}$ provided (A.3) holds, that the variable $X_{1}$ has better prediction power. In the main text, we still adopt 4.5 instead of (A.3). The main reasons are as follows. All our discussions are under the SCM with interventions setting.
Firstly, as further shown in Section 4.2, the cases where 4.5 fails to hold are degenerate cases. When the interventions are nondegenerate, there always exists a maximum invariant set $S^{\star}$, i.e., 4.5 holds. This means, (A.3) is somewhat “marginally” weaker than 4.5.
The second reason is the lack of semantic meaning of $S^{\star}$ under this case. When the interventions are non-degenerate, $S^{\star}$ can be interpreted as “contemporary/pragmatic direct causes” that can be expressed as direct causes + unaffected children + parents of unaffected children in Proposition A.2, such a variable set also has certain robust transfer learning properties as stated in Proposition 4.4. All the above semantic meanings are valid even if the interventions are insufficient. However, when the interventions are degenerate such that 4.5 may not hold but (A.3) may hold, e.g., Example 1, all the two properties will no longer hold. If the true causal mechanism is $X_{1}\to Y\to X_{2}$, then it is possible to construct data generating process such that $S^{\star}$ can be either $\{1\}$ or $\{2\}$ in (A.3).
A.5 Discussion on the Nondegenerate Intervention Condition
The conditions (1) and (2) in 4.8 are imposed to eliminate some degenerate cases. To illustrate the intuitions why such two conditions are needed, and how such a condition will hold in general. We consider the following two examples.
Introduction of condition (1)
From a high-level viewpoint, the introduction of condition (1) is to eliminate the cases where though there are shifts in condition distributions among different environments, it happens that there are no shifts in conditional expectations. This can be illustrated in the following example.
Example 1.
Consider the following canonical model also presented in Example 4.1 in Fan et al., (2023).
$X_{1}^{(e)}$ $\leftarrow\sqrt{0.5}U_{1}$
$Y^{(e)}$ $\leftarrow X_{1}^{(e)}+\sqrt{0.5}U_{3}$
$X_{2}^{(e)}$ $\leftarrow s^{(e)}Y^{(e)}+U_{2}$
where $U_{1},U_{2},U_{3}$ are independent standard normal variables, and =$\mathcal{E}=\{1,2\}$. We let $e=1$ be the observational environment and $e=2$ be the intervention environment where the linear effect of $Y$ on $X_{2}$ are intervened ($s^{(1)}\neq s^{(2)}$). We also focus on the regime where $s^{(1)}+s^{(2)}\neq 0$ such that running least squares will lead to a biased solution.
In the above model, we can see that
$Y^{(e)}|X_{2}^{(e)}\sim\mathcal{N}\left(\frac{s^{(e)}}{(s^{(e)})^{2}+1}X_{2}^{(e)},\frac{1}{(s^{(e)})^{2}+1}\right)$
It is easy to check under the case of no-degenerated child ($s^{(1)}+s^{(2)}\neq 0$) and faithfulness on $\widetilde{M}$ ($s^{(1)}\neq s^{(2)}$). We have
[…]
or in other words, $Y\perp\!\!\!\perp E|X_{2}$. However, when $s^{(1)}=1/s^{(2)}=s$, the following holds
$\mathbb{E}[Y^{(1)}|X_{2}^{(1)}=x]=\frac{s^{(1)}}{(s^{(1)})^{2}+1}x=\frac{s}{s^{2}+1}x=\frac{s^{(2)}}{(s^{(2)})^{2}+1}x=\mathbb{E}[Y^{(2)}|X_{2}^{(2)}=x]$
The introduction of 4.8 (1) is to rule out the cases where $s^{(1)}=1/s^{(2)}=s$. And it is easy to see when $s^{(1)}$ and $s^{(2)}$ are independently generated from some prior distribution that is absolute continuous with respect to Lebesgue measure on $\mathbb{R}$, i.e., $S^{(1)},S^{(2)}\sim p_{s}$, then
$\mathbb{P}\left[S^{(1)}S^{(2)}=1\right]=0.$
Introduction of condition (2).
The condition (2), that the faithfulness condition on $\widetilde{M}$, is to eliminate the cases where though the interventions are applied, it happens that such interventions do not make an impact on the variables intervened. The following example presents such an example.
Example 2.
Consider the case where =$\mathcal{E}=\{1,2\}$, and the data generating process is as follows
$Y^{(e)}$ $\leftarrow U_{3}$
$X^{(e)}_{1}$ $\leftarrow Y^{(e)}+e+U_{1}$
$X_{2}^{(e)}$ $\leftarrow 0.5Y^{(e)}-sX^{(e)}_{1}+e+U_{2}.$
where $U_{1},U_{2},U_{3}$ are independent standard normal variables, $s\neq 0.5$ is a fixed parameter. We let $e=1$ be the observational environment and $e=2$ be the intervention environment where shifts in mean are applied to the variables $X_{1}$ and $X_{2}$.
In the above case, we have $S^{\star}=\mathtt{pa}(3)=\emptyset$, and there exists a effective simultaneous intervention on $(X_{1},X_{2})$. However, such an intervention will not affect $X_{2}$ if and only if $s=1$ because its direct effect on $X_{2}$ and the indirect effect passing through $X_{1}$ get canceled provided $s=1$. To be specific, $X_{2}^{(e)}$ can be written as
$X_{2}^{(e)}=0.5Y^{(e)}-s(Y^{(e)}+e+U_{1})+e+U_{2}=(0.5-s)Y^{(e)}-sU_{1}+U_{2}+e(1-s).$
This implies that
$Y\perp\!\!\!\perp E|X_{2}$
[…]
$\mathbb{P}\left[S=1\right]=0.$
A.6 The Complete Statement of Proposition 4.4
Specifically, we construct a unified SCM $(X,Y,E)\sim\bar{M}(\bar{\mathcal{S}},\nu)$ based on $M^{(0)}$ and new environment $M^{(t)}$ as follows:
$E$ $\leftarrow\text{Uniform}(\{0,t\})$
$X_{j}$$\leftarrow\begin{cases}\bar{f}_{j}(X_{\mathtt{pa}(j)},U_{j}):=f^{(0)}_{j}(X_{\mathtt{pa}(j)},U_{j})&\qquad\forall j\in[d]\setminus I\\ \bar{f}_{j}(X_{\mathtt{pa}(j)},E,U_{j}):=f^{(t)}_{j}(X_{\mathtt{pa}(j)},U_{j})&\qquad\forall j\in I\end{cases}$
$Y$ $\leftarrow\bar{f}_{d+1}(X_{\mathtt{pa}}(d+1),U_{d+1}):=f_{d+1}(X_{\mathtt{pa(d+1)}},U_{d+1}).$
We suppose the following condition similar to 4.8 holds in the constructed graph.
Condition A.1.
[…]
$\forall~{}\text{Disjoint}~{}A,B,C\subseteq[d+2],\qquad Z_{A}\perp\!\!\!\perp Z_{B}|Z_{C}~{}~{}\overset{(a)}{\Longrightarrow}~{}~{}Z_{A}\perp\!\!\!\perp_{\bar{G}}Z_{B}|Z_{C},$
where $Z_{A}\perp\!\!\!\perp_{\widetilde{G}}Z_{B}|Z_{C}$ means the node set $A$ and $B$ and d-separated conditioned on $C$ in the graph $\bar{G}=G(\bar{M})$.
We are ready to give a complete statement of Proposition 4.4.
Proposition A.3 (Formal Statement of Proposition 4.4).
Under the setting of Theorem 4.3, for a new environment $t$ with SCM $M^{(t)}=\{\mathcal{S}^{(t)},\nu\}$ satisfying $f_{j}^{(t)}\equiv f_{j}^{(0)}$ for any $j\in[d+1]\setminus I$, i.e., only $X_{I}$ is intervened, we also have $\mathbb{E}[Y^{(t)}|X_{S_{\star}}^{(t)}]\equiv\mathbb{E}[Y^{(0)}|X_{S_{\star}}^{(0)}]$. Suppose further that A.1 holds for the constructed SCM $\bar{M}$. Then $S_{\star}$ is the unique largest set whose conditional expectation is transferable, i.e., for any $S\subseteq[d]$ such that $\mathbb{E}[Y^{(t)}|X_{S_{\star}\cup S}^{(t)}]\neq\mathbb{E}[Y^{(t)}|X_{S_{\star}}^{(t)}]$, one has $\mathbb{E}[Y^{(t)}|X_{S}^{(t)}]\neq\mathbb{E}[Y^{(0)}|X_{S}^{(0)}]$.
A.7 Technical Related Works
Adversarial Estimation. Adversarial estimation is introduced in Goodfellow et al., (2014) for generative modeling. There is some literature on formalizing the adversarial estimation from a statistical perspective and providing statistical analysis for this family of generative modeling methods (Liang,, 2021; Singh et al.,, 2018; Biau et al.,, 2020). It is also used as a tool for estimating conditional moment models, for example, instrumental variable regression (Dikkala et al.,, 2020), estimating the (implicit) influence function (Chernozhukov et al.,, 2020; Hirshberg & Wager,, 2021).
Appendix B Generic Results and Its Applications
B.1 Main Result for the General FAIR Least Squares Estimator
This section is designed to offer a unified main result characterizing when the FAIR least squares estimator can identify the target regression function together with a non-asymptotic $L_{2}$ error bound for general $(\mathcal{G},\mathcal{F})$. To impose statistical complexity on the function classes we used, we introduce the definition of localized population Rademacher complexity, described as follows.
Definition 5 (Localized Population Rademacher Complexity).
For a given radius $\delta>0$, function class $\mathcal{H}$, and distribution $\nu$, define
$R_{n,\nu}(\delta;\mathcal{H})=\mathbb{E}_{X,\varepsilon}\left[\sup_{h\in\mathcal{H},\|h\|_{L_{2}(\nu)}\leq\delta}\left|\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}h(X_{i})\right|\right],$
where $X_{1},\ldots,X_{n}$ are i.i.d. samples from distribution $\nu$, and $\varepsilon_{1},\ldots,\varepsilon_{n}$ are i.i.d. Rademacher variables taking values in $\{-1,+1\}$ with equal probability which are also independent of $(X_{1},\ldots,X_{n})$.
Condition B.1 (Function Class).
Suppose the following holds for the function class $\mathcal{G}$ and $\mathcal{F}$ we use:
  • (1).
    It is uniformly bounded by $B\geq 1$, i.e., $\sup_{h\in\mathcal{G}\cup\mathcal{F}}\|h\|_{\infty}\leq B$.
  • (2).
    $0\in\mathcal{F}$ and the statistical complexity of the function classes $\mathcal{G}+\mathcal{F}:=\{g+f:g\in\mathcal{G},f\in\mathcal{F}_{S_{g}}\}$ is upper-bounded by $\delta_{n}$. In particular, there exists some quantity $11/n\leq\delta_{n}<$ 1 such that
    $R_{n,\mu^{(e)}}(\delta;\partial\mathcal{G})\leq B\delta_{n}\delta\qquad\text{and}\qquad R_{n,\mu^{(e)}}(\delta;\partial(\mathcal{G}+\mathcal{F}))\leq 2B\delta_{n}\delta$
    for any $e\in\mathcal{E}$ and $\delta\in[\delta_{n},2B]$, where =$\partial\mathcal{H}=\{h-h^{\prime}:h,h^{\prime}\in\mathcal{H}\}$.
Note that when $-\mathcal{G}=\mathcal{G}$, $R_{n,\mu^{(e)}}(\delta;\partial\mathcal{G})=R_{n,\mu^{(e)}}(\delta;\mathcal{G})$. The above three assumptions 4.1, 4.3, B.1 are standard in the theoretical analysis of regression. Now, we introduce the specific assumption in our multi-environment regression setting.
Condition B.2 (Invariance and Identification).
For any $S$, let $\overline{\mathcal{G}_{S}}\supseteq\mathcal{G}_{S}$, $\overline{\mathcal{F}_{S}}\supseteq\mathcal{F}_{S}$ be closed subspaces of $\Theta_{S}$ satisfying $\overline{\mathcal{G}_{S}}\subseteq\overline{\mathcal{F}_{S}}$. In this case, we can define $\Pi_{\mathcal{A}}(h)=\mathop{\mathrm{argmin}}_{a\in\mathcal{A}}\|a-h\|_{2}$ and $\Pi^{(e)}_{\mathcal{A}}(h)=\mathop{\mathrm{argmin}}_{a\in\mathcal{A}}\|a-h\|_{2,e}$ when $\mathcal{A}\in\{\overline{\mathcal{F}_{S}},\overline{\mathcal{G}_{S}}\}$ and $h\in\Theta_{S}$. Suppose the following holds:
  • 1.
    (Invariance) There exists some index set $S^{\star}\subseteq[d]$ such that
    $\forall e\in\mathcal{E}\qquad\Pi^{(e)}_{\overline{\mathcal{F}_{S^{\star}}}}(m^{(e,S^{\star})})=\Pi_{\overline{\mathcal{G}_{S^{\star}}}}(\bar{m}^{(S^{\star})}):=g^{\star}$
  • 2.
    (Heterogeneity) For each $S\subseteq[d]$, if $\mathsf{b}_{\mathcal{G}}(S)>0$, then $\bar{\mathsf{d}}_{\mathcal{G},\mathcal{F}}(S)>0$, where
    $\mathsf{b}_{\mathcal{G}}(S)=\|\Pi_{\overline{\mathcal{G}_{S\cup S^{\star}}}}(\bar{m}^{(S\cup S^{\star})})-g^{\star}\|_{2}^{2}~{}~{}\text{and}~{}~{}\bar{\mathsf{d}}_{\mathcal{G},\mathcal{F}}(S)=\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\|\Pi_{\overline{\mathcal{F}_{S}}}^{(e)}(m^{(e,S)})-\Pi_{\overline{\mathcal{G}_{S}}}(\bar{m}^{(S)})\|_{2,e}^{2}.$ (B.1)
  • 3.
    (Nondegenerate Covariate) For any $S\subseteq[d]$ such that $S^{\star}\setminus S\neq\emptyset$, we have $\inf_{g\in\overline{\mathcal{G}_{S}}}\|g-g^{\star}\|_{2}^{2}\geq s_{\min}$ for some constant $s_{\min}>0$.
The first condition “invariance” specifies the target regression function $g^{\star}$ of interests and states the invariance structure imposed for our theoretical analysis. It relaxes the general conditional expectation invariance (1.2) when $\overline{\mathcal{F}_{S}}\subsetneq\Theta_{S}$. Two leading examples are (1) the fully nonparametric class $\overline{\mathcal{G}_{S}}=\overline{\mathcal{F}_{S}}=\Theta_{S}$, and (2) linear class $\overline{\mathcal{G}_{S}}=\overline{\mathcal{F}_{S}}=\{f(x)=\beta_{S}^{\top}x_{S}:\beta_{S}\in\mathbb{R}^{|S|}\}$. In the first example, we are interested in estimating the invariant conditional expectation $g^{\star}=m^{\star}$, and the invariance condition requires the conditional expectation invariance (1.2), that
$\forall e\in\mathcal{E}\qquad m^{(e,S^{\star})}(x)=m^{\star}(x_{S^{\star}}).$
In the second example, when the covariance matrices $\mathbb{E}[X^{(e)}(X^{(e)})^{\top}]$ across all the environments are all positive definite, we are interested in estimating the invariant linear predictor $g^{\star}(x)=x^{\top}\beta^{\star}$, and such the “invariance” condition only requires that
$\forall e\in\mathcal{E}\qquad\beta^{(e,S^{\star})}\equiv\beta^{\star}\qquad\text{where}~{}~{}~{}~{}\beta^{(e,S^{\star})}=\mathop{\mathrm{argmin}}_{\beta\in\mathbb{R}^{d},\beta_{(S^{\star})^{c}}=0}\mathbb{E}[|Y^{(e)}-\beta^{\top}X^{(e)}|^{2}],$
that is, the best linear predictors constrained on $S^{\star}$ among all the environment are the same. In this case, the conditional expectations $m^{(e,S^{\star})}(x)$ can be nonlinear or different.
The second condition “heterogeneity” is for identification and is fundamental to derive the population-level strong convexity with respect to $g^{\star}$. The two quantities in (B.1) are general forms of the bias mean and the bias variance, respectively. We refer to $\mathsf{b}_{\mathcal{G}}(S)$ as the bias mean because $\mathsf{b}_{\mathcal{G}}(S)$ is the precise bias of the estimator that regress $Y$ on $X_{S}$ when $S^{\star}\subseteq S$ using all the data. This can be formally presented in the following proposition, which asserts that in the absence of our proposed regularizer, a vanilla least squares estimator will not consistently estimate $g^{\star}$.
Proposition B.1 (Inconsistency of Least Squares Estimator).
Let $S$ be an index set such that $S^{\star}\subseteq S\subseteq[d]$. Assume 4.1, 4.3, B.1B.2 hold, and $\mathsf{b}_{\mathcal{G}}(S)>0$. Suppose further that $U\delta_{n,\log n}+\inf_{g\in\mathcal{G}_{S}}\|g-\Pi_{\overline{\mathcal{G}_{S}}}(\bar{m}^{(S)})\|_{2}=o(1)$, where $U$ and $\delta_{n,t}$ are two constants defined in Theorem B.2 below. Then the estimator $\widehat{g}_{\mathtt{R}}$ that minimizes (3.6) in $\mathcal{G}_{S}$ satisfies, for large enough $n$,
$0.99\leq\frac{\|\widehat{g}_{\mathtt{R}}-g^{\star}\|_{2}^{2}}{\mathsf{b}_{\mathcal{G}}(S)}\leq 1.01$
with probability at least $1-\{C_{y}(\sigma_{y}+1)+1\}n^{-100}$.
On the other hand, our proposed FAIR estimator will not converge to the biased solution under the condition “heterogeneity”. The condition “heterogeneity” is an abstraction of the “identification” condition in previous subsections, for example, 4.5 for FAIR-NN.
The last condition “nondegenerate covariate” ensures that the target regression function $g^{\star}$ cannot be exactly fitted by any function $g$ whose dependent variable set $S_{g}$ does not cover $S^{\star}$. It reduces to be “non-collinearity” when $\mathcal{G}$ is linear.
Now we are ready to state the main result regarding the statistical rate of convergence of our estimator $\widehat{g}$ to $g^{\star}$, that is,
$\|\widehat{g}-g^{\star}\|_{2}=\left\{\int(\widehat{g}-g^{\star})^{2}\bar{\mu}_{x}(dx)\right\}^{1/2}.$
Theorem B.2 (Main Result for the FAIR Estimator with $\ell_{2}$ Loss).
Assume 4.1, 4.3, and B.1B.2 hold. Define the critical threshold
$\gamma^{\star}:=\sup_{S\subseteq[d]:\mathsf{b}_{\mathcal{G}}(S)>0}\frac{\mathsf{b}_{\mathcal{G}}(S)}{\bar{\mathsf{d}}_{\mathcal{G},\mathcal{F}}(S)}.$
There exists some universal constant $C$ such that, for any $\gamma\geq 8\gamma^{\star}$, the following holds:
(1) General $L_{2}$ error rate. Let $t>0$ be arbitrary. Define general approximation errors with respect to the function class $\mathcal{G}$ and $\mathcal{F}$ as
$\delta_{\mathtt{a},\mathcal{G}}=\inf_{g\in\mathcal{G}_{S^{\star}}}\|g-g^{\star}\|_{2}~{}~{}~{}~{}\text{and}~{}~{}~{}~{}\delta_{\mathtt{a},\mathcal{F},\mathcal{G}}(S)=\sqrt{\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\sup_{g\in\mathcal{G}:S_{g}=S}\inf_{f\in{\mathcal{F}_{S_{g}}}}\|\Pi_{\overline{\mathcal{F}_{S}}}^{(e)}(m^{(e,S)})-g-f\|_{2,e}^{2}},$
and the stochastic error as $\delta_{n,t}=\delta_{n}+\{(\log(nB|\mathcal{E}|)+t+1)/n\}^{1/2}$, where $\delta_{n}$ is the quantity in B.1. Let $U=B(B+\sigma_{y}\sqrt{\log(n|\mathcal{E}|)})$, then
$\|\widehat{g}-g^{\star}\|_{2}\leq C(1+\gamma)\left(U\delta_{n,t}+\delta_{\mathtt{a},\mathcal{G}}+\delta_{\mathtt{a},\mathcal{F},\mathcal{G}}(S_{\widehat{g}})+\delta_{\mathtt{a},\mathcal{F},\mathcal{G}}(S^{\star})+\delta_{\mathtt{opt}}\right).$ (B.2)
with probability at least $\mathfrak{p}=1-6e^{-t}-2C_{y}(\sigma_{y}+1)n^{-100}$.
(2) Faster $L_{2}$ error rate. Moreover, if
$\delta_{\mathtt{opt}}^{2}+\sup_{S\subseteq[d]}\delta^{2}_{\mathtt{a},\mathcal{F},\mathcal{G}}(S)+\delta^{2}_{\mathtt{a},\mathcal{G}}+UB\delta_{n,t}\leq\left\{1\land\frac{s_{\min}}{\gamma+1}\land\left(\frac{\gamma}{\gamma+1}\inf_{S:\bar{\mathsf{d}}_{\mathcal{G},\mathcal{F}}(S)>0}\bar{\mathsf{d}}_{\mathcal{G},\mathcal{F}}(S)\right)\right\}/C$ (B.3)
then the following holds, with probability at least $\mathfrak{p}$,
$\|\widehat{g}-g^{\star}\|_{2}\leq C\left(U\delta_{n,t}+\delta_{\mathtt{a},\mathcal{G}}+\delta_{\mathtt{a},\mathcal{F},\mathcal{G}}^{\star}+\delta_{\mathtt{opt}}\right),$ (B.4)
where $\delta_{\mathtt{a},\mathcal{F},\mathcal{G}}^{\star}=\{\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\sup_{g\in\mathcal{G}}\inf_{f\in\mathcal{F}_{S_{g}}}\|g^{\star}-g-f\|_{2,e}^{2}\}^{1/2}$.
Theorem B.2 generalizes Theorem 4.4 in Fan et al., (2023) to a broad spectrum of $(\mathcal{G},\mathcal{F})$ configurations. After specifying the function class $(\mathcal{G},\mathcal{F})$, one can further derive the corresponding identification condition by calculating $(\mathsf{b}_{\mathcal{G}}(S),\bar{\mathsf{d}}_{\mathcal{G},\mathcal{F}}(S))$ and establish a high probability bound on the $L_{2}$ error by substituting approximation errors $(\delta_{\mathtt{a},\mathcal{G}},\delta_{\mathtt{a},\mathcal{F},\mathcal{G}}(S),\delta_{\mathtt{a},\mathcal{F},\mathcal{G}}^{\star})$ and stochastic error $\delta_{n}$ for the function class $(\mathcal{G},\mathcal{F})$. In particular, when $\mathcal{G}$ and $\mathcal{F}$ are restricted to the linear function class, they not only match but also significantly improve the result in Fan et al., (2023); see Section B.6. All the results in Table 2 are direct corollaries of our abstract result Theorem B.2.
It is required that $\gamma$ should be greater than a constant-level critical threshold $8\gamma^{\star}$ for consistent estimation of $g^{\star}$. Theorem B.2 further establishes a crude instant-dependent and oracle-type error bound (B.2) that holds for arbitrary $n\geq 2$ and scales linearly with $\gamma$. Furthermore, when the stochastic error and approximation errors all go to 00 as $n$ increases and $n$ is large enough such that (B.3) holds, we have (B.4), which improves the $L_{2}$ error bound (B.2) in two aspects – the error bound is no longer dependent on either $\gamma$ or other $m^{(e,S)}$ with $S\neq S^{\star}$. The quantities in the RHS of (B.3) can be interpreted as the smaller of (1) the signal of true important variables and (2) the signal of heterogeneity. When one of these signals is weak, one can expect to demand more data to differentiate whether it is signal or noise.
One important ingredient in the FAIR estimator is the choice of regularization hyper-parameter $\gamma$ that promotes the invariance. Theorem B.2 offers some insights on choosing $\gamma$. Firstly, $\gamma\geq C\gamma^{\star}$ is required such that it will correctly identify $g^{\star}$ from a population-level perspective. Second, it will influence the $L_{2}$ error rate when $n$ is not large enough such that (B.3) does not hold. Furthermore, the final $L_{2}$ error rate (B.4) when $n$ is large enough is independent of $\gamma$. This indicates that the estimator’s performance is somewhat not very sensitive to the choice of hyper-parameter $\gamma$. In this case, one can adopt a slightly conservative large $\gamma$ to meet the population condition $\gamma\geq C\gamma^{\star}$.
B.2 Extension to the General Risk Loss $\ell$ under the Nonparametric Setting
Condition B.3 (Risk Loss).
Define $\mathcal{V}=[\inf_{g\in\mathcal{G}\cup\{g^{\star}\}}\sup_{L}\{g(X)\geq L,\bar{\mu}_{x}\text{-}a.s.\},\sup_{g\in\mathcal{G}\cup\{g^{\star}\}}\inf_{U}\{g(X)\leq U,\bar{\mu}_{x}\text{-}a.s.\}]$ be the value that $g(X)$ takes, and $\mathcal{Y}=[\sup_{l}\{Y\geq l,\bar{\mu}_{x}\text{-}a.s.\},\inf_{u}\{Y\leq u,\bar{\mu}_{x}\text{-}a.s.\}]$ be the value that $Y$ takes. The loss $\ell(\cdot,\cdot)$ satisfies
  • (1)
    $\ell(y,v)<\infty$ for any $y\in\mathcal{Y}$ and $v\in\mathcal{V}$ and twice continuously differentiable in $\mathcal{Y}\times\mathcal{V}$. $\frac{\partial\ell(y,v)}{\partial v}=(v-y)\psi(v)$ for some continuously differentiable $\psi(v):\mathbb{R}\to\mathbb{R}$.
  • (2)
    There exists some universal constant $\zeta\geq 1$ such that
    $|\psi(v)|\leq\zeta\qquad\text{and}\qquad{\zeta^{-1}}\leq\frac{\partial^{2}\ell}{\partial v^{2}}(Y,v)\leq{\zeta}\qquad\forall v\in\mathcal{V}\text{ and }\bar{\mu}\text{-}a.s.~{}.$
The assumptions on risk loss in B.3 is standard: (1) ensures that $\ell$ is well-defined on optimal solutions and linear combination of them, (2) requires that the population-level global minima is conditional mean, (3) guarantees that the loss function is strongly convex and smooth in the domain, and satisfies $|\ell(y,v)-\ell(y,v^{\prime})|\leq\zeta|y-\widetilde{v}||v-v^{\prime}|$ for some universal constant $\zeta$, which slightly relaxes the Lipschitz condition in Farrell et al., (2021) and Foster & Syrgkanis, (2019).
We now state the invariance and identification condition when the general risk loss is adopted.
Condition B.4 (Invariance and Identification for General Risk Loss).
Suppose the following holds
  • 1.
    (Invariance) There exists some index set $S^{\star}\subseteq[d]$ such that
    $\forall e\in\mathcal{E}\qquad m^{(e,S^{\star})}=\bar{m}^{(S^{\star})}=:m^{\star}$
  • 2.
    (Heterogeneity) For each $S\subseteq[d]$, if $\mathsf{b}(S)>0$, then $\bar{\mathsf{d}}(S)>0$, where
    $\mathsf{b}(S):=\|\bar{m}^{(S\cup S^{\star})}-m^{\star}\|_{2}^{2},\qquad\bar{\mathsf{d}}(S):=\frac{1}{|\mathcal{E}|}\sum_{e=1}^{m}\|m^{(e,S)}-\bar{m}^{(S)}\|_{2,e}^{2}.$ (B.5)
  • 3.
    (Nondegenerate Covariate) For any $S\subseteq[d]$ such that $S^{\star}\setminus S\neq\emptyset$, we have $\inf_{g\in\Theta_{S}}\|g-m^{\star}\|_{2}^{2}\geq s_{\min}$ for some constant $s_{\min}>0$.
We are now ready to state the main result in this case.
Theorem B.3 (Main Result for the FAIR Estimator with General Risk Loss).
Assume 4.1,4.3, B.1, and B.3B.4 hold. Define the critical threshold
$\gamma^{\star}:=\sup_{S\subseteq[d]:\mathsf{b}(S)>0}\frac{\mathsf{b}(S)}{\bar{\mathsf{d}}(S)}.$
There exists some universal constant $C$ such that, for any $\gamma\geq 8\zeta^{2}\gamma^{\star}$, the following holds:
(1) General $L_{2}$ error rate. Let $t>0$ be arbitrary. Define general approximation errors with respect to the function class $\mathcal{G}$ and $\mathcal{F}$ as
$\delta_{\mathtt{a},\mathcal{G}}=\inf_{g\in\mathcal{G}_{S^{\star}}}\|g-m^{\star}\|_{2}~{}~{}~{}~{}\text{and}~{}~{}~{}~{}\delta_{\mathtt{a},\mathcal{F},\mathcal{G}}(S)=\sqrt{\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\sup_{g\in\mathcal{G}:S_{g}=S}\inf_{f\in{\mathcal{F}_{S_{g}}}}\|m^{(e,S)}-g-f\|_{2,e}^{2}},$
and the stochastic error as $\delta_{n,t}=\delta_{n}+\{(\log(nB|\mathcal{E}|)+t+1)/n\}^{1/2}$, where $\delta_{n}$ is the quantity in B.1. Let $U=B(B+\sigma_{y}\sqrt{\log(n|\mathcal{E}|)})$, then
$\|\widehat{g}-m^{\star}\|_{2}\lor\|\widehat{g}-m^{\star}\|_{n}\leq C(\zeta+\gamma)\zeta\left(U\delta_{n,t}+\delta_{\mathtt{a},\mathcal{G}}+\delta_{\mathtt{a},\mathcal{F},\mathcal{G}}(S_{\widehat{g}})+\delta_{\mathtt{a},\mathcal{F},\mathcal{G}}(S^{\star})+\delta_{\mathtt{opt}}\right).$ (B.6)
with probability at least $\mathfrak{p}=1-6e^{-t}-2C_{y}(\sigma_{y}+1)n^{-100}$.
(2) Faster $L_{2}$ error rate. Moreover, if
$\delta_{\mathtt{opt}}^{2}+\sup_{S\subseteq[d]}\delta^{2}_{\mathtt{a},\mathcal{F},\mathcal{G}}(S)+\delta^{2}_{\mathtt{a},\mathcal{G}}+UB\delta_{n,t}\leq\left\{1\land\frac{s_{\min}}{(\gamma+\zeta)\zeta}\land\left(\frac{\gamma}{\gamma+\zeta}\inf_{S:\bar{\mathsf{d}}_{\mathcal{G},\mathcal{F}}(S)>0}\bar{\mathsf{d}}_{\mathcal{G},\mathcal{F}}(S)\right)\right\}/C$ (B.7)
then the following holds, with probability at least $\mathfrak{p}$,
$\|\widehat{g}-m^{\star}\|_{2}\lor\|\widehat{g}-m^{\star}\|_{n}\leq C\zeta^{2}\left(U\delta_{n,t}+\delta_{\mathtt{a},\mathcal{G}}+\delta_{\mathtt{a},\mathcal{F},\mathcal{G}}^{\star}+\delta_{\mathtt{opt}}\right),$ (B.8)
where $\delta_{\mathtt{a},\mathcal{F},\mathcal{G}}^{\star}=\{\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\sup_{g\in\mathcal{G}}\inf_{f\in\mathcal{F}_{S_{g}}}\|m^{\star}-g-f\|_{2,e}^{2}\}^{1/2}$.
B.3 Key Ideas and Proof Sketch of Theorem B.2
We first introduce some additional notations. Let
$\mathsf{A}^{(e)}(g,f^{(e)})$ $=\mathbb{E}\left[\{Y^{(e)}-g(X^{(e)})\}f^{(e)}(X^{(e)})-\frac{1}{2}\{f^{(e)}(X^{(e)})\}^{2}\right]$
$\widehat{\mathsf{A}}^{(e)}(g,f^{(e)})$ $=\frac{1}{n}\sum_{i=1}^{n}\{Y_{i}^{(e)}-g(X_{i}^{(e)})\}f^{(e)}(X_{i}^{(e)})-\frac{1}{2}\{f^{(e)}(X_{i}^{(e)})\}^{2}.$
Define the population-level pooled risk and FAIR estimator loss as
$\mathsf{R}(g)=\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\mathbb{E}\left[\frac{1}{2}|Y^{(e)}-g(X^{(e)})|^{2}\right]\qquad\text{and}\qquad\mathsf{Q}_{\gamma}(g,f^{\mathcal{E}})=\mathsf{R}(g)+\gamma\mathsf{J}(g,f^{\mathcal{E}})$
We will use the following theorem establishing approximate strong convexity with respect to $g^{\star}$.
Theorem B.4.
Assume B.2 hold, $\ell(y,v)=\frac{1}{2}(y-v)^{2}$. Let $\delta\in(0,1)$ be arbitrary. Then the following holds, for any $\gamma\geq 4\delta^{-1}\gamma^{\star}$,
$\mathsf{Q}_{\gamma}(g,f^{\mathcal{E}})-\mathsf{Q}_{\gamma}(\widetilde{g},\widetilde{f}^{\mathcal{E}})\geq$ $\frac{1-\delta}{2}\|g-\widetilde{g}\|_{2}^{2}+\frac{\gamma}{4}\bar{\mathsf{d}}_{\mathcal{G},\mathcal{F}}(S)+\frac{\gamma}{2}\|g-\Pi_{\overline{\mathcal{G}_{S}}}(\bar{m}^{(S)})\|_{2}^{2}$
$~{}~{}~{}~{}~{}~{}-\frac{\gamma}{2|\mathcal{E}|}\sum_{e\in\mathcal{E}}\|f^{(e)}-\{\Pi_{\overline{\mathcal{F}_{S}}}^{(e)}(m^{(e,S)})-g\}\|_{2,e}^{2}-(\delta^{-1}+\gamma/2)\|\widetilde{g}-g^{\star}\|_{2}^{2}$
for any $g\in\mathcal{G}$, $\widetilde{g}\in\mathcal{G}_{S^{\star}}$ and $S_{\widetilde{g}}=S^{\star}$, $f^{\mathcal{E}}\in\{\overline{\mathcal{F}_{S_{g}}}\}^{|\mathcal{E}|}$, and $\widetilde{f}^{\mathcal{E}}\in\{\overline{\mathcal{F}_{S^{\star}}}\}^{|\mathcal{E}|}$.
Recall our definition of
$\delta_{n,t}=\delta_{n}+\sqrt{\frac{t+\log(nB|\mathcal{E}|)+1}{n}}\qquad\text{and}\qquad U=B(B+\sigma\sqrt{\log(n|\mathcal{E}|)})$
The first proposition establishes instance-dependent error bounds on
$\Delta_{\mathsf{R}}(g,\widetilde{g}):=\{\widehat{\mathsf{R}}(g)-\widehat{\mathsf{R}}(\widetilde{g})\}-\{{\mathsf{R}}(g)-\mathsf{R}(\widetilde{g})\},$
and is standard in nonparametric regression literature.
Proposition B.5 (Instance-dependent error bounds for pooled risk).
Suppose 4.1,4.3, B.1 hold. There exists some universal constant $C$ such that for any $\eta>0$ and $t>0$, the following event
$\forall g,\widetilde{g}\in\mathcal{G},~{}~{}|\Delta_{\mathsf{R}}(g,\widetilde{g})|\leq CU\left\{\delta_{n,t}^{2}+\delta_{n,t}\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\|g-\widetilde{g}\|_{2,e}\right\}$
occurs with probability at least $1-3e^{-t}-C_{y}(\sigma_{y}+1)n^{-100}$.
The analysis of the focused adversarial invariance regularizer is more involved. The next proposition establishes the instance-dependent error bound for the regularizer. We define
$\Delta_{\mathsf{A}}^{(e)}(g,\widetilde{g},f^{(e)},\widetilde{f}^{(e)})=\mathsf{A}^{(e)}(f,g^{(e)})-\mathsf{A}^{(e)}(\widetilde{f},\widetilde{g}^{(e)})-\left\{\widehat{\mathsf{A}}^{(e)}(f,g^{(e)})-\widehat{\mathsf{A}}^{(e)}(\widetilde{f},\widetilde{g}^{(e)})\right\}$
and
$\mathcal{M}(\mathcal{G},\mathcal{F})=\left\{(g,\widetilde{g},f,\widetilde{f}):g,\widetilde{g}\in\mathcal{G}~{}\text{and}~{}f\in\mathcal{F}_{S_{g}},\widetilde{f}\in\mathcal{F}_{S_{\widetilde{g}}}\right\}.$
Proposition B.6 (Instance-dependent error bounds for regularizer).
Suppose 4.1, 4.3, B.1 hold. There exists some universal constant $C$ such that for any $t>0$, the following event
$\forall e\in\mathcal{E},$ $~{}\forall(g,\widetilde{g},f^{(e)},\widetilde{f}^{(e)})\in\mathcal{M}(\mathcal{G},\mathcal{F}),$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}|\Delta_{\mathsf{A}}^{(e)}(g,\widetilde{g},f^{(e)},\widetilde{f}^{(e)})|\leq CU\left(\delta_{n,t}\left(\|\widetilde{g}-g\|_{2,e}+\|\widetilde{g}+\widetilde{f}^{(e)}-g-f^{(e)}\|_{2,e}\right)+\delta_{n,t}^{2}\right)$
occurs with probability at least $1-3e^{-t}-C_{y}(\sigma_{y}+1)n^{-100}$.
We first utilize Proposition B.6 in a way that $g$ and $\widetilde{g}$ are the same. In this case, the optimization problem of $\max$$-\mathcal{F}$ in one single environment $e\in\mathcal{E}$ for fixed $g\in\mathcal{G}$ is similar to least squares regression that fits the target regression function
$\Pi_{\overline{\mathcal{F}_{S}}}^{(e)}(m^{(e,S)})-g.$
Thus one can establish high probability error bounds on the $\|\cdot\|_{2,e}$ norm between the empirical loss maximizer $\widehat{f}_{g}^{(e)}$ and the above target function in terms of statistical error $\delta_{n,t}$ and approximation error rate $\delta_{\mathtt{a},\mathcal{F},\mathcal{G}}(e,S_{g})$, defined as
$\delta_{\mathtt{a},\mathcal{F},\mathcal{G}}(e,S):=\sup_{g\in\mathcal{G}:S_{g}=S}\inf_{f\in\mathcal{F}_{S}}\|\Pi_{\overline{\mathcal{F}_{S}}}^{(e)}(m^{(e,S)})-g-f\|_{2,e}$
We formally present the above intuition in the following instance-dependent error bound in Proposition B.7 in a way that the optimization gap term is maintained in the error bound.
Proposition B.7 (Instance-dependent characterization of approximately optimal discriminator).
Let $20<\eta<1/$ 2 be arbitrary, under the event defined in Proposition B.6, the following holds,
$\forall e\in\mathcal{E},\forall g\in\mathcal{G},\forall f^{(e)}\in\mathcal{F}_{S_{g}},$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\|\Pi_{\overline{\mathcal{F}_{S}}}(m^{(e,S)})-g-f^{(e)}\|_{2,e}^{2}\leq\frac{2\eta^{-1}+2-4\eta}{1-2\eta}\delta^{2}_{\mathtt{a},\mathcal{F},\mathcal{G}}(e,S_{g})+\frac{2\eta^{-1}+4}{1-2\eta}C^{2}U^{2}\delta_{n,t}^{2}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\frac{4}{1-2\eta}\left\{\sup_{\breve{f}\in\mathcal{F}_{S_{g}}}\widehat{\mathsf{A}}^{(e)}(g,\breve{f})-\widehat{\mathsf{A}}^{(e)}(g,f^{(e)})\right\}$
where $C$ is the universal constant defined in Proposition B.6. Averaging over all the $e\in\mathcal{E}$, we obtain
$\forall g\in\mathcal{G},~{}\forall f^{\mathcal{E}}\in\{\mathcal{F}_{S_{g}}\}^{|\mathcal{E}|},$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\|\Pi_{\overline{\mathcal{F}_{S}}}(m^{(e,S)})-g-f^{(e)}\|_{2,e}^{2}\leq\frac{2\eta^{-1}+2-4\eta}{1-2\eta}\delta^{2}_{\mathtt{a},\mathcal{F},\mathcal{G}}(S_{g})+\frac{2\eta^{-1}+4}{1-2\eta}C^{2}U^{2}\delta_{n,t}^{2}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\gamma^{-1}\frac{4}{1-2\eta}\left\{\sup_{\breve{f}^{\mathcal{E}}\in\{\mathcal{F}_{S_{g}}\}^{|\mathcal{E}|}}\widehat{\mathsf{Q}}_{\gamma}(g,\breve{f})-\widehat{\mathsf{Q}}_{\gamma}(g,f^{\mathcal{E}})\right\}$
Now we are ready to prove Theorem B.2.
For the proof of (2) faster $L_{2}$ rate, we will divide the proof into two main steps as follows.
  1. 1.
    In the first step, we establish a variable selection property claim that when the Eq. B.3 holds, and the events defined in Proposition B.5 and B.6 occurs, then $\widehat{S}$ satisfies
    $\forall e\in\mathcal{E}\qquad\Pi_{\overline{\mathcal{F}_{\widehat{S}}}}^{(e)}(m^{(e,\widehat{S})})=g^{\star}$
    using proof by contradiction that any $g$ such that such that the above constrain is violated in $S_{g}$, will not be the approximate solution of the minimax optimization $\inf_{g}\sup_{f^{\mathcal{E}}}\widehat{\mathsf{Q}}_{\gamma}(g,f^{\mathcal{E}})$. This can be summarized as the following Proposition B.8.
  2. 2.
    In the second step, we proceed conditioned on the above claim and derive a sharp $L_{2}$ error bound. To derive a sharp error bound, we combine (1) the approximate strong convexity with respect to $g^{\star}$, i.e., Theorem B.4, (2) the instance-dependent error bound for $\mathsf{J}$ and $\mathsf{R}$, i.e., Proposition B.5 and B.6, and (3) the key fact that, if the claim in step 1 holds, then
    $\|\widetilde{g}+\widetilde{f}_{\widetilde{g}}^{(e)}-g-f_{g}^{(e)}\|_{2,e}$ $\leq\|\widetilde{g}+\widetilde{f}_{\widetilde{g}}^{(e)}-g^{\star}+\Pi_{\overline{\mathcal{F}_{S_{g}}}}^{(e)}(m^{(e,S_{g})})-g-f_{g}^{(e)}\|_{2,e}$
    $\lesssim\|\widetilde{g}+\widetilde{f}_{\widetilde{g}}^{(e)}-g^{\star}\|_{2,e}+\|g^{\star}-g-f_{g}^{(e)}\|_{2,e}$
    $\lesssim\delta_{n,t}+\delta_{\mathtt{a},\mathcal{F},\mathcal{G}}^{\star}.$
The proof of (1) is similar to the second step in the proof of (2), but now we no longer have $g^{\star}=\Pi_{\overline{\mathcal{F}_{S_{g}}}}^{(e)}(m^{(e,S_{g})})$. The key challenge here is to establish an upper bound on $\|g^{\star}-\Pi_{\overline{\mathcal{F}_{S_{g}}}}^{(e)}(m^{(e,S_{g})})\|_{2,e}$ without imposing other population-level condition like Condition 7 in an early version of Fan et al., (2023). Instead, we will use the following instance-dependent bound, that
$\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\|g^{\star}-\Pi_{\overline{\mathcal{F}_{S_{g}}}}^{(e)}(m^{(e,S_{g})})\|_{2,e}^{2}\leq C\left((1+\gamma^{\star})\bar{\mathsf{d}}_{\mathcal{G},\mathcal{F}}(S_{g})+\|g-g^{\star}\|_{2}^{2}\right)$
Such a bound is a population-level instance-dependent bound in that both the R.H.S. and L.H.S. are dependent on the function $g$.
Proposition B.8.
Under the event defined in Proposition B.6 and B.5, we have the event
$\mathcal{A}_{+}:=\left\{\forall e\in\mathcal{E}\qquad\Pi_{\overline{\mathcal{F}_{\widehat{S}}}}^{(e)}(m^{(e,\widehat{S})})=g^{\star}\qquad\text{for}~{}~{}\widehat{S}=S_{\widehat{g}}\right\}$ (B.9)
occurs if the condition (B.3) with some large universal constant $C$ holds.
B.4 Applications of Theorem B.2 and Connection to the Predecessors
We present some examples here, sorted by the potential approximation capability of the function class $(\mathcal{G},\mathcal{F})$.
Example 3 (Linear $\mathcal{G}$, Linear $\mathcal{F}$).
The simplest case is that $\mathcal{G}$ and $\mathcal{F}$ are all linear function classes, that
$\mathcal{G}=\mathcal{F}=\{h(x)=\beta^{\top}x:\beta\in\mathbb{R}^{d}\}:=\mathcal{H}_{\mathtt{lin}}(d).$
The objective takes on a form that closely resembles the EILLS objective proposed in Fan et al., (2023). To see this, the EILLS objective is expressed as $\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\widehat{\mathbb{E}}[|Y^{(e)}-g(X^{(e)})|^{2}]+\frac{\gamma}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\|\widehat{r}_{g}^{(e)}\|_{2}^{2}$ where $\widehat{r}_{g}^{(e)}=\widehat{\mathbb{E}}[\{Y^{(e)}-g(X^{(e)})\}X_{S_{g}}^{(e)}]$. If we take the supremum over all the $f^{(e)}\in\mathcal{F}_{S_{g}}$ with $e\in\mathcal{E}$, the objective in (3.5) transforms into
$\sup_{f^{\mathcal{E}}\in\{\mathcal{F}_{S_{g}}\}^{|\mathcal{E}|}}\widehat{\mathsf{Q}}_{\gamma}(g,f^{\mathcal{E}})=\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\widehat{\mathbb{E}}[|Y^{(e)}-g(X^{(e)})|^{2}]+\frac{\gamma}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}(\widehat{r}_{g}^{(e)})^{\top}\{\widehat{\mathbb{E}}[X_{S}^{(e)}(X_{S}^{(e)})^{\top}]\}^{-1}(\widehat{r}_{g}^{(e)}).$
It slightly stabilizes the EILLS objective in that the regularizer has a matched moment index compared with the pooled least squares loss; see a detailed explanation and theoretical justification in Section B.6.
Example 4 (Linear $\mathcal{G}$, Augmented Linear $\mathcal{F}$).
Consider the case where $\mathcal{F}$ is potentially larger than $\mathcal{G}$, that is, $\mathcal{G}=\mathcal{H}_{\mathtt{lin}}(d)$ and =$\mathcal{F}=\{f(x)=\beta^{\top}x+\beta_{\phi}^{\top}\bar{\phi}(x):\beta,\beta_{\phi}\in\mathbb{R}^{d}\}:=\mathcal{H}_{\mathtt{alin}}(d,\phi)$, where $\bar{\phi}(x)=(\phi(x_{1}),\ldots,\phi(x_{d}))$ applies a transformation function $\phi:\mathbb{R}\to\mathbb{R}$ to each entry of the vector $x$.
The proposed estimator utilizes both the heterogeneity among different environments and the strong prior knowledge that the true regression function admits linear form. It bridges the EILLS estimator in Fan et al., (2023) and the Focused GMM estimator in Fan & Liao, (2014) when the instrumental variables are $[X_{S},\bar{\phi}(X_{S})]$ and reduces to an improved version of the latter when $|\mathcal{E}|=1$.
Example 5 (Linear $\mathcal{G}$, Neural Network $\mathcal{F}$).
We consider a more algorithmic version of Example 4 that uses neural networks to automatically learn the transformation function, that is, $\mathcal{G}=\mathcal{H}_{\mathtt{lin}}(d)$ and $\mathcal{F}=\mathcal{H}_{\mathtt{nn}}(d,L_{f},N_{f},B_{f})$ with neural network architecture hyper-parameters of $(L_{f},N_{f},B_{f})$.
The above three estimators focus on linear $\mathcal{G}$, the simplest structural function class. We now consider a more complicated structural function class when we know the invariant association admits additive form.
Example 6 (Additive Neural Network $\mathcal{G}$, Neural Network $\mathcal{F}$).
We let $\mathcal{G}=\mathcal{H}_{\mathtt{ann}}(d,L_{g},N_{g},B_{g}):=\{g(x)=\mathrm{Tc}_{B_{g}}(\sum_{j=1}^{d}g_{j}(x_{j})):g_{j}\in\mathcal{H}_{\mathtt{nn}}(1,L_{g},N_{g},\infty)\}$ and $\mathcal{F}=\mathcal{H}_{\mathtt{nn}}(d,L_{f},N_{f},B_{f})$. Here $(L_{g},N_{g},B_{g})$ and $(L_{f},N_{f},B_{f})$ are all neural network architecture hyper-parameters.
Finally, we present the most algorithmic estimator, the FAIR-NN estimator, in which both $\mathcal{G}$ and $\mathcal{F}$ are realized by fully-connected neural networks with no additional imposed structures.
Example 7 (Neural Network $\mathcal{G}$, Neural Network $\mathcal{F}$).
We let $\mathcal{G}=\mathcal{H}_{\mathtt{nn}}(d,L_{g},N_{g},B_{g})$ and $\mathcal{F}=\mathcal{H}_{\mathtt{nn}}(d,L_{f},N_{f},B_{f})$ with neural network architecture hyper-parameters $(L_{g},N_{g},B_{g})$ and $(L_{f},N_{f},B_{f})$.
[…]
Example 3 $\mathcal{H}_{\mathtt{lin}}(d)$ $\mathcal{H}_{\mathtt{lin}}(d)$ $\mathcal{G}\asymp\mathcal{F}$ FAIR-Linear Theorem B.10
Example 7 $\mathcal{H}_{\mathtt{nn}}(d,L_{g},N_{g},B_{g})$ $\mathcal{H}_{\mathtt{nn}}(d,L_{f},N_{f},B_{f})$ $\mathcal{G}\asymp\mathcal{F}$ FAIR-NN Theorem 4.1
Example 4 $\mathcal{H}_{\mathtt{lin}}(d)$ $\mathcal{H}_{\mathtt{alin}}(d,\phi)$ $\mathcal{G}\ll\mathcal{F}$ FAIR-AugLinear Theorem B.11
Example 5 $\mathcal{H}_{\mathtt{lin}}(d)$ $\mathcal{H}_{\mathtt{nn}}(d,L_{f},N_{f},B_{f})$ $\mathcal{G}\ll\mathcal{F}$ FAIR-NNLinear Theorem B.12
Example 6 $\mathcal{H}_{\mathtt{ann}}(d,L_{g},N_{g},B_{g})$ $\mathcal{H}_{\mathtt{nn}}(d,L_{f},N_{f},B_{f})$ $\mathcal{G}\ll\mathcal{F}$ FAIR-ANN Theorem B.9
Table 2: A Glimpse of Estimators
Our framework requires $\mathcal{G}\subseteq\mathcal{F}$. We can divide the above estimators into two main categories that (1) $\mathcal{G}$ has roughly the same representation power as $\mathcal{F}$, denoted as $\mathcal{G}\asymp\mathcal{F}$, and (2) $\mathcal{F}$ has at least as good representation power as $\mathcal{G}$, denoted as $\mathcal{G}\ll\mathcal{F}$. For the former, our framework uses only heterogeneity among different environments to identify the invariant association. For the latter, our framework utilizes both the heterogeneity and strong prior structural assumption that the invariant association cannot be significantly better approximated by $\mathcal{F}$ than by $\mathcal{G}$ to jointly identify the invariant association. We summarize the proposed estimators above and divide them into these two categories in Table 2.
B.5 FAIR-ANN: Bridging Invariance and Additional Structural Knowledge
We next consider the estimator that utilizes both heterogeneity and the strong structural assumption that the invariant association $m^{\star}$ admits additive form to identify $m^{\star}$, which can be summarized as the following assumption.
Condition B.5 (Invariance and Nondegenerate Covariate for FAIR-ANN).
There exists some set $S^{\star}$ and $m^{\star}:\mathbb{R}^{|S^{\star}|}\to\mathbb{R}$ such that $m^{(e,S^{\star})}(x)\equiv m^{\star}(x_{S^{\star}})=\sum_{j\in S^{\star}}m^{\star}_{j}(x_{j})$ for any $e\in\mathcal{E}$. Moreover, for any $S\subseteq[d]$ with $S^{\star}\setminus S\neq\emptyset$, $\inf_{m\in\Theta_{S}}\|m-m^{\star}\|_{2}^{2}\geq s_{\min}>0$.
Condition B.6.
There exists some constant $C_{a}$ such that
$\left\|\sum_{j=1}^{d}m_{j}(x_{j})\right\|_{2}^{2}\geq C_{a}^{-1}\sum_{j=1}^{d}\|m_{j}(x_{j})\|_{2}^{2}\qquad\forall(m_{1},\ldots,m_{d})\in\prod_{j=1}^{d}\Theta_{\{j\}}~{}\text{with}~{}\int m_{j}(x_{j})\bar{\mu}_{x}(dx)\equiv 0.$
The above condition is referred to as the nonparametric version of the restricted strong convexity condition, which is widely used in the theoretical analysis for nonparametric high-dimension additive models (Van de Geer,, 2008; Raskutti et al.,, 2012; Yuan & Zhou,, 2016). This condition is imposed to let $\prod_{j\in S}\Theta_{\{j\}}$ be a closed subspace of $\Theta_{S}$, where we can define
$A_{S}(h)=\mathop{\mathrm{argmin}}_{u\in\prod_{j\in S}\Theta_{\{j\}}}\|h-u\|_{2},$
which finds a unique additive function dependent on $x_{S}$ that fits $h$ best in $\|\cdot\|_{2}$ norm.
Condition B.7 (Identification for FAIR-ANN).
For any $S\subseteq[d]$ such that $\bar{\mu}(\{m^{\star}\neq A_{S\cup S^{\star}}(\bar{m}^{(S\cup S^{\star})})\})>0$, either of the two holds: (1) there exists some $e,e^{\prime}\in\mathcal{E}$ such that $(\mu^{(e)}\land\mu^{(e^{\prime})})(\{m^{(e,S)}\neq m^{(e^{\prime},S)}\})>0$, or (2) $\bar{\mu}(\{\bar{m}^{(S)}\neq A_{S}(\bar{m}^{(S)})\})>0$.
With network hyper-parameter $N,L$, we realize the $\mathcal{G}$ and $\mathcal{F}$ as
$\mathcal{G}=\mathcal{H}_{\mathtt{ann}}(d,L,N,b_{m})\qquad\text{and}\qquad\mathcal{F}=\mathcal{H}_{\mathtt{nn}}(d,L+2,2dN,2b_{m}).$ (B.10)
Similarly to the choice of for FAIR-NN (4.1), the choice of $\mathcal{F}$ is to ensure $\mathcal{G}-\mathcal{G}\subseteq\mathcal{F}$.
Theorem B.9 (Optimal Rate for FAIR-ANN Least Squares Estimator).
Assume 4.1,4.3, 4.4, and B.5B.7 hold. Assume further that all the conditional moments $\{m^{(e,S)}\}_{e\in\mathcal{E},S\subseteq[d]}$ are $(\beta^{\prime},C^{\prime})$-smooth for some $\beta^{\prime}>0$ and $C^{\prime}>0$, and $\delta_{\mathtt{opt}}=o(1)$. Consider the FAIR-ANN estimator that solves (3.8) with $\ell(y,v)=\frac{1}{2}(y-v)^{2}$ using $\gamma\geq 8\gamma^{\star}_{\mathtt{AN}}$ with
$\gamma_{\mathtt{AN}}^{\star}:=\sup_{S\subseteq[d]:\bar{\mu}(\{m^{\star}\neq A_{S\cup S^{\star}}(\bar{m}^{(S\cup S^{\star})})\})>0}\frac{\|m^{\star}-A_{S\cup S^{\star}}(\bar{m}^{(S\cup S^{\star})}))\|_{2}^{2}}{\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\|m^{(e,S)}-A_{S}(\bar{m}^{(S)}))\|_{2,e}^{2}},$ (B.11)
and function class (B.10) with $L,N$ satisfying $LN\asymp\{n(\log n)^{8\beta^{\star}-3}\}^{\frac{1}{2(2\beta^{\star}+1)}}$ and $(\log n)/(N\land L)=o(1)$. Then, we have (1) $\gamma_{\mathtt{AN}}^{\star}\leq\gamma_{\mathtt{NN}}^{\star}$, and (2) for $n$ large enough, the following event occurs with probability at least $1-\widetilde{C}n^{-100}$
$\sup_{\begin{subarray}{c}m^{\star}=\sum_{j\in S^{\star}}m^{\star}_{j}(x_{j})~{}\text{with}~{}m_{j}^{\star}\in\mathcal{H}_{\mathtt{HS}}(1,\beta^{\star},C^{\star})\\ \|m^{\star}\|_{\infty}\leq b_{m}\end{subarray}}\|\widehat{g}-m^{\star}\|_{2}\leq\widetilde{C}\left\{\delta_{\mathtt{opt}}+\left(\frac{\log^{7}n}{n}\right)^{-\frac{\beta^{\star}}{2\beta^{\star}+1}}\right\},$ (B.12)
where $\widetilde{C}$ is a constant that depends on $(C_{1},d,\beta^{\star},C^{\star},\sigma_{y},C_{y},b_{x},b_{m})$ but independent of $\gamma,\delta_{\mathtt{opt}}$ and $n$.
The choice of $N,L$, and the convergence rate align with FAIR-NN with $\alpha^{\star}=\beta^{\star}$. Given the strong structural prior knowledge that the true regression function is additive, FAIR-ANN requires weaker identification condition B.7 and also smaller critical threshold of $\gamma$. In particular, B.7 requires that for any $S$ such that regressing $Y$ on $X_{S\cup S^{\star}}$ via additive models yields biased estimation, there should be either (1) a shift in conditional moments $m^{(e,S)}$ across different environments, or (2) one of the conditional moments $m^{(e,S)}$ is non-additive. This characteristic is called the “double identifiable” property since meeting either of these conditions can consistently estimate $m^{\star}$. Notably, the critical threshold $\gamma^{\star}_{\mathtt{AN}}$ can be smaller than that of the FAIR-NN estimator. A small $\gamma$ can be adopted if either the signal of violating the additive structure or the signal of heterogeneity is strong.
B.6 Theoretical Analysis for Linear $\mathcal{G}$
In this section, we apply our result in Theorem B.2 to the cases where the target regression function $g^{\star}$ is linear. As such, we use linear function class $\mathcal{H}_{\mathtt{lin}}(d)$ as our predictor function class $\mathcal{G}$. Our theorem suggests that enhancing the potential approximation ability of the discriminator function class $\mathcal{F}$ will result in (1) a stronger condition on invariance, and (2) a weaker identification condition and a reduced choice of critical threshold $\gamma^{\star}$.
B.6.1 Linear $\mathcal{F}$
We first consider the case where we use linear discriminator function class $\mathcal{F}=\mathcal{H}_{\mathtt{lin}}(d)$. We introduce some notations used in linear regression and state some standard regularity conditions used in linear regression and are also imposed in Fan et al., (2023).
Condition B.8.
Suppose the following holds:
  • (1)
    The data satisfies 4.1 with $|\mathcal{E}|\leq n^{C_{1}}$ for some constant $C_{1}$.
  • (2)
    The covariance matrix $\Sigma^{(e)}=\mathbb{E}[X^{(e)}(X^{(e)})^{\top}]\in\mathbb{R}^{d\times d}$ in each environment satisfies $\lambda(\Sigma^{(e)})\geq\kappa_{L}$ for some constant $\kappa_{L}>0$.
  • (3)
    Define the pooled covariance matrix ${\Sigma}:=|\mathcal{E}|^{-1}\sum_{e\in\mathcal{E}}\Sigma^{(e)}$. There exists some positive constant $C_{x},\sigma_{x}$ such that
    $\forall e\in\mathcal{E},~{}\forall v\in\mathbb{R}^{d}~{}\text{with}~{}\|v\|_{2}=1,~{}\forall t\in[0,\infty),\qquad\mathbb{P}\left(|v^{\top}({\Sigma})^{-1/2}X^{(e)}|\geq t\right)\leq C_{x}e^{-t^{2}/(2\sigma_{x}^{2})}$
  • (4)
    4.3 holds.
Under B.8 that the covariance matrices are all positive definite, we can define
$\beta^{(e,S)}=\mathop{\mathrm{argmin}}_{\beta\in\mathbb{R}^{d}:\beta_{S^{c}}=0}\mathbb{E}[|Y^{(e)}-\beta^{\top}X^{(e)}|^{2}]$
We can state the invariance and identification condition in this case.
Condition B.9 (Invariance in Linear $\mathcal{G}$ and Linear $\mathcal{F}$).
There exists some $S^{\star}\subseteq[d]$ and $\beta^{\star}\in\mathbb{R}^{d}$ with $\beta^{\star}_{(S^{\star})^{c}}=0$ and $\min_{j\in S^{\star}}|\beta_{j}^{\star}|=\beta_{\min}>0$ such that
$\forall e\in\mathcal{E}\qquad\beta^{(e,S)}=\beta^{\star}.$ (B.13)
Let $\varepsilon^{(e)}=Y^{(e)}-(\beta^{\star})^{\top}X^{(e)}$, the above invariance equality (B.13) is equivalent to that $X_{S^{\star}}$ are exogenous across all the environments, that is,
$\forall e\in\mathcal{E}\qquad\mathbb{E}[\varepsilon^{(e)}X_{S^{\star}}^{(e)}]=0$
Condition B.10 (Identification for Linear $\mathcal{G}$ and Linear $\mathcal{F}$).
For any $S\subseteq[d]$ with $\sum_{e\in\mathcal{E}}\mathbb{E}[X_{S}^{(e)}\varepsilon^{(e)}]\neq 0$, there exists $e,e^{\prime}\in\mathcal{E}$ such that $\beta^{(e,S)}\neq\beta^{(e^{\prime},S)}$.
We are ready to state the result using truncated linear function class with bounded $L_{2}$ norm, that is,
$\mathcal{H}_{\mathtt{lin}}(d,B_{1},B_{2})=\left\{f(x)=\mathrm{Tc}_{B_{2}}(\beta^{\top}x):\beta\in\mathbb{R}^{d},\|\Sigma^{1/2}\beta\|_{2}\leq B_{1}\right\}.$
Theorem B.10 (Linear $\mathcal{G}$ and Linear $\mathcal{F}$).
Suppose B.8B.10 hold, and we choose
$\mathcal{G}=\mathcal{H}_{\mathtt{lin}}(d,C_{2},C_{2}\sqrt{\log n})\qquad\text{and}\qquad\mathcal{F}=\mathcal{H}_{\mathtt{lin}}(d,2C_{2},2C_{2}\sqrt{\log n})$
with some constant $C_{2}\geq 2(\sigma_{x}\lor 1)\max_{e\in\mathcal{E},S\subseteq[d]}\|\Sigma^{1/2}\beta^{(e,S)}\|_{2}$. Then, there exists some constant $\widetilde{C}$ that only depends on $(C_{1},C_{2},\sigma_{x},C_{x},\sigma_{y},C_{y})$ such that the FAIR least squares estimator using the above function class and hyper-parameter $\gamma$ satisfying $\gamma\geq 8\gamma_{\mathtt{LL}}^{\star}=8\sup_{S:\mathsf{b}_{\mathtt{LL}}(S)>0}\mathsf{b}_{\mathtt{LL}}(S)/\bar{\mathsf{d}}_{\mathtt{LL}}(S)$, where
$\begin{split}\mathsf{b}_{\mathtt{LL}}(S)&=\bigg{\|}\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\mathbb{E}[X_{S\cup S^{\star}}^{(e)}\varepsilon^{(e)}]\bigg{\|}_{(\bar{\Sigma}_{S\cup S^{\star}})^{-1}}^{2}\leq(\kappa_{L})^{-1}\bigg{\|}\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\mathbb{E}[X_{S}^{(e)}\varepsilon^{(e)}]\bigg{\|}_{2}^{2},\\ \bar{\mathsf{d}}_{\mathtt{LL}}(S)&=\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\|\beta^{(e,S)}_{S}-\beta^{(S)}_{\dagger}\|_{\Sigma_{S}^{(e)}}^{2}\geq\kappa_{L}\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\|\beta^{(e,S)}-\bar{\beta}^{(S)}\|_{2}^{2}\end{split}$ (B.14)
with $\beta_{\dagger}^{(S)}=(\bar{\Sigma}_{S})^{-1}\{\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\mathbb{E}[X^{(e)}_{S}Y^{(e)}]\}$ and $\bar{\beta}^{(S)}=\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\beta^{(e,S)}$, satisfies, with probability at least $1-\widetilde{C}n^{-100}$,
$\forall n\geq 3\qquad\|{\Sigma}^{1/2}(\beta_{\widehat{g}}-\beta^{\star})\|_{2}\leq\widetilde{C}(1+\gamma)\sqrt{\frac{d\log^{5}(n)}{n}},$ (B.15)
for $\widehat{g}(x)=\mathrm{Tc}_{B}(\beta_{\widehat{g}}^{\top}x)$. Moreover, if $d=o((1+\gamma^{2})n/(\log^{6}n))$, then for large enough $n$, we further have
$\|{\Sigma}^{1/2}(\beta_{\widehat{g}}-\beta^{\star})\|_{2}\leq\widetilde{C}\sqrt{\frac{d\log^{5}(n)}{n}}$ (B.16)
Remark 5.
We present the results using truncated function classes, and there exist poly-$\log n$ factors in the non-asymptotic $L_{2}$ error bounds. These are for technical convenience such that we can directly apply our result Theorem B.2 which focuses on uniformly bounded function classes. Indeed, one can use a finer analysis and obtain the $\ell_{2}$ error bound
$\sqrt{\frac{d+\log n}{n}}$
using unbounded linear function class.
The obtained results in Theorem B.10 align with (up to $\log(n)$ factors) and offer significant enhancements over Theorem 2 & 3 from Fan et al., (2023). Firstly, the “invariance” condition gets relaxed, we only assume that the noise $\varepsilon^{(e)}$ and the true important variables $X_{S^{\star}}^{(e)}$ are uncorrelated rather than conditional independent across different environments. Meanwhile, the identification condition B.10 exactly matches that in Fan et al., (2023) (refer to Condition 5 therein), and the choice of critical threshold $\gamma^{\star}$ gets reduced as indicated by the inequality in (B.14) and given that $\kappa_{L}=O(1)$. Such an improvement can be attributed to the term $-\frac{1}{2}\{f^{(e)}\}^{2}$ in our minimax regularization that stabilizes the objective. To see this, consider $\beta$ with $\mathrm{supp}(\beta)=S^{\star}$, the population-level EILLS objective can be written as
$(\beta-\beta^{\star})^{\top}\Sigma(\beta-\beta^{\star})+\gamma\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}(\beta-\beta^{\star})_{S^{\star}}^{\top}(\Sigma^{(e)}_{S^{\star}})^{2}(\beta-\beta^{\star})_{S^{\star}},$
where a square of the covariance matrix appears in the regularizer. This does not match what it is in the empirical risk part and will make the objective less stable. Meanwhile, the population-level FAIR objective with sup-$f$ in this case is
$(1+\gamma)(\beta-\beta^{\star})^{\top}\Sigma(\beta-\beta^{\star}),$
which the problem of mismatched covariance matrix order disappears.
We’ve also refined the non-asymptotic $L_{2}$ error bounds. On the one hand, we can derive the error bound without further imposing stronger population-level conditions (Condition 7 required by Theorem 3 in Fan et al., (2023)). On the other, the faster $\ell_{2}$ error bound for sufficiently large $n$ remains independent of the hyper-parameter $\gamma$ we choose. These refinements result from our tighter characterization of the instance-dependent error bounds compared to the ones in Fan et al., (2023); see the discussion on technical novelties in Section B.3.
B.6.2 Augmented Linear $\mathcal{F}$
Here we consider the case where the discriminator function class $\mathcal{F}$ is potentially larger than the predictor function class $\mathcal{G}$. We introduce the following notations. We let $[x,y]$ be the concatenation of two vectors $x\in\mathbb{R}^{d_{1}}$ and $y\in\mathbb{R}^{d_{2}}$ as a $d_{1}+d_{2}$ dimensional vector. For each $S\subseteq[d]$, we define $\widetilde{X}_{S}^{(e)}=[X_{S}^{(e)},\bar{\phi}(X_{S}^{(e)})]\in\mathbb{R}^{2|S|}$, $\widetilde{\Sigma}^{(e)}_{S}=\mathbb{E}[\widetilde{X}_{S}^{(e)}(\widetilde{X}_{S}^{(e)})^{\top}]\in\mathbb{R}^{(2|S|)\times(2|S|)}$ and let $\widetilde{X}^{(e)}=\widetilde{X}^{(e)}_{[d]}$ and $\widetilde{\Sigma}^{(e)}=\widetilde{\Sigma}^{(e)}_{[d]}$. We impose additional regularity conditions due to the incorporation of basis function $\phi$.
Condition B.11.
There exists some constant $\widetilde{\kappa}_{L}>0$ such that $\lambda_{\min}(\widetilde{\Sigma}^{(e)})\geq\widetilde{\kappa}_{L}$ for any $e\in\mathcal{E}$. Moreover, define $\widetilde{\Sigma}:=|\mathcal{E}|^{-1}\sum_{e\in\mathcal{E}}\widetilde{\Sigma}^{(e)}$. There exists some positive constant $C_{\widetilde{x}},\sigma_{\widetilde{x}}$ such that
$\forall e\in\mathcal{E},~{}\forall v\in\mathbb{R}^{2d}~{}\text{with}~{}\|v\|_{2}=1,~{}\forall t\in[0,\infty),\qquad\mathbb{P}\left(|v^{\top}(\widetilde{\Sigma})^{-1/2}\widetilde{X}^{(e)}|\geq t\right)\leq C_{\widetilde{x}}e^{-t^{2}/(2\sigma_{\widetilde{x}}^{2})}$
Under B.11 such that the covariance matrix for $\widetilde{X}$ are positive definite, we can define
$\widetilde{\beta}^{(e,S)}=[\breve{\beta},\breve{\beta}^{\phi}]~{}~{}~{}~{}\text{with}~{}~{}~{}~{}(\breve{\beta},\breve{\beta}^{\phi})=\mathop{\mathrm{argmin}}_{({\beta},\beta^{\phi})\in(\mathbb{R}^{d})^{2},\beta_{S^{c}}=\beta_{S^{c}}^{\phi}=0}\mathbb{E}[|Y^{(e)}-\beta^{\top}X^{(e)}-(\beta^{\phi})^{\top}\bar{\phi}(X^{(e)})|^{2}],$
and $\widetilde{\beta}^{(e,S)}_{S}=[\breve{\beta}_{S},\breve{\beta}^{\phi}_{S}]$ be a $2|S|$-dimensional vector. The invariance and identification conditions in this case are as follows.
Condition B.12 (Invariance in Linear $\mathcal{G}$ and Augmented Linear $\mathcal{F}$).
There exists some $S^{\star}\subseteq[d]$ and $\beta^{\star}\in\mathbb{R}^{d}$ with $\beta^{\star}_{(S^{\star})^{c}}=0$ and $\min_{j\in S^{\star}}|\beta_{j}^{\star}|=\beta_{\min}>0$ such that
$\forall e\in\mathcal{E}\qquad\widetilde{\beta}^{(e,S)}=[\beta^{\star},0].$ (B.17)
Let $\varepsilon^{(e)}=Y^{(e)}-(\beta^{\star})^{\top}X^{(e)}$ be the noise, the above invariance equality (B.13) is equivalent to that both $X_{S^{\star}}$ and $\bar{\phi}(X_{S^{\star}})$ are uncorrelated with noise across all the environments, that is,
$\forall e\in\mathcal{E}\qquad\mathbb{E}[\varepsilon^{(e)}X_{S^{\star}}^{(e)}]=\mathbb{E}[\varepsilon^{(e)}\bar{\phi}(X_{S^{\star}}^{(e)})]=0$
Condition B.13 (Identification for Linear $\mathcal{G}$ and Augmented Linear $\mathcal{F}$).
For any $S\subseteq[d]$ with $\sum_{e\in\mathcal{E}}\mathbb{E}[X_{S}^{(e)}\varepsilon^{(e)}]\neq 0$, either (1) there exists some $e\in\mathcal{E}$ such that $\widetilde{\beta}^{(e,S)}\neq[\beta^{(e,S)},0]$, or (2) there exists $e,e^{\prime}\in\mathcal{E}$ such that $\beta^{(e,S)}\neq\beta^{(e^{\prime},S)}$.
For technical convenience, we also used truncated function class the discriminator class, defined as $\mathcal{H}_{\mathtt{alin}}(d,\phi,B)=\{\widetilde{f}=\mathrm{Tc}_{B}(f):f\in\mathcal{H}_{\mathtt{alin}}\}$.
Theorem B.11 (Linear $\mathcal{G}$ and Augmented Linear $\mathcal{F}$).
Suppose B.8, B.11B.13 hold, and we choose
$\mathcal{G}=\mathcal{H}_{\mathtt{lin}}(d,C_{2},C_{2}\sqrt{\log n})\qquad\text{and}\qquad\mathcal{F}=\mathcal{H}_{\mathtt{alin}}(d,\phi,2C_{2}\sqrt{\log n})$
with some constant $C_{2}\geq 2(\sigma_{\widetilde{x}}\lor 1)\max_{e\in\mathcal{E},S\subseteq[d]}\|\widetilde{\Sigma}^{1/2}\widetilde{\beta}^{(e,S)}\|_{2}$. Then, there exists some constant $\widetilde{C}$ that only depends on $(C_{1},C_{2},\sigma_{\widetilde{x}},C_{\widetilde{x}},\sigma_{y},C_{y})$ such that the FAIR least squares estimator using the above function classes and hyper-parameter $\gamma$ satisfying $\gamma\geq 8\gamma^{\star}_{\mathtt{LA}}=8\sup_{S:\mathsf{b}_{\mathtt{LL}}(S)>0}\mathsf{b}_{\mathtt{LL}}(S)/\bar{\mathsf{d}}_{\mathtt{LA}}(S)$, where
[…]
satisfies the $L_{2}$ error bound (B.15) with probability at least $1-\widetilde{C}n^{-100}$. Moreover, if $d=o((1+\gamma^{2})n/(\log^{6}n))$, for large enough $n$, the error bound (B.16) also holds with probability at least $1-\widetilde{C}n^{-100}$.
We can see that the proposed estimator utilizes both the heterogeneity among different environments and strong prior knowledge that the true regression function admits linear form to help the identification. It bridges the EILLS estimator in Fan et al., (2023) and the Focused GMM (FGMM) estimator in Fan & Liao, (2014) when the instrumental variables are $[X_{S},\bar{\phi}(X_{S})]$ and hence has some advantages over the individual ones. We illustrate this as follows.
  • 1.
    When there are multiple environments $|\mathcal{E}|>1$, the identification condition B.13 is weaker to both the EILLS and FGMM estimators. In particular, a consistent estimate $\beta^{\star}$ is attainable if incorporating variables $x_{j}$ with $\sum_{e\in\mathcal{E}}\mathbb{E}[X_{j}^{(e)}\varepsilon^{(e)}]\neq 0$ will result in either (1) a shift in the best linear predictor across environments or (2) the fitted residuals is strongly correlated with some nonlinear basis. We refer to this property as “double identifiable” property, given satisfying either condition can lead to the consistent estimation of the true parameter. Furthermore, the critical threshold $\gamma^{\star}$ can be smaller than that of the EILLS estimator according to the inequality $\bar{\mathsf{d}}_{\mathtt{LA}}(S)\geq\bar{\mathsf{d}}_{\mathtt{LL}}(S)$. This implies that the estimation is sample efficient, which allows for a small $\gamma$, if either the signal of nonlinear basis or the signal of heterogeneity is strong.
  • 2.
    If there is only one environment $|\mathcal{E}|=1$, it reduces to an estimator similar to the FGMM estimator. Consistent estimation remains feasible in this case but completely impossible for EILLS estimator. Moreover, the identification condition, in this case, resembles and relaxes that in Fan & Liao, (2014).
At the same time, it should be noted that the above advantages over the EILLS estimator (linear $\mathcal{F}$) are at the cost of imposing stronger invariance condition B.12, which assures that the noise should not only be uncorrelated with $X_{j}^{(e)}$ but also be uncorrelated with $\phi(X_{j}^{(e)})$ for any $j\in S^{\star}$ and $e\in\mathcal{E}$.
B.6.3 Neural Network $\mathcal{F}$
We impose some regularity conditions on the regression function.
Condition B.14.
There exists some constant $(C_{m},\sigma_{m})$ such that $m^{(e,S)}$ is $C_{m}$ Lipschitz and $|m^{(e,S)}(0)|\leq C_{m}$ for any $e\in\mathcal{E}$ and $S\subseteq[d]$ and
$\mathbb{P}(|m^{(e,S)}(X_{S}^{(e)})|\geq t)\leq C_{m}e^{-t^{2}/(2\sigma_{m}^{2})}\qquad\forall t\in[0,\infty)$
In this case, we consider the strongest invariance condition together with the weakest identification when the predictor function class $\mathcal{G}$ is linear.
Condition B.15 (Invariance in Linear $\mathcal{G}$ and Neural Network $\mathcal{F}$).
There exists some $S^{\star}\subseteq[d]$ and $\beta^{\star}\in\mathbb{R}^{d}$ with $\beta^{\star}_{(S^{\star})^{c}}=0$ and $\min_{j\in S^{\star}}|\beta_{j}^{\star}|=\beta_{\min}>0$ such that
$\forall e\in\mathcal{E}\qquad\mathbb{E}[Y^{(e)}|X_{S^{\star}}^{(e)}]\equiv(\beta^{\star})^{\top}X^{(e)}$ (B.19)
Condition B.16 (Identification for Linear $\mathcal{G}$ and Neural Network $\mathcal{F}$).
For any $S\subseteq[d]$ with $\sum_{e\in\mathcal{E}}\mathbb{E}[X_{S}^{(e)}\varepsilon^{(e)}]\neq 0$, either (1) there exists some $e\in\mathcal{E}$ such that $\mu^{(e)}(\{m^{(e,S)}\neq X^{\top}\beta^{(e,S)}\})>0$, or (2) there exists $e,e^{\prime}\in\mathcal{E}$ such that $\beta^{(e,S)}\neq\beta^{(e^{\prime},S)}$.
Theorem B.12 (Linear $\mathcal{G}$ and Neural Network $\mathcal{F}$).
Suppose B.8, B.14B.16 hold, and we choose the function classes $\mathcal{G}=\mathcal{H}_{\mathtt{lin}}(d,C_{2},C_{2}\sqrt{\log n})$ and $\mathcal{H}_{\mathtt{nn}}(d,\log^{d}n,\log^{d}n,C_{2}\sqrt{\log n})$ with some constant $C_{2}\geq(1\lor\sigma_{x}\lor\sigma_{m})\max_{e\in\mathcal{E},S\subseteq[d]}\|\Sigma^{1/2}\beta^{\star}\|_{2}$. Then, there exists some constant $\widetilde{C}$ that only depends on $(C_{1},C_{2},d,\sigma_{m},C_{m},\sigma_{y},C_{y},\sigma_{x},C_{x})$ such that the FAIR estimator using the above function classes and hyper-parameter $\gamma$ satisfying $\gamma\geq 8\gamma^{\star}_{\mathtt{LN}}=8\sup_{S:\mathsf{b}_{\mathtt{LL}}(S)>0}\mathsf{b}_{\mathtt{LL}}(S)/\bar{\mathsf{d}}_{\mathtt{LN}}(S)$, where
$\bar{\mathsf{d}}_{\mathtt{LN}}(S)=\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\|m^{(e,S)}-(\beta^{(S)}_{\dagger})^{\top}x_{S}\|_{2,e}\geq\mathsf{d}_{\mathtt{LA}}(S),$ (B.20)
satisfies, for large enough $n$,
$\qquad\|\beta_{\widehat{g}}-\beta^{\star}\|_{2}\leq\widetilde{C}(\log^{d+3}n)n^{-1/2}$
with probability at least $1-\widetilde{C}n^{-100}$.
The estimator can be viewed as an advanced version of the one using $\mathcal{F}=\mathcal{H}_{\mathtt{alin}}(d,\phi)$. It leverages neural networks to search for appropriate basis function $\phi$ with strong signals. With the proper choice of the neural network hyper-parameters, the estimator still maintains a parametric optimal rate (up to logarithmic factors). Additionally, it requires a weaker identification condition as described by B.16 and reduced critical threshold $\gamma^{\star}$ according to the inequality $\bar{\mathsf{d}}_{\mathtt{LN}}(S)\geq\bar{\mathsf{d}}_{\mathtt{LA}}(S)$ in Theorem B.12.
Appendix C Omitted Parts in Experiments
C.1 Pseudo-code of the Gradient Descent Ascent Algorithm
Algorithm 1 FAIR Gradient Descent Ascent Training
1:SGD Hyper-parameters: iteration steps $T$, batch size $m$, predictor/discriminator iter steps $T_{g}$/$T_{f}$.
2:FAIR Hyper-parameters: invariance regularization $\gamma$.
3:Annealing Hyper-parameters: Initial $\tau_{0}$ and final $\tau_{T}$.
4:Models: predictor $g(x;\theta)$, discriminators $\{f^{(e)}(x;\phi^{(e)})\}_{e\in\mathcal{E}}$, gate $w$.
5:Input: data $\{\mathcal{D}^{(e)}\}_{e\in\mathcal{E}}$ with $\mathcal{D}^{(e)}=\{(x_{i}^{(e)},y_{i}^{(e)})\}_{i=1}^{n}$ from $|\mathcal{E}|$ environments, loss function $\ell(\cdot,\cdot)$.
6:Output: Parameters of the prediction model: $w$ and $\theta$
7:
8:Initialize $\theta,\{\phi^{(e)}\}_{e\in\mathcal{E}}$ with random weights
9:Set $w=0$
10:
11:for $t\in\{1,\ldots,T\}$ do
12:     Set $\tau_{t}=\tau_{0}\times(\tau_{T}/\tau_{0})^{t/T}$
13:     for $t_{f}\in\{1,\ldots,T_{f}\}$ do $\triangleright$ Discriminator Ascent
14:         Sample $\{u_{i,1},u_{i,2}\}_{i=1}^{d}$ from $\mathrm{Gumbel}(0,1)$
15:         Calculate $a=(a_{1},\ldots,a_{d})$ with $a_{j}=V(w_{j},\tau_{t},u_{j,1},u_{j,2})$, where $V(\cdot)$ is defined in (6.1).
16:         for $e\in\mathcal{E}$ do $\triangleright$ Update $f^{(e)}$
17:              Sample minibatch of $m$ examples $\{(x^{(e,i)},y^{(e,i)})\}_{i=1}^{m}$ from $\mathcal{D}^{(e)}$.
18:              Update the discriminator by ascending its stochastic gradient:
$\nabla_{\phi^{(e)}}\frac{\gamma}{m}\sum_{i=1}^{m}\left[\{y^{(e,i)}-g(x^{(e,i)})\}f_{\phi^{(e)}}(x^{(e,i)})-\frac{1}{2}\{f_{\phi^{(e)}}(\phi^{(e)})\}^{2}\right]$
              where
$g(x)=g(a(w)\odot x;\theta)\qquad\text{and}\qquad f_{\phi^{(e)}}(x)=f(a(w)\odot x;\phi^{(e)})$
19:         end for
20:     end for
21:     for $t_{g}\in\{1,\ldots,T_{g}\}$ do $\triangleright$ Predictor Descent
22:         Sample $\{u_{i,1},u_{i,2}\}_{i=1}^{d}$ from $\mathrm{Gumbel}(0,1)$
23:         Calculate $a=(a_{1},\ldots,a_{d})$ with $a_{j}=V(w_{j},\tau_{t},u_{j,1},u_{j,2})$, where $V(\cdot)$ is defined in (6.1).
24:         for $e\in\mathcal{E}$ do $\triangleright$ Enumerate Environments
25:              Sample minibatch of $m$ examples $\{(x^{(e,i)},y^{(e,i)})\}_{i=1}^{m}$ from $\mathcal{D}^{(e)}$.
26:              Calculate loss as function of $\theta$ and $w$, that is
$L^{(e)}(\theta,w)$ $=\frac{\gamma}{m}\sum_{i=1}^{m}\left[\{y^{(e,i)}-g_{w,\theta}(x^{(e,i)})\}f_{w}(x^{(e,i)})-\frac{1}{2}\{f_{w}(x^{(e,i)})\}^{2}\right]$
$~{}~{}~{}~{}~{}~{}+\frac{1}{m}\sum_{i=1}^{m}\left[\ell\left(y^{(e,i)},g_{w,\theta}(x^{(e,i)})\right)\right]$
              where
$g_{w,\theta}(x)=g(a(w)\odot x;\theta)\qquad\text{and}\qquad f_{w}(x)=f(a(w)\odot x;\phi^{(e)})$
27:         end for
28:         Update the predictor weights $w,\theta$ by descending its stochastic gradient:
$\nabla_{(\theta,w)}\sum_{e\in\mathcal{E}}L^{(e)}(\theta,w)$
29:     end for
30:end for
C.2 Detailed Simulation Configuration
C.2.1 Linear Model with $p=15$
Data Generating Process.
The data-generating process is similar to that described in Section 6.2.1. We also let $|\mathcal{E}|=2$, and use the same procedure to generate parent-children relationship and structural assignment except that (1) we use $d=15$ and let the variable $Z_{8}$ be $Y$; and (2) we enforce that $Y$ has at least $3$ parents and $3$ children (3) the structural assignment for variable $Y$ is
$Y^{(8)}=Z^{(8)}\leftarrow\sum_{k\in\mathtt{pa}(8)}C_{8,k}Z_{k}^{(e)}+C_{8,8}\varepsilon_{8},$
that is we let the variance noise to be the same for the two environments. This is because we will include ICP in our simulation comparisons, which requires conditional distribution invariance.
Implementation.
We use the same configurations in the implementation of FAIR-GB and FAIR-RF. We also use fixed $\gamma=36$ for all the FAIR family estimators including EILLS. It is worth noticing that ICP, anchor regression, and IRM introduce an additional hyper-parameter, we pick it in an oracle way for them: that is, we enumerate all the candidate hyper-parameters and select the one that minimizes the $L_{2}$ estimation error. We report the performance for $n\in\{100,200,500,800,1000\}$.
Discussion of Results.
For anchor regression and IRM, their performance and the corresponding relationships w.r.t. Pool-LS are similar to the 12 variable illustrations in Fan et al., (2023). The anchor regression is almost the same as Pool-LS because it is essentially the same as standard least squares when the environments are discrete: indeed, in $|\mathcal{E}|=2$, it just runs least squares with a difference intercept for the interventional environment $e=1$. The IRM is better than vanilla least squares by slightly decreasing the bias, while the performance improvement is negligible compared with the bias it has.
For ICP, the performance is even worse than pooled least squares because it collapses to conservative solutions like 00. Note that we apply interventions to all the variables in environment $e=1$, under which it is possible for ICP to identify $\beta^{\star}$ and $S^{\star}$ when $n=\infty$. The large estimation error it depicts is due to its inefficiency in estimation.
We can also see that the performance of FAIR-BF and FAIR-RF are similar, demonstrating the effectiveness of our proposed gradient descent ascent algorithm with Gumbel approximation. The performance of FAIR-GF and FAIR-RF is slightly better than EILLS. This is because the FAIR estimator is essentially doing the most efficient pooled least squares when it selects the correct variable.
C.2.2 Nonlinear Model
Data Generating Process.
For the structural assignment, we let $\varepsilon_{i}^{(e)}=\varepsilon_{i}$ for $i\leq 5$ and $\varepsilon_{i}^{(e)}=C_{i,i}^{(e)}\varepsilon_{i}$ where $(\varepsilon_{1},\ldots,\varepsilon_{26})$ are independent $\mathrm{Uniform}([-1.5,1.5])$ random variables to let the covariates to be uniformly bounded and $C_{i,i}^{(e)}$ are scalars that are randomly generated in each trial. $\varepsilon_{0}$ is standard normal distributed that is independent of $(\varepsilon_{1},\ldots,\varepsilon_{26})$.
For the assignments for the children of $Y$, we let $f_{i,0}^{(e)}(u)=C_{i,0}^{(e)}\tanh(u)$, where $C_{i,0}^{(e)}$ are scalars that is randomly sampled from $\mathrm{Uniform}([-1.5,1.5])$ for $e=0$ and $\mathrm{Uniform}([-5,5])$ for $e=1$, the noise level $C_{i,i}^{(e)}$ is a scalar generated from $\mathrm{Uniform}([1,1.5])$. For the assignments for other variables $X_{i}$ with $i\geq 10$, we let $f_{i,j}^{(e)}(u)=C_{i,j}^{(e)}h_{i,j}^{(e)}(u)$ where $h_{i,j}^{(e)}$ are randomly picked from the function set $\{\tanh(x),\sin(x),\cos(x)\}$, the noise level $C_{i,i}^{(e)}$ is a scalar generated from $\mathrm{Uniform}([2,3])$. For $m_{1}^{\star}$, it is $\sum_{k=1}^{5}f_{0,j}(x)$ with $f_{0,j}(x)$ randomly picked from $\{\tanh(x),\sin(x),\max(0,x),x\}$.
Implementation.
For the FAIR-NN implementation using Gumbel approximation, we also run gradient descent ascent using the Adam optimizer using a learning rate of 1e-3, batch size $64$. The number of iterations is $70k$ for $m_{1}^{\star}$ and $80k$ for $m_{2}^{\star}$. In each iteration, one gradient descent update of the neural network parameters in $g$ and the Gumbel logits parameter $w$ is conducted followed by three gradient ascent updates of the neural network parameters in $f^{0}$ and $f^{(1)}$. We also use fixed $\gamma=36$. The implementation details for the estimators are:
  • (1)
    Pool-LS: it simply runs least squares on the full covariate $X$ using all the data.
  • (2)
    FAIR-GB: Our FAIR-NN estimator with Gumbel approximation, its prediction on the test dataset is evaluated by averaging the predictions over $100$ Gumbel samples.
  • (3)
    FAIR-RF: it first selects the variables $x_{j}$ in the fitted model in (2) with ${\mathrm{sig}(w_{j})}>t$, i.e., $\widehat{S}=\{j:{\mathrm{sig}(w_{j})}>t\}$, and runs least squares again on $X_{\widehat{S}}$ using all the data. Here we let $t=0.6$ for $n\leq 2000$ and $t=0.9$ for $n>2000$.
  • (4)
    Oracle: it runs least squares on $X_{S^{\star}}$ using all the data.
For FAIR-GB, we report the estimated MSE for the model in the last iteration. For other estimators, we also run gradient descent using Adam optimizer for 10k iterations. We report the estimated MSE for the model with early stopping regularization: that is, we report the estimated MSE of the model that has the smallest validation error, and the validation data is sampled independently and identically to the training data with sample size $n_{\mathrm{valid}}=\lfloor 3n/7\rfloor$.